A Complete Bibliography of Publications in the ACM Symposia on Theory of Computing (STOC) for

1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.05

Title word cross-reference

#P [BW91, SV99c]. #P-complete [BW91, SV99c].

(log n)^{2−o(1)} [GH95]. + 1 [DN95]. .879 [GW94a]. 1 [ANTSV99, Mul90]. 2 [DN95, GW94a, HdLT98, KP94c]. 3 [AK94, DG96, GK98, PSZ97, VV96]. 4/3 [CG91]. = [IW97]. ^{0}[p^{k}] [MP97]. E [IW97]. ϵ [CEG^{+}93, LV92]. j [Fei95]. K [Fre98b, Gro91, ARR98, AKTT97, AABV95, BKPS98, BCV96, BR92, BCV95, BRV96, BD95, CK92, CCGG98, CGTS99, CT91, CT96, cDF97, FW93, GH94, JL93, KRR91, KM94b, KP94b, Mul90, SZ93]. L_{2} [Ajt98]. ln n [Fei96]. log log N [AM92b]. m [Gro94, Sze90]. N [AM92b, BBBT97, CK92]. ≠ [Mul97]. O(congestion + dilation + log^{1+ϵ} N) [OR97]. O(log k) [GH94]. O(n^{2}) [BK96, KS93]. O(n^{3}) [Ben94]. O(n log^{3} m) [CH97]. O(n log n) [Ma96]. O(nm) [NI96]. ω [Saf92]. Ω(n^{2}) [GK97]. p [Ajt94]. R^{d} [MLS92, Mul90]. s [BR91, BK96]. SL ⊆ L^{4/3} [ATSWZ97]. st [EP95]. ⊆ [Nis92]. t [BR91, BK96, BG93, HS93a]. t + 1 [GM93]. W [BFH94].

- [BR91, BK96]. -approximation [GW94a]. -approximations [LV92]. -automata [Saf92]. -body [CK92]. -chains [FW93]. -CNF [BKPS98]. -colorable [AK94]. -competitive [BBBT97]. -connectivity [EP95]. -D [VV96]. -DNF [BCV96]. -edge [HdLT98]. -function [Fei95]. -level [DN95]. -median [CCGG98, CGTS99]. -medians [ARR98]. -MST [BCV95]. -nearest-neighbors [CK92]. -nets [CEG^{+}93]. -node [CT96]. -processor [AM92b]. -resilient [BG93, HS93a]. -sat [Fre98b]. -server [KRR91, KP94b, Gro91]. -set [cDF97, SZ93]. -shredders [CT96]. -size [Ma96]. -space [DG96]. -Steiner [BD95]. -term [BR92]. -time [AM92b]. -tours [AKTT97]. -trees [AABV95]. -vertex [CT91]. -way [ANTSV99]. -wise [KM94b].

ability [KM96a]. abstract [AMO96, Ajt96, AFS98, Ajt98, Ajt99, ABO96, ABAR99, ALMZ96, AFGV97, AF92, APSPS92, AKP92, AAFL96, ABKU94, BFR92, BPR96, Bea96a, Bea97b, BC92, BCK98, BT94, BC97, BRV96, BFH94, BCFM98, CO99, CRS94, CGTS99, CM99, CT96, CG97, CGP^{+}98b, DJP^{+}92, DMR99, DGS98, ECBD97, FL92b, FL92a, FKN94, FK97, Fei98, FS98, FY92, GIS92, GH95, GOP94, GMNS99, HR97, IK99, JS94, Kal92, Kil92, LY92, LV92, Lin92b, LV97, Ma96, MOOY92, MRST97, Mul97, Nao96, NR97, NS96, NR99, OS97, PY94, PSR94, PL98, PSTW97, Roj99, Saf92, STA97, STV99, TS96, Tre97, Tre98, Tre99, VY94, VR92, Wig94]. AC [Sit94]. ACC [MP97]. accelerated [VY94]. acceleration [MP98]. acceptance [Saf92]. access [AL94a, Mil94]. achieving [DMR99]. acknowledgment [DGS98]. acyclic [FS98]. Adapting [APSPS92]. Adaptive [AKOR98, Bea96a, Dam98, AAT94, DKKP95, Gam99]. Adaptively [CFGN96]. Adding [KP95]. Additive [KW95]. Adversarial [BKR^{+}96, AKOR98, Gam99]. advice [CBFH^{+}93, HS96]. affordable [FK94]. against [NY90, RS93, Yao95]. agreement [ADLS91, CR93, GM93, SZ93]. algebraic [AB92, Bas96, BPR96, GKV94, GKadHS96, HR98, Mul97, OS99, SW98]. algebraic-geometric [OS99, SW98]. algorithm [AKR91, AM92b, BBBT97, BRV96, Bod93, CKR98a, CGTS99, CY94, CEI96, DSC95, DR91, Fil97, Gal92, GH94, Gro91, Gro96, HP95, Kal92, KS93, KM94a, Kar96b, KVV90, KS99, KT94, Lin92b, LSW98, LN93, MS93b, NSS98, Sla96, ST97, Way99, WGMV93]. Algorithmic [NR99, Var97, MR98]. Algorithms [CKR98b, CT91, DG96, AB92, AMO96, AAT94, AW91, ALM90, AW96, AL94b, AS97b, BCF^{+}91, Bak93, BR91, BFR92, BFKV92, BDBK^{+}90, BDBK97, BC97, BGRS90, BGR93, BP90, BOR99b, Bsh98, BKRW98, CRS94, CCGG98, CT96, CS92, CM91, CFLS93, CDRS90, CadHSV96, CGG^{+}95, DI96, DHW93, EGIS93, FK96, FM91, FW93, Gab94, GKT99, GI91, GLM^{+}95, GKS90, GW94a, Gol93, Gro98, GKR^{+}99, Hag94, HK95, HdLT98, INW94, Ind99b, JS94, KK90, KKS^{+}99, KRR91, KM96a, KR95, KST90, KRRS94, KL96, Kle97, Kle98, KMvS94, La 94, LSM^{+}91, LNS98, MS92, McC96, MP94, OR97, Pan95, PS92, Rao92, RMR^{+}93, STA97, Sit94]. Aligning [JLW94]. all-optical [RU94]. all-pairs-shortest-path [Sei92]. Allocating [KRT97]. allocation [ABF93, BNCK^{+}95, BCPV93, CS92]. allocations [ABKU94]. Almost [TS98, AAFL96, Zwi98]. almost-satisfying [Zwi98]. Alpha [La 94]. Alpha-algorithms [La 94]. Alphabet [ABF92]. alternating [KV98]. alternation [TW98]. always [GJ97]. amazing [Wig94]. among [Pel94, dBMS93]. amortised [ABAR99]. amplification [DI99]. analog [Maa93]. analyses [CCG^{+}91, GLM^{+}95]. Analysis [LMSS98, BCK98, BFJ^{+}94, CJSW93, DI96, JS94, RS93, SS90a, Sla96]. ancestors [BKRW98]. ancient [BFKV92]. Angles [DV93]. angular [Mal92]. annual [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97, ACM98, ACM99]. any [DI99, LV98, LH97]. application [BFL96, BGG95]. applications [AST90, AS97a, AAF^{+}93, Bak93, BGLR93, BGLR94, BDBK97, BLR90, CK92, CRS93, CDRS90, DM91, FMdB99, GR99, GVY93, Ira97, Kal91, KP95, KMY90, LY94, MV91, Mic97, NN90, PRV99, WZ93, Zwi99b]. approach [AFS94, BCK98, BFU97, FMdB99, Gab91, HR98, OS99]. Approximability [KTW96, Aro98, BS94, GG98, KSW97, Tre97]. Approximate [AAMR93, GVY93, IM98, KMS99, LN90, Rei97, ABCG93, CCGL99, EKR99, GLR^{+}91, Hub98, KOR98, LSW98, Pat92, Ree92]. Approximately [LV97]. Approximating [ALS97, BNGNS99, BK96, Fei98, LR97b, RS98, AMS96, Bar98, Fei96, HR94, LY93, NW99, Pan95]. Approximation [ARR98, AK99, cDF97, MHSR94, STA97, AKR91, AKK95, AKTT97, AL94b, AABV95, BGLR94, BJL^{+}91, BCV95, BRV96, BOR99b, CKR98a, CCGG98, CGTS99, CM99, Coh94, CFLS93, GH94, GW94a, GKR^{+}99, JP99, JSOS99, JLW94, KM94a, Kar95, KR95, KST90, KS92, KL96, LSM^{+}91, LV91, LN93, NSS98, PR90, RMR^{+}93, SV99c, ST97, WGMV93]. Approximations [EGL^{+}92, KT95b, Mat91, BGLR93, CRS94, KV92, LV92, Sri95]. arbitrary [AZ99, Bar98, BKL96, Moh96, RT96]. arborescences [Gab91]. arguments [CD97, Kil92]. arising [JS94, Kar93, KL96]. arithmetic [AJ93, BHH92, GK98]. arrangements [AGR95]. array [AHHP94, AHP95, Die97, KLO96]. arrays [ABO92, BU96, CMS93]. aspects [MR98]. assembly [KD95, KD96]. assignments [Zwi98]. associated [MPR94]. assumptions [Hås90]. asymmetric [MNSW95]. Asymptomatic [GHK92]. Asymptotic [MP98]. Asynchronous [BOCG93, AAMR93, APSPS92, ACS94, BG93, CR93, HS93a]. Atlanta [ACM99]. attacks [NY90]. attribute [Blu90, Dam98, Ser99]. attribute-efficient [Dam98, Ser99]. augmentation [CT96, PSTW97]. Augmenting [Ben94]. authentication [BCK98]. automata [AI99, ANTSV99, CHPW94, FKR^{+}93, HK90, HKPV95, KV98, Saf92]. Automatic [ALMZ96]. Average [VR92, Wan95, ABV95, CCG^{+}91, CJSW93, JRS94, LK99]. Average-case [Wan95, CCG^{+}91, CJSW93]. axis [KR99]. axis-parallel [KR99].

Backing [BAP99]. Balanced

[ABKU94, FM92]. balancing [ACMR95, AAMR93, AAF^{+}93, GLM^{+}95, PW93, ST97]. ballot [BT94]. Baltimore [ACM90]. Bandwidth [BNCK^{+}95, ALMZ96, BKRV98, Fei98, KRT97]. banyans [RS98]. barrier [FW90]. base [BS92]. based [AMO96, Bea97b, BK93, EGIS93, FGY96, Mil96, MTTW95, SV93]. bases [BS99]. basis [CEI96, Lak90]. BCH [Pan97]. be [BBF98, GadHW90, KST90, NS93, Nao96]. beat [WZ93]. best [CJSW93]. Better [Alb97, JSV94]. Betti [Bas96, Fri96, Yao94]. between [BGIP99, CCG^{+}91, MNV94, Oka96]. Beyond [BFH94, SW98]. bias [NN90]. Biased [ABK^{+}92, BN93]. bibliography [Ang92]. Biconnectivity [KV92, HdLT98, Rau94]. bilinear [NW95]. bin [CCG^{+}91, CJSW93]. binary [ST90]. biology [Kar93]. bipartite [BGK^{+}96, KL98a, KVV90]. bipartiteness [GR98]. birds [RMR^{+}93]. Bit [DMR99, BHK98, BH98]. bits [FK95a, HS96]. black [HR98]. BLASTING [FW90]. block [VS90]. blocking [AZ98]. body [CK92]. boolean [Blu90, BBR92, EP96, KSW97, Luk99, NS92, Pat92, PSZ97, TN94]. boosting [KM96a]. both [CHPW94]. bottleneck [Erg95, KK90]. bound [ANTSV99, AHP95, BG91, CBDFS96, CCGL99, EP95, Fre90, GKadHS96, GK97, GK98, Pon95, SW98, WZ93]. bounded [AD91, BFH94, CKLW99, Cle90, DK99, DW92, FW94, GR97, GR98, HR94, MS92, Nis90, OW90, PY94, RR99, Sze90]. bounded-degree [HR94]. bounded-width [Cle90]. Bounding [CI95, HBW95, Bas96, Zuc90]. Bounds [ADLS91, Maa93, Alb97, Asp97, BGK^{+}96, BFL96, BIK^{+}92, BF99, BLY92, BPR95, BOR99a, CEG^{+}93, Cha95, Epp95, EP96, FS98, FK98, Gál98, GKV94, Gri98, IK99, KLM^{+}95, KRR91, KM95, KMRZ93, La 90, Mil94, Mul94, Nis91, PSZ97, PT93, RSZ99, YA94]. box [HR98]. BPP [IW97]. branching [Pon95, RWY97, Tha98]. breaking [ACS94, DMR99, MOOY92, SV94]. British [ACM92]. broadcast [FY95, KL98b, KZ98]. Bubbles [DKKP95]. bursty [AKOR98, KRT97]. butterfly [KNT94, MS92]. buy [CKLW99]. buy-and-hold [CKLW99]. Byzantine [CR93, GM93, MR97].

cabbage [HP95]. cache [BGR93, CK99b]. caching [CKZ99, Ira97, KLL^{+}97]. cactus [DN95]. calculus [BGIP99, RWY97]. California [ACM93]. Can [BD92, BRSU93, GadHW90, NS93]. Canada [ACM92, ACM94]. cannot [JL93]. canonization [Lin92b]. capacitated [CKR98b]. capacities [KST90]. carcass [DV94]. Carlo [BCF^{+}91]. carvings [KV92]. case [AD97, ABAR99, BR95, BS96, CCG^{+}91, CJSW93, JRS94, VR92, Wan95]. case/average [AD97]. categories [Sha91]. catenable [KT96]. catenation [KT95a]. central [BV97]. certificates [CT91]. chains [CLP99, CJSW93, DMR99, FW93, Fil97, LW95]. changing [ECBD97]. channels [Mau91]. characteristic [Bas96]. characteristics [ECBD97]. characterization [BC92, Gál98, KM96b, RS97]. characterizations [DFK^{+}99]. Characterizing [KW93, KOR96, BFJ^{+}94]. characters [BPWY96]. chasing [PRV99]. checkable [BGLR93, BGLR94, CFLS93, KPT97]. checkers [EKK^{+}98, Yao90]. Checking [ABCG93, BFLS91, LV98, FGY96]. checkpoints [BGRS90]. Chinese [GRS99]. Choosing [AFS94]. chosen [NY90]. chromatic [Fei95]. ciphertext [NY90]. Circuit [JRS94, AAF^{+}93, BFL96, CR96, CMadH^{+}98, KW93]. circuits [AKN98, AJ93, BNT95, GK93, GH95, GK98, Gro94, IPS93, IW97, KP94c, KOR96, MRST97, PZ92, PSZ97, PR93, RW90, Sze90]. clairvoyant [ECBD97]. classes [BDBK97]. classical [BCW98, Raz99]. classification [KSW97]. Clique [FM91, Hås96]. clock [PSR94]. closed [BRS91a, NTS95, SS90a]. closure [AFS98, KK90, KS99]. clustering [BOR99b, CCFM97]. CNF [BKPS98]. code [Hås96]. codes [LR97a, LMS^{+}97, LMSS98, OS99, Pan97, SW98, Spi95]. coding [AYZ94, ANTSV99, RS94, Sch93, Var97]. coefficients [BPR95]. Coherent [Yao90, Yao95]. coin [Asp97, BN93, NS96, RSZ99]. coin-flipping [Asp97, RSZ99]. coins [CKK^{+}95]. collapses [KvM99, Ogi96]. collecting [AABV95]. collective [RSZ99]. collide [AKR91]. Color [AYZ94]. Color-coding [AYZ94]. colorable [AK94]. coloring [AK94, Fu96, PS92, RSST96, SV93, KL96]. Columbia [ACM92]. Combinatorial [BV97, FW93, Fri96, Ind99a, KP95, Kar93, LLSZ93, Way99]. Combining [KPRS91, FSSW97]. come [McC96]. commitment [DIO98]. commitments [AAFL96]. Commodity [Bea97b, AL94b].

Commodity-based [Bea97b]. common [BKRW98]. Communication [FY92, Goo96, BHK98, BL92, BCW98, Die97, HS96, KNR95, KOR95, KLO96, Lin92a, MNV94, MNSW95, NS96, NW91, PRW97, PRV99, Raz99, Sch93, Sze90]. Communication-efficient [Goo96]. commutative [Nis91]. Compact [MPS.S99]. compacting [AHL90]. comparing [KLP^{+}97]. Comparison [BK93]. Comparison-based [BK93]. Competitive [AKP92, ABF93, BFR92, BRIS91, AL94b, BBBT97, Gro91]. competitiveness [AW96]. complement [NTS95]. complete [ARV94, BW91, DadH90, DT92, KSW97, MHSR94, SV99c]. completeness [BFH94, Kil91, Ko91, Wan95]. completion [SW99]. complex [Pan95, Rei97]. complexes [DG96]. Complexity [AV91, FHKM99, AMO96, AAI^{+}97, ABO96, AMS96, AHHP94, BBTN92, BHK98, BNT95, BL92, BCE^{+}95, BKPS98, BMR90, Bea96b, BV97, BMO90b, BV93, BS99, CCGL99, CR96, CK93, CGP^{+}98b, DJP^{+}92, Die97, DMR99, FY92, GOP94, GM95, Gri98, JRS94, JLW94, Kel92, KNR95, Kur94, KLO96, Lak90, LY94, MvzG94, Maa93, MNT90, MNSW95, MP94, NW99, NW91, NW95, PC99, PSY90, PY94, PRV99, Raz99, Roj99, Ser99, SS92, Sze90, TN94, Var97, Yao94]. composite [BBR92]. composition [BS92, BDBK97, Bsh98]. comprehensible [DW92]. compressed [FT95]. compression [FM91]. computability [DM91, HS93a, HS94]. Computation [AV91, ABO97, Bea97a, BOCG93, BGL^{+}93, BCW98, yC92, CFGN96, CO99, ER90, FKN94, FS97, FY92, GR99, HS94, Kap99, KPRR92, Nis91, Ost90, RS94, RR99, SV99b]. Computational [Ang92, Cha94, GOP94, Ser99, Bea96a, KP94c, MvzG94, Maa93, MNT90, RRS95]. Computationally [CG97]. computations [ABCG93, BFLS91, Bea96b, BL93, BG93, BSS97, KPS90, Nis90]. computed [NS93]. computer [CJSW93]. computers [CP90]. Computing [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97, ACM98, ACM99, AGR95, ABO92, BPR96, FPRU90, Fri96, KL90, Vor98, vzGS92, Bab97, Bas96, BS99, FL92b, Hag94, KPRS91, Kel92, Lak90, Ree92]. concave [AK91b]. concept [BDBK97]. concepts [BGM^{+}96]. Concurrent [DNS98, DW92, KST90, NBY97, Sha91]. condition [AF92, Saf92]. conditional [GHK92]. conductance [LK99]. congestion [CMadH^{+}98]. conjecture [CG91, Col90, DT92, KP94b, KLO96]. connected [CK93]. Connection [CKZ99]. connections [KRT97]. connectivity [BR91, Ben94, CT91, CT96, CDKT93, DV94, Edm93, EP95, Gab91, GI91, GH95, HdLT98, KR95]. conquer [Mat91]. consensus [Asp97]. considerations [KS96]. Consistent [KLL^{+}97]. Constant [Hag91, Jon93, ABO97, BMO90a, BCV95, BRV96, CGTS99, Gal92, GadHW90, MV91, MOOY92, RS97, Rei97, ST97]. constant-factor [BCV95, BRV96, CGTS99, ST97]. constant-space [MOOY92]. Constant-time [Hag91, Gal92]. constants [Pla99]. constrained [ART95]. constraint [FV93, KSW97, LV92]. constraints [KP95, KM93]. Constructing [BPWY96, BS91, KM93, LNS98]. Construction [Tre99, BFU92, DR91, Har94, KM94b, Kos94, KD95, KD96, LLSZ93, Mul90, NR97, SV94, WZ93, dBDS94]. constructions [NN90]. constructive [HS94, Zuc96]. containment [Mil96]. Contention [DHW93, MPR94, RU95]. continuous [BS96, CCG^{+}91]. continuous-motion [BS96]. control [AMO96, OR97]. Convergence [AMO96]. Converting [MV91]. convex [CEG^{+}93, Sch99b, dBMS93]. coordination [AHS91]. correcting [Pan97, Spi95]. correction [SW98]. correctness [AF92]. Correlated [Bea96b]. cost [BHK98, KP95, Way99]. Counting [AHS91, BW91, vzGKS93, Ajt94, Hub98, KP92, LV97]. cover [cDF97, Fei96, Sla96, Wil96, Zuc90]. Covering [AKTT97, KR99, Sri95]. criteria [ST97]. critical [PSTW97]. Cryptographic [Kha93, RS93, FGY96, Mau91]. cryptography [Bea97b, BGG95, DDN91, Mic97]. cryptosystem [AD97]. cryptosystems [NY90]. cube [CCGL99]. curse [IM98]. curve [BV97, MVO91]. curves [vzGKS93]. cut [CKR98a, GVY93, Kar94, KKS^{+}99, PT93, GW94a, Kar96b, KL96, Zwi99b]. cuts [BK96, DJP^{+}92, DN95, KS93, KM94a, Kar96a, MLS92, PP93, Rao92]. cutting [BPR95, Fu96]. cycles [AYZ94, BFS91]. cyclotomic [BSS97].

D [VV96]. daily [CKLW99]. Dallas [ACM98]. dark [Edm99]. data [BFR92, CDKT93, FG95, GIKM98, GHPT99, Kah91, KZ98, Mic97, MNSW95, Rau94]. database [Gro96]. debate [CFLS93]. Decidability [HK90, BT90, HR97]. decidable [HKPV95]. Decision [Fle93, Kle98, Yao94, BLY92, EP96, GKV94, GKadHS96, HR97, KM96a, KM91]. decodable [Spi95]. Decoding [SW98, OS99, Pan97]. decomposition [CK92, DT92, HR98, KPR93, PS92]. decompositions [Bod93]. dedicated [NW95]. defense [RS93]. definite [BKRV98, KPRS91]. deflection [BU96]. degree [AS97a, GR97, GR98, HR94, KRY94, NS92, Pat92, RS97, SSZ95]. degrees [BGIP99]. Delaunay [MTTW95]. delay [AZ99, DGS98, Gol93]. delays [HBW95, RU95]. Dense [ANTSV99, AKK95, AS97b]. density [LMSS98]. dependable [KPRS91]. dependencies [FKST93]. depends [VY94]. Depth [AJ93, GH95, GK98, IPS93, KP92, KP94c, PSZ97, Pla92, RW90]. Derandomization [KM94a, LV98]. derandomizing [IW97]. derived [KSW97]. Descriptive [GM95]. Design [DK99, BCK98, Ind99a, Kar94, NR99]. designs [LMSS98]. detect [BD92]. detecting [BBF98]. Determinism [Ajt99]. Deterministic [BR91, CR96, CP90, NI96, Sch93, Vis90, AM92b, CCGG98, HdLT98, KW93, LSW98, LV91, MS93b, NZ93]. determinization [Saf92]. diagrams [AHL90, Mul90]. diameter [KT95b, MS93b]. dictionary [DadH90]. Diego [ACM93]. Digital [DLN96]. digraphs [KK90]. dimension [AM92b, KM95, LLSZ93, RS95, Vor98]. dimensional [ABF92, BKL96, BOR99a, BOR99b, CK92, CK93, Gup99, KOR98, Lak90, MS93b]. dimensionality [IM98]. dimensions [Kle97, MT90, MTTW95, Roj99]. Diophantine [Roj99, VR92]. Direct [PRW97]. directed [BFR^{+}98, EP95, MRST97]. discreet [SS90b]. discrepancies [CCG^{+}91]. discrete [BSS97, CCG^{+}91, KMR^{+}94, SS90b]. disjoint [AKW96, BFU92, GKR^{+}99, KT95b, Kle98]. disk [AGL98, KL98b, VS90]. disks [BS96]. dispatching [FMdB99]. dispersers [SSZ95, TS98]. displacement [OS99]. distance [BGL^{+}93, DK99, Var97]. distinct [BGIP99]. Distinguishing [ACY95]. distribute [DadH90]. Distributed [RT96, AW96, Asp97, AKP92, ABF93, ACS94, BFR92, CS92, ER90, FMY98, HR97, KLL^{+}97, KLadH92, PS92, RS94]. distribution [BR95, BGM^{+}96, Kha93]. distribution-free [BGM^{+}96]. distribution-specific [Kha93]. distributions [CRS94, CCG^{+}91, EGL^{+}92, KMR^{+}94]. divide [Mat91]. divide-and-conquer [Mat91]. Dixon [KS96]. DNA [KD95, KD96]. DNF [BCV96, BR92, BFJ^{+}94, Bsh96, LV91]. do [Bei92, DSC95, JL93, Jon93]. Dobrushin [PL98]. does [GJ97]. doesn’t [BH98]. domination [FW94]. dominators [BKRW98]. down [KT95a, KM96a]. dual [WGMV93]. Dynamic [BU96, GT91, AAMR93, APSPS92, AL94b, BFU97, CCFM97, Col90, DKKP95, DGS98, EGIS93, FG95, GI91, GIS92, HK95, HdLT98, KS99, PU99, Rau94, SV99a]. dynamical [ARV94].

easier [FK98, KL98a, Nao96]. easy [BBF98]. edge [BFU92, CY94, DN95, Gab91, GI91, GKR^{+}99, HdLT98, NI96, Rao92]. edge-cuts [DN95]. edge-disjoint [GKR^{+}99]. edge-splitting [NI96]. Effective [Kal91]. Efficient [ACS94, BGLR93, BGLR94, CS92, FK96, FKR^{+}93, Gab94, GMNS99, Kap99, KLadH92, Kea93, KPS90, KPRR92, KL96, KOR98, LNS98, LLSZ93, LW95, Ost90, RU94, VV96, BP92, BL93, CD97, Dam98, DR91, DW92, FMY98, Goo96, KK90, Kil92, KD95, KD96, NN90, OS99, ORV94, Rei95, SV99a, Ser99]. Efficiently [RSST96]. eigenproblem [PC99]. eigenvalue [WZ93]. Eigenvalues [CY97]. eighth [ACM96]. eights [HI95]. election [CL93, ORV94, RSZ99, Zuc96]. elections [BT94]. element [CRS93]. elliptic [MVO91]. embedded [PL98]. Embedding [Gup99, Moh96, KKS^{+}99]. embeddings [Fei98, Sha91]. emptiness [KV98]. enclosure [Mil96]. encodable [Spi95]. end [AZ99, KOR95]. end-to-end [AZ99, KOR95]. enforcing [DLN96]. engines [SV99b]. Entropy [KK92, PRV99]. environments [AH97]. equality [ST90]. equality-testing [ST90]. equals [GSV98]. equation [JS94, Pan97]. equations [ABO96, Bar92, Ko91, MP98, Pla99]. equivalence [AD97, HK90, Luk99]. equivocation [Bea96a]. Erdos [Bab97]. error [ABO97, DFK^{+}99, FK94, KMS99, Pan97, RS97, RRV99, SW98, Spi95]. error-correcting [Pan97, Spi95]. error-correction [SW98]. error-probability [DFK^{+}99, RS97]. errors [AD91, BKL96, BK93, GRS99]. essential [BIKM99]. estimates [BLY92]. estimators [SV99c]. Euclid [Tre97]. Euclidean [ARR98, ADM^{+}95, DG96, Gup99, LMS98]. Euler [Bas96]. Evaluation [Nao96, NP99, PU99, Rei97]. even [MRST97]. every [AAFL96]. evolution [FK96]. evolutionary [BPWY96, FKW93, JLW94, KLP^{+}97]. Exact [Hub98]. exchange [BCK98]. Excluded [KPR93, AST90]. exclusion [KMRZ93, LN90, YA94]. execution [ECBD97]. executions [KPRS91]. Existence [BFU92]. expander [BFU92, BFU97, LR97a]. Expanders [WZ93, BCL94]. expansion [Bab91]. expansive [LS96]. expert [CBFH^{+}93]. Explicit [SSZ95, DR91, WZ93]. Exploiting [CK99b, BCL94]. Exploring [AH97, BFR^{+}98]. Exponential [BIK^{+}92, PSZ97, Raz99, Saf92, GK98, IW97, Ko91]. exponentiated [KW95]. expressive [ABFR91]. extended [AMO96, Ajt96, AFS98, Ajt98, Ajt99, ABO96, ABAR99, ALMZ96, AFGV97, AF92, APSPS92, AKP92, AAFL96, ABKU94, BFR92, BPR96, Bea96a, Bea97b, BC92, BCK98, BT94, BC97, BRV96, BFH94, BCFM98, BS91, CO99, CRS94, CGTS99, CM99, CT96, CG97, CGP^{+}98b, DJP^{+}92, DMR99, DGS98, ECBD97, FL92b, FL92a, FKN94, FK97, Fei98, FS98, FY92, GIS92, GH95, GOP94, GMNS99, HR97, IK99, JS94, Kal92, Kil92, LY92, LV92, Lin92b, LV97, Ma96, MOOY92, MRST97, Mul97, Nao96, NR97, NS96, NR99, OS97, PY94, PSR94, PL98, PSTW97, Roj99, Saf92, STA97, STV99, TS96, Tre97, Tre98, Tre99, VY94, VR92]. extension [PL98]. extensions [BW91, McC96]. external [AFGV97, FG95, VV96]. extra [Bei92]. Extracting [RRV99, TS96]. extractors [RRV99, Tre99, Zuc96]. Extremal [BGK^{+}96].

face [AAFL96]. faces [AGR95]. facility [STA97]. factor [BCV95, BRV96, CGTS99, Sha93, ST97]. Factoring [Adl91, KS95, vzGS92]. factors [BS92, Jon93, Pan95]. fairness [BCPV93, Saf92]. false [KLO96]. families [Gol93, GW94b]. Fast [BCF^{+}91, BR92, BSS97, CR93, CT96, CL93, EKR99, HI95, KMvS94, LSM^{+}91, McC96, ADT95, Gro96, Gro98, KPRS91, Vis90]. Faster [KRRS94, LK99, Pan97, Rao92, Spi96, KST90]. Fault [ABO97, KP94a, Lin92a, CS92, KNT94, LNS98, Ma96]. Fault-tolerant [ABO97, KP94a, LNS98, Ma96]. faults [CMS93]. faulty [ABO92]. Feasibility [Bar92]. feasible [ABO96]. field [LLMP90, LV98, MVO91]. fields [BS91, CK92, KS95, Sho90]. fifth [ACM93]. file [ABF93]. fill [NSS98]. fill-in [NSS98]. financial [CKLW99, PZ92]. find [ABAR99, AW91, CEI96, Kar97, La 90, Mil94]. Finding [BFS91, KL98a, LMW99, PP93, Ree92, Zwi98, AYZ94, Bod93, FKW93, Gab91, KT94, KMvS94, Rao92]. finger [Col90]. finite [AI99, Bab91, BD92, BS91, CHPW94, FKR^{+}93, HK90, KS95, MVO91, Sho90, YL91]. Finiteness [MS93a]. first [ACM99, GHK92]. first-order [GHK92]. fit [CJSW93]. fixed [AM92b]. flipping [Asp97, RSZ99]. flips [NS96]. floorplans [He97]. flow [AMO96, AL94b, AALR99, GVY93, Kar94, KTW96, KST90, KPR93, Kle98, LSM^{+}91, LR97b, McC96, PW93, PT93, Way99]. flows [CY97, KP95, Kar97, KL98a, PT93]. FLP [BG93]. folding [CGP^{+}98b, HI95]. forbidden [DR91]. forest [BS92]. forms [Kal91, NW95]. formulae [Bsh96]. formulas [BKPS98, BCV96, BR92, BHH92, Cle90, Fu96]. formulation [MTTW95]. forward [CadHSV96, SV99a]. four [LV97, RSST96]. four-coloring [RSST96]. Fourier [Bea97a, BFJ^{+}94, BT95, BSS97, HH99, KM91]. fourth [ACM92]. framework [Gro98]. free [ADT95, AW91, BT94, BGM^{+}96, HS94, LH97, SZ93]. Frege [MP97]. frequency [AMS96]. Frobenius [vzGS92]. full [ORV94]. Fully [GI91, GIS92, GM93, FG95, HdLT98, Kar95, KS99, Rau94]. fully-dynamic [FG95, HdLT98]. function [DDFY94, DI99, Fei95, FGY96]. functional [JS94, KT96]. Functions [Sze90, AK91b, BBR92, BIKM99, BC92, BD92, Blu90, BT95, CMR98, CFLS93, Erg95, FN91, GLR^{+}91, GW94b, Ind99a, LLW91, Luk99, MvzG94, NS92, Pat92, Rom90, Yao90]. Fundamental [CCG^{+}91]. Further [MR98]. FUSION [FW90].

game [KMvS94]. games

[DM91, FK97, FW94, Kil91, LY94, NS96]. gaps [BGIP99]. gates [Bei92, Gro94, KP94c, Sze90]. GCD [PRW97, Rei95]. General [KLP^{+}97, AKTT97, BGM^{+}96, CM99, EGL^{+}92, GSV98, Kil91]. Generalized [BG93, AKR91, Way99, WGMV93]. Generating [Ajt96, Wil96]. generation [Bab91, FMY98, MTTW95, Nao96, PL98, Sha93]. generator [Erg95]. generators [Hås90, Nis90, Sit94, STV99, Tre99]. Generic [AV91]. genetics [RRS95]. genome [Kar93]. genus [DR91]. geodesic [Kap99]. Geometric [Epp95, BDBK97, BRS91b, BGM^{+}96, FMdB99, KKS^{+}99, LNS98, Mat91, OS99, SW98, Tre97]. geometrical [RS98]. geometry [Cha94, Roj99]. Georgia [ACM99]. given [KM93]. global [ST97]. Goemans [Kar96b]. Goemans-Williamson [Kar96b]. good [ADM^{+}97, HR94, Kar96b]. gradient [KW95]. Graph [DT92, KvM99, Sch99a, BFL96, DV94, DN95, EGIS93, FM91, FHKM99, Fei95, HK95, KV92, SV93]. graphs [AST90, ACG93, AYZ94, AK94, Bab91, BGK^{+}96, BF93, BFR^{+}98, BD95, BFU92, BFU97, CK99a, CGP98a, CY97, CDRS90, DV93, Gab94, Gol93, GR97, GR98, HR94, Har91, He97, Kar97, KL98a, KRS93, KRRS94, LMSS98, Mal92, Moh96, NI96, PP93, PL98, Rao92, RS98, RS95, RSST96, Spi96]. grasping [KMY90]. Greed [HR94]. greedy [Sla96]. grid [MPS.S99]. grids [MPS95]. Gröbner [Lak90]. Groebner [CEI96]. group [CCGG98]. groups [BCF^{+}91, Bab91, BS92, Bea97a, yC92, Fre98b, KL90, Wan95]. guarantees [AABV95].

half [Kle98]. half-disjoint [Kle98]. Ham [MLS92]. Ham-sandwich [MLS92]. Hamiltonian [BFS91, Har91]. Hamming [CCGL99, Tre97]. hard [Ajt96, Ajt98, AKK95, Aro98, CFLS93, Fu96, Sha93]. Hardness [CNS99, BFH94, DI99, FL92b, GKR^{+}99, Hås96, Kha93, LY93]. hardness-amplification [DI99]. hardware [BRSU93]. harmonic [Gro91]. hash [CMR98]. hashed [GadHW90]. hashing [ADM^{+}97, GW94b, IMRV97, KLL^{+}97, LS96, MNT90, MV91, SS90a]. heads [JL93, JSV94]. heap [Fre90]. heaps [Fre98a]. heat [SV99b]. help [AK91a, BHK98, Bei92, BH98, BRSU93]. help-bit [BH98]. heterogeneous [BNGNS98]. Hidden [Mul91, AL94a, BFS91]. hiding [ALMZ96, FS90]. hierarchical [BCL94]. hierarchies [LH97]. hierarchy [BFH94, CNS99, KvM99, Ogi96, Sch99a, Tha98]. high [ALMZ96, AF92, BOR99a, BOR99b, DKKP95, KT95b, Kle97, KOR98, LLSZ93, MV91]. high-diameter [KT95b]. high-performance [AF92]. high-speed [DKKP95]. history [Sip92]. hitting [LLSZ93]. hold [CKLW99]. holographic [PS94]. Holyer [DT92]. Honest [GSV98, CO99]. honest-looking [CO99]. Honest-verifier [GSV98]. hot [KLL^{+}97]. hybrid [HKPV95]. hydrophilic [HI95]. hydrophobic [HI95]. hydrophobic-hydrophilic [HI95]. hyper [ALS97]. hyper-rectangles [ALS97]. hyperbolic [yC92]. hypercube [CP90]. hypercubic [Pla92]. Hypergraph [Luk99, Kel92]. hypergraphs [FY95]. hypothesis [BS91].

I/O [VS90]. ideal [Lak90]. identities [LV98]. if [CO99, IW97]. immunity [CL93]. implications [Fre98a]. Impossibility [FK95a, BG93]. impossible [SZ93]. Improved [AS97a, AL94b, AABV95, BS94, CRS94, CEG^{+}93, CM91, FS98, IK99, JP99, KR95, MV97, PS92, PT93, Rau94, Sri95, CKR98a, Gál98, LMSS98]. inapproximability [Hås97]. inclusion [LN90]. inclusion-exclusion [LN90]. Incremental [BGG95, CCFM97, DV94, DN95, La 94, dBDS94]. independence [Ajt94, Wig94]. independent [ABF92, BS99, BCFM98, EGL^{+}92, HR94, KM94b, Kel92, Mau91]. indexed [KZ98]. indistinguishable [FS90]. inequalities [CM91]. inequality [CM91]. Infinite [DM91, Blu90, Har91]. influence [Bab97]. Information [Fre98a, AL94a, BIKM99, BGL^{+}93, CBDFS96, CCFM97, CG97, CL93, DLN96, FPRU90, Fre90, FW90, GIKM98, IK99, ORV94, OS97, RSZ99]. information-theoretic [CBDFS96, IK99]. inhibition [Phi93]. input [AZ98]. instances [Ajt96, AKK95]. integer [Hag91, Pon95]. integers [Adl91, BSS97]. Integral [Ko91]. integration [AK91b]. interactive [DIO98, Sch93]. interconnection [CMadH^{+}98]. interior [VY94]. Interpolation [Ind99a, BT90]. interruptible [Fil97]. intersection [BRS91a]. intractability [VR92]. inverting [FK96, FN91]. investments [PZ92]. irrational [CK97]. irreducibility [Kal91]. irregular [LMSS98]. isolation [CRS93]. isomorphism [Luk99, Spi96]. issues [CK93].

Jacques [NW95]. JAG [Edm93]. job [AKP92, CM99, FS98, JSOS99, SW99]. jobs [ECBD97, FKST93]. jumping [NBY97]. June [ACM95].

key [AD97, BR95, BCK98, FMY98, NY90, Pan97]. keys [GadHW90]. knapsack [GK97]. know [DSC95]. Knowledge [ABV95, Bea96a, BMO90a, BMO90b, BP92, CD97, DNS98, GOP94, GSV98, Kil92, KPT97, Oka96, SZ93].

language [Sha91]. languages [Sha91]. lanky [ADM^{+}95]. Laplacians [Fri96]. Large [KD95, KD96, AYZ94, LY94, Sch99b]. Large-scale [KD95, KD96]. latency [ALMZ96, AK99, BCC^{+}94]. lattice [Ajt96, BS99, GG98, KV97, KRS93]. layout [AKW96]. layouts [MPS.S99]. lazy [dBDS94]. leader [CL93, ORV94, RSZ99, Zuc96]. leader-election [CL93]. learn [HRPW95]. learnability [Bsh96, KMR^{+}94]. Learning [BCV96, Blu90, BHH92, KM91, Ang92, AFS94, AL94a, ALS97, BDBK97, BR92, BFJ^{+}94, BGM^{+}96, Bsh98, CBDFS96, Dam98, FKR^{+}93, Kea93, KM96a, Kha93, LLW91, Maa93, Ser99, Sit94]. least [BKRW98]. Leighton [KST90]. Leighton-Rao [KST90]. lemma [IW97, MR98, STV99]. Lempel [FT95]. length [GY95]. level [DN95, MPS.S99]. levels [Mul90]. Lifting [CLP99]. lightest [Zwi99a]. limited [BNT95, HS96, SS90a, TN94]. limits [GG98]. line [ALM90, AAF^{+}93, BFL96, BC97, BGR93, Cha95, CDRS90, KVV90, LLW91]. Linear [BLY92, BJL^{+}91, BKRW98, BGIP99, CD97, PY93, Spi95, AB92, AM92b, Ajt99, ABO96, ADM^{+}97, AHNR95, BS92, Bod93, BW91, CM91, Coh94, CL93, Die97, Erg95, FS97, Kar96a, KW95, KT94, KOR96, KLO96, LLW91, LN93, Mil96, Moh96, Mul94, PS94, RW90, dBMS93]. linear-array [Die97, KLO96]. Linear-time [BKRW98, Spi95, KT94]. linearity [Tre98]. linearly [AD91, BS99]. linked [BAP99]. listing [Gol93]. lists [BAP99, KT95a, KT96]. load [ACMR95, AAMR93, AAF^{+}93, GLM^{+}95, PW93]. Local [Bab91, AL94b, DI96, cDF97, GLM^{+}95, GKS90, MR98, OR97, PSY90, ST97, Vor98]. Locality [IMRV97, SV93, BRIS91, FK95b]. Locality-preserving [IMRV97]. locally [NS93]. locating [BGRS90]. location [GT91, Pel94, STA97]. Log [KOR95, AK91b, SS90b]. log-concave [AK91b]. Log-space [KOR95]. logarithmic [ABV95, CP90, HdLT98, Pla92]. logarithms [MVO91]. logic [GHK92]. logics [Val99]. logspace [Lin92b, NZ93, NTS95]. long [Hås96]. looking [CO99]. loss [LMS^{+}97]. loss-resilient [LMS^{+}97]. Louisiana [ACM91]. Lovasz [Fei95]. Low [KRY94, AS97a, CMadH^{+}98, FK94, Gup99, LMSS98, RS97, Roj99]. low-congestion [CMadH^{+}98]. low-degree [AS97a, RS97]. Lower [Asp97, BFL96, BPR95, BOR99a, Cha95, EP96, GKV94, KLM^{+}95, KRR91, KMRZ93, La 90, Mil94, Mul94, Nis91, RSZ99, ANTSV99, AHP95, BGK^{+}96, BIK^{+}92, BG91, CCGL99, EP95, Epp95, FK98, Gál98, GKadHS96, GK97, GK98, Gri98, PSZ97, Pon95, Zuc90]. Luby [NR97]. Luby-Rackoff [NR97].

machine [AAF^{+}93, BKRW98, KLadH92, KTW96]. machines [ACY95, BNGNS99, CKK^{+}95, JP99, La 90, Lei92, LR97b, Mil94, SW99, YL91]. made [ADT95, BSW99]. maintaining [KS99]. maintenance [DV94, DN95]. majority [Bei92, GK93]. Majorizing [SV99c]. Makespan [JSOS99]. Making [AAFL96, BP92, FK97]. malleable [DIO98, DDN91]. management [BFR92]. manifolds [DG96]. Many [MPS95, RMR^{+}93, HRPW95, LMW99]. Many-to-one [MPS95]. map [CGP98a]. Mapping [Kar93, BFR^{+}98]. maps [Sch99b, vzGS92]. markets [CKLW99]. Markov [CKR^{+}99, CLP99, CJSW93, Fil97, LW95]. Maryland [ACM90]. matching [ABF92, Bak93, BG91, CRS93, CH97, FT95, FMdB99, GKT99, Gal92, IS93, JL93, KL98a, KVV90, Kos94, RW90, Vis90]. Matchings [KRS93, ACG93]. matrices [FL92b]. Matrix [HS93b, ABO96, LSW98, PC99, PY93, VR92]. matroid [Epp95, Gab91]. matroids [FM92]. matter [Jon93]. MAX [Kar96b, GVY93, PT93, GW94a, KL96, Zwi99b]. max-flow [GVY93, PT93]. maxima [GKS90]. maximal [AS97b, Kel92]. maximization [KSW97]. Maximum [FW93, ER90, GKT99, Kar97, KL98a, McC96, Odl94]. May [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97, ACM98, ACM99, Nao96]. measurements [Yao95]. mechanical [Gro96, Gro98]. mechanism [NR99]. median [CCGG98, CGTS99, NW99]. medians [ARR98]. meets [Tre97]. membership [AK91a, BCV96, GKV94]. memory [AFGV97, DHW93, KLadH92, NW95, VV96]. Menger [GY95]. merging [Hag94]. mesh [AKW96]. message [Lei92]. method [AYZ94, FMdB99, MTTW95, VY94]. methodology [CGH98]. Methods [Lei92, ALMZ96]. metric [Cla97, HS93b, Ind99b]. metrical [BBBT97]. metrices [Bar98]. metrics [Bar98, Gup99, LMS98]. Metropolis [DSC95]. might [BBF98, KST90]. migration [AALR99]. Min [BCFM98, GVY93, PT93]. min- [GVY93]. min-cut [PT93]. Min-wise [BCFM98]. minimal [FKN94, Mil96]. minimization [JSOS99, LY92, LY93]. Minimizing [AGL98, AALR99, KTW96, SW99]. Minimum [Kar96a, AK99, AABV95, BK96, BCC^{+}94, BKRV98, DN95, GH94, Hag94, HdLT98, KS93, KM94a, KKS^{+}99, KT94, LV92, NSS98, PP93, Var97, Way99]. minimum-quotient [PP93]. minimum-weight [AABV95]. minor [AST90]. minors [KPR93]. missing [BHK98]. mix [GJ97]. mixed [AKN98]. mixing [CLP99, LK99]. mod [Sze90, Gro94]. model [BKL96, DN95, EP95, FKW93, FKN94, HI95, Kah91, ORV94, RSZ99]. models [PRW97]. moderate [ART95]. Modified [PR93]. Modular [AW96, Fu96, BCK98]. modulo [Ajt94, BBR92, BGIP99, KP94c]. Molecular [SV99b, Kar93]. moments [AMS96]. monadic [FV93, AFS98]. Monotone [FV93, GH95, RW90, BGK^{+}96, BT95, Gál98]. Monte [BCF^{+}91]. Montreal [ACM94]. Morgenstern [NW95]. Motion [ART95, BS96, Kah91, KRR91, Pel94]. motion-planning [KRR91]. moving [Mul91]. MST [BCV95, BRV96, BKRW98, Tre97]. much [BRSU93]. Multi [CMS93, FMdB99, AHS91, AL94b, CK92, CFGN96, GVY93, Ira97, MPS.S99, RMR^{+}93]. multi-commodity [AL94b]. multi-dimensional [CK92]. multi-level [MPS.S99]. Multi-method [FMdB99]. multi-objective [RMR^{+}93]. multi-party [CFGN96]. multi-processor [AHS91]. Multi-scale [CMS93]. multi-size [Ira97]. multibutterflies [MV97]. multicast [AS97b]. Multicasting [BNGNS98]. multicommodity [KP95, KPR93, LSM^{+}91, PT93]. multidimensional [IMRV97]. multifingered [KMY90]. Multiple [IS93, BNGNS99, KP95]. multiplication [PZ92, Pon95]. multiplicity [HK90]. multipointer [FK95b]. multiprocessor [CM99, ECBD97]. multiprocessors [AF92]. multistage [CMadH^{+}98]. multitape [HK90]. multithreaded [BL93]. multivariate [Erg95, MP98, Sha93]. multiway [CKR98a, DJP^{+}92, KKS^{+}99]. mutual [KMRZ93, YA94].

narrow [BSW99]. Natural [RR94]. Navigating [BRS91b]. navigation [AWZ96]. NC [KM94a, Mul97]. Near [GKR^{+}99, AK91b, CBDFS96, CY94, Coh94, Kar96a, Rei97]. near-linear [Coh94, Kar96a]. Near-optimal [GKR^{+}99]. Nearest [Cla97, BOR99a, CK92, CCGL99, IM98, Kle97, KOR98]. nearest-neighbor [CCGL99, Kle97]. Nearly [PS94, BS92, CP90, EP95, MV91, Pla92]. nearly-constant [MV91]. Nearly-linear [PS94]. necessary [Rom90]. necessity [BP90]. needed [HRPW95]. negation [BNT95, TN94]. negation-limited [BNT95, TN94]. neighbor [BOR99a, CCGL99, Cla97, Kle97, KOR98]. neighborhoods [Sch92]. neighbors [CK92, IM98]. nets [CEG^{+}93, Maa93]. network [ALM90, DadH90, GHPT99, INW94, Kar94, Kar95, KPR93, Ma96, PS92, Phi93, PW93, Pla92, WGMV93]. networks [AKR91, AAMR93, ALMZ96, AZ98, AHS91, APSPS92, AL94b, BNGNS98, CMadH^{+}98, DK99, DKKP95, ER90, Gam99, HBW95, KLM^{+}95, KM95, KT95b, KP92, Lin92a, MS93a, MS92, MPS.S99, RT96, RU94, TN94]. neural [KM95, Maa93, MS93a]. Nevada [ACM95]. NEXPTIME [Pla99]. ninth [ACM97]. NNJAG [EP95]. nobody [CO99]. Node [AKW96, CT96]. Node-disjoint [AKW96]. Noether [Kal91]. Noise [BGM^{+}96, CBDFS96, Kea93]. Noise-tolerant [BGM^{+}96, CBDFS96, Kea93]. noisy [EP96]. Non [DIO98, DDN91, ECBD97, LS96, MP94, Ajt99, BS94, Gam99, GG98, KW93, Nis91]. non-adaptive [Gam99]. non-approximability [BS94, GG98]. Non-clairvoyant [ECBD97]. non-commutative [Nis91]. non-determinism [Ajt99]. non-deterministic [KW93]. Non-expansive [LS96]. Non-interactive [DIO98]. Non-malleable [DDN91, DIO98]. Non-standard [MP94]. nonadaptive [Dam98]. nonapproximability [KTW96]. nonblocking [ALM90]. noncommutative [AJ93]. Nondeterministic [HS96, ACY95, BL92, CHPW94]. nonisomorphism [KvM99]. Nonmonotonic [Fei99]. nonpreemptive [CG91]. nonresidues [BS91]. note [CD97, Kil92]. notion [BCPV93]. NP [AFS98, Ajt98, AKK95, Aro98, BCE^{+}95, BBF98, BFH94, DFK^{+}99, OW90, RS97, Sip92]. NP-completeness [BFH94]. NP-hard [Ajt98, AKK95, Aro98]. NPC [DT92]. number [HS96, LLMP90]. numbers [Adl91, AB92, BBR92, Bas96, CK97, Fei95, Fri96, GM95, Yao94]. numerical [BLR90, MTTW95].

O [VS90]. objective [RMR^{+}93]. Oblivious [Mic97, NP99, Ost90]. obstacles [ART95, dBMS93]. Occam [BP90]. off [AKW96, Cha95, Gab94, GW94b, Gro94]. off-line [Cha95]. offs [Edm93, IPS93, MNV94]. often [JS94]. old [Fle93, PRW97]. On-line [ALM90, AAF^{+}93, BC97, BGR93, LLW91, BFL96, CDRS90, KVV90]. once [BHH92, Pon95, RWY97]. One [BH98, BIKM99, Rom90, CMR98, DI99, FL92a, JL93, KNR95, MPS95, NS96, RMR^{+}93, TW98]. one-round [FL92a, KNR95]. One-way [BIKM99, Rom90, CMR98, DI99, JL93]. Online [AS97b, BGRS90, LY92, PW93, Alb97, BDBK^{+}90, DM91, FKST93, Gro91]. only [VY94]. operation [HK95, SS92]. optical [BFL96, RU94]. Optimal [BF99, CKLW99, FKST93, GKS90, Hag94, Har94, Pan95, PSTW97, VS90, AKM^{+}93, CR93, CRS93, CY94, Cle90, CDKT93, CGG^{+}95, EP95, FKW93, Gal92, GKR^{+}99, Hås97, HI95, KVV90, Mat91, TS98, Zuc96]. Optimality [FW94, ABAR99]. optimistic [AMO96]. optimization [DI96, cDF97, Epp95, KP95]. oracle [CGH98]. order [GHK92, Mul90]. ordering [BKRV98]. orientations [MRST97]. Orleans [ACM91]. other [AYZ94, BKRV98, Mul94, Zwi99b]. outage [GMNS99]. Output [Mul90, AZ98]. Outward [Zwi99b]. overcoming [Erg95, KK90].

P [IW97, Mul97, Sip92]. Packet [AZ99, AKOR98, Fei99, Gam99, HBW95, MPS95, OR97, RT96, ST97]. packet-routing [HBW95]. packing [CCG^{+}91, CJSW93, Gab91, LV92, Pul95, Sri95]. Page [Ira97]. pages [Ira97]. paging [BRIS91, FK95b, KL98b]. pairing [Fre98a]. pairs [Sei92, Zwi99a]. pairwise [DK99, Wig94]. Parallel [ACMR95, yC92, AGL98, AW91, BG91, CT91, CGG^{+}95, FKST93, Gal92, Goo92, Goo96, Hag91, Hag94, Har94, JP99, KK90, KPS90, KPRS91, KPRR92, Kel92, KS92, KR99, Lei92, LR97b, LN93, MNV94, MV91, Mul94, Pan95, Raz95, Rei95, SW99, VS90]. parameterized [Bak93]. Parametric [AHV93, AM92a, Epp95, McC96]. partial [FY95]. partially [Mau91]. parties [CO99]. partition [FHKM99, MTTW95]. partitions [FM91]. party [BR95, CFGN96, Kil91]. Paso [ACM97]. path [ALM90, BFU97, HS93b, KLMR95, KRRS94, Sei92]. paths [AKW96, AYZ94, BFU92, Coh94, GKR^{+}99, Har91, Kap99, KS92, KT95b, Kle98, Zwi99a, dBMS93]. pattern [Bak93, CH97, Kos94, Vis90]. patterns [CK99b, IS93]. Paul [Bab97]. PCP [DFK^{+}99, RS97]. PCPs [EKR99, Tre98]. pebble [BFR^{+}98]. average-case [AD97]. correcting [BLR90, GLR^{+}91]. space [FN91]. Pennsylvania [ACM96]. Perfect [BMO90a, Mau91, ABV95, CRS93, Fil97, RSZ99]. perfection [CL93]. perfection-information [CL93]. Perfectly [CMR98]. performance [AF92]. permanent [FL92b]. Permanents [MRST97, LSW98]. permutation [BCF^{+}91, MvzG94]. permutations [ACG93, BCFM98, HP95, NR97]. Persistent [KT95a]. Pfaffian [MRST97]. phenomena [Fei99, MPR94]. Philadelphia [ACM96]. pick [AAFL96]. Piecewise [dBMS93]. pigeonhole [BIK^{+}92, RWY97]. PL [Ogi96]. Planar [CGP98a, Goo92, DV93, EGIS93, FW93, He97, KK90, KRRS94, KT95b, Lin92a, Mal92, PP93, Rao92, RSST96, Sch99b]. planarity [GIS92, La 94]. plane [AKTT97, BCV95, CK93, GH94]. planes [BPR95, Fu96]. planning [ART95, KRR91, KLMR95, Pel94]. players [FW94]. point [CK92, FW93, GT91, Mul91, Pel94, Rei97, VY94]. point-sets [CK92]. Pointer [NBY97, BKRW98, La 90, PRV99]. pointer-machine [BKRW98]. points [AKTT97, KV97]. policies [Gam99]. Poly [HdLT98, AM92b]. Poly-logarithmic [HdLT98]. polygon [Goo92, Mil96]. polygons [KR99]. polyhedra [Sch99b]. polyhedron [GKV94]. Polylog [Coh94, BBBT97, Pan95, SSZ95]. Polylog-time [Coh94]. polylogarithmic [BFLS91, HK95]. polymorphic [BPWY96]. Polynomial [AB92, AKK95, Gol93, KM95, AKTT97, BR91, BT90, BGIP99, CNS99, CM99, GM93, HP95, JSOS99, Kar95, KvM99, KOR95, LV98, LSW98, MP98, NP99, NSS98, OW90, Pan95, Rei95, Rei97, Sch99a, Way99]. polynomial-time [KvM99]. polynomially [DFK^{+}99]. polynomially-small [DFK^{+}99]. polynomials [ABFR91, BBR92, GLR^{+}91, KS95, NS92, Pat92, Sha93, vzGS92]. polytime [BC92]. population [RRS95]. positive [LN93]. potential [CK92]. power [ABFR91, BDBK^{+}90, BFR^{+}98, CI95, CHPW94, FL92a, GMNS99, KP94c, Maa93, Wig94]. powerful [TW98]. PP [BRS91a]. Practical [LMS^{+}97, KST90]. practice [DGS98]. PRAM [KLadH92]. precise [BSS97]. predecessor [BF99]. prediction [KW95]. predictors [FSSW97]. preemption [BNCK^{+}95, CI95]. preemptive [CG91]. preliminary [AK94, ARV94, BU96, BFU97, CK92, CMR98, CGH98, CY94, CHPW94, DG96, DLN96, Fei96, GW94b, Goo92, Goo96, KPRR92, Kos94, La 94, MHSR94, Pat92]. presence [AD91, ADLS91, BPWY96, BK93]. preserving [DI99, IMRV97]. primal [WGMV93]. primal-dual [WGMV93]. primes [BGIP99]. primitive [Sho90]. principle [BIK^{+}92, RWY97]. principles [Ajt94]. privacy [FY95, GIKM98, KOR96]. Private [OS97, Bea96b, BIKM99, CG97, GR99, GIKM98, IK99, NS96]. prize [AABV95]. prize-collecting [AABV95]. Probabilistic [Kar91, ACY95, BGLR94, CNS99, CMR98, CHPW94, FS97, MPR94]. Probabilistically [CFLS93, KPT97, BGLR93]. probabilities [GHK92]. probability [CRS94, DFK^{+}99, MV91, NN90, RS97]. problem [AKR91, Ajt98, AW91, AL94b, BFKV92, BF99, BCC^{+}94, BCV95, BRV96, BN93, CGTS99, Die97, DR91, FMRT90, GH94, GHPT99, HR98, Kar95, KTU90, KT95b, Kle98, La 90, MS93b, NSS98, PRW97, Phi93, Sei92, Var97, Wan95]. problems [AHL90, Ajt96, AKK95, Aro98, ARR98, AK99, BFL96, BCE^{+}95, BC97, BLR90, BKRV98, BFH94, BOR99a, BOR99b, CRS93, CGG^{+}95, DM91, FHKM99, FL92a, FMdB99, GI91, GG98, GKR^{+}99, Ind99b, Kar94, KP95, Kar93, KSW97, KR95, KPR98, LSM^{+}91, LY93, MHSR94, Mil94, Mul94, PS92, Pul95, SV99c, STA97, Sri95, VR92, WGMV93, YA94, Zwi99b]. Proceedings [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97, ACM98, ACM99]. process [GJ97]. processing [KLMR95]. processor [AM92b, AHS91, CG91]. product [PRW97]. products [Fei95]. program [FGY96, Gál98, KPRR92, Tha98, Yao90]. programmed [BOW92]. programming [AB92, AM92b, LN93, Mil96, Mul94, PY93, Zwi99b]. programs [BGK^{+}96, Cle90, Gál98, KL96, Pon95, RWY97]. progress [BCPV93]. projections [vzGKS93]. Proof [CG91, DT92, FL92a, FK95a, Mul97]. proofs [BKPS98, BGLR93, BGLR94, BSW99, BPR95, CEI96, CJSW93, CD97, Fu96, Kil92, KPT97, KvM99, MP97, Oka96, PS94, RR94, RWY97]. properties [GW94b, PU99]. Property [GR97]. Proportionate [BCPV93]. Protecting [GIKM98]. protection [BGG95, DLN96]. protein [CGP^{+}98b, HI95]. protocol [CL93]. protocols [BMR90, BCK98, CMadH^{+}98, FS90, FK94, KLL^{+}97, MPR94, SV99a, Yao95]. Provably [BR95, NY90]. prover [FL92a, FK94, FK95a]. provers [BP92]. Pseudo [Hås90, ALS97, NR97, Tre99]. Pseudo-random [Hås90, ALS97, NR97, Tre99]. Pseudorandom [Nis90, Sit94, STV99].

Pseudorandomness [INW94, Bea96b]. PSPACE [ARV94, CFLS93, MHSR94]. PSPACE-complete [ARV94, MHSR94]. PSPACE-hard [CFLS93]. PTAS [KM96b, SW99]. Public [NY90, NS96, AD97, SZ93]. Public-key [NY90, AD97]. Purely [KT96]. Pushing [BS96].

QoS [PU99]. quadratic [ARV94, Bar92, Ko91]. quality [GW94b]. quality-size [GW94b]. quantifier [TW98]. Quantifiers [PR90]. Quantitative [KMY90]. Quantum [AKN98, Bea97a, BV93, BCW98, HH99, ABO97, AI99, ANTSV99, Gro96, Gro98, NW99, Raz99, SV99b, Yao95]. quasi [CNS99, Kos94]. quasi-polynomial [CNS99]. quasi-real-time [Kos94]. Quebec [ACM94]. queries [AK91a, AWZ96, BCV96, BR92, Cla97, CDKT93, HRPW95, Kea93, Tre98]. query [AHL90, BFJ^{+}94, KLMR95, NW99]. query-retrieval [AHL90]. question [Sip92]. queueing [BKR^{+}96, Gam99]. queues [MS92]. quickly [Ree92, Wil96]. quorum [MR97]. quotient [KL90, PP93].

Rackoff [NR97]. radical [Lak90]. RAM [SS92]. RAMs [Ajt99, Ost90]. Ramsey [Sch99a]. Random [CDRS90, Kar94, PL98, Sch99b, AK94, ALS97, ABK^{+}92, Bab91, BF93, BKPS98, BCL94, BOW92, BFU97, CGH98, FL92b, FK95a, FKR^{+}93, GW94b, Hås90, KLL^{+}97, Kar97, KM94b, Mil94, NR97, Odl94, TS96, Tre99, Wil96]. randomization [BDBK^{+}90, DM91, KPRR92]. Randomized [BL92, CMadH^{+}98, Fei95, FK95b, GK97, Gri98, HK95, KLMR95, ACMR95, Ajt98, Asp97, Ben94, CI95, CH97, GKS90, GKadHS96, Kal92, Kar95, KRR91, KS92, KT94, KMvS94, KNR95, KMRZ93, dBDS94]. Randomness [CRS93, Zuc96, CK97, RRV99, RR99, SS90a, TS96]. Randomness-optimal [CRS93, Zuc96]. range [AWZ96, Cha95, VV96]. rank [ABO96]. ranks [PR93]. Rao [KST90]. rapidly [GJ97]. rate [AMO96, ER90]. rates [FK94]. ratio [BD95, PT93]. rationality [PY94]. Ray [AM92a]. reach [ADLS91]. Read [RWY97, BHH92, NBY97, Pon95, Tha98]. read-k-times [Tha98]. Read-once [RWY97, BHH92, Pon95]. Real [Kos94, AHV93, BNGNS99, Bar92, BD92, CKK^{+}95, GM95, NS92]. Real-time [Kos94, AHV93, BNGNS99]. real-valued [BD92]. reals [ABCG93, FK98]. reasoning [AHV93]. Receipt [BT94]. Receipt-free [BT94]. Recognition [RS95]. recognizing [MvzG94]. reconfiguring [CMS93]. Reconstructing [BKL96]. recovery [GMNS99]. rectangles [ALS97, KR99, LLSZ93]. rectangular [IS93, RWY97]. rectilinear [KR99]. recurrence [Kar91]. recursion [BC92]. recursion-theoretic [BC92]. recursive [KT95a]. Recycling [Tre98, FK95a, RR99]. reducibility [OW90]. Reducing [AAI^{+}97, CK97, MVO91, RRV99]. reduction [AJ93]. reductions [AAI^{+}97, Ajt98]. reference [BRIS91, FK95b]. regions [CK93, LMW99]. regular [DI99, Spi96]. regularities [CK99b]. Reinventing [CDKT93]. related [ARR98, BOR99a, CRS93, CP90, GKR^{+}99, Mil94, NW99, YA94]. relations [Kar91]. relationships [Oka96]. relative [BCE^{+}95, KMS99]. relaxations [BKRV98, Zwi99b]. reliability [Kar95]. relieving [KLL^{+}97]. remaindering [GRS99]. removal [Mul91]. removing [IM98]. repeated [FW94]. repetition [Raz95]. replacement [CK99b, Ira97]. representations [KT96, ST90]. Representing [BBR92]. require [RW90]. requirements [AZ99]. requires [IW97, NBY97]. resilience [CR93]. resilient [BG93, HS93a, KPRR92, LMS^{+}97]. resolution [BSW99, MPR94, Mal92, RU95]. resolved [Die97]. resource [BCPV93, CS92, PSTW97]. respect [Mul91]. respecting [Fei98]. restriction [Mil96]. restrictions [CR96]. result [BG93]. resultants [KS96]. results [AZ98, BS94, FK95a, Fle93, GKR^{+}99, Hås97, KTW96, MS93a, PRW97]. Retraction [FS97]. retrieval [AHL90, BIKM99, CCFM97, CG97, GIKM98, IK99]. retrospective [Cha94]. returns [CKLW99]. reversals [HP95]. revisited [CGH98, NR97, NW91]. Riemann [BS91]. Rigorous [FN91]. rings [DMR99]. RL [Nis92]. RNC [Ben94]. roadmaps [BPR96]. Robot [AWZ96, ART95, KLMR95]. Robust [FMY98, Val99, FKW93, FGY96, KPS90, LH97]. roots [Sho90]. rotations [Zwi99b]. round [BMR90, FL92a, KNR95, NS96]. Rounding [CCGG98, KKS^{+}99, Zwi99b]. Rounds [NW91, BMO90a, GM93]. Routing [ACG93, AKOR98, AZ99, AAF^{+}93, AL94b, BFL96, BRSU93, BU96, CKR98b, CMadH^{+}98, CadHSV96, DKKP95, Fei99, Gam99, HBW95, Lei92, MS92, MV97, MPS95, Pip93, RU94, SV99a, ST97]. RSA [FMY98]. RSA-key [FMY98]. rules [LW95]. running [VY94].

salesmen [AABV95]. Sample [Sch92, KM93, Ser99]. samples [BD92]. Sampling [AK91b, KV97, Fil97, HH99, Hub98, Kar94, Kar97, PRV99, Sch99b, Vis90, Zuc96]. San [ACM93]. sandwich [MLS92]. sat [Fre98b, GW94a, BCV96]. Sat- [BCV96]. satisfaction [FV93, KSW97]. Satisfiability [Pla99]. satisfying [KM93, Zwi98]. Scalable [BCL94, SV99b]. scale [CMS93, KD95, KD96, SV99b]. scaling [LSW98]. Scheduling [Edm99, GHPT99, Alb97, AAF^{+}93, AKP92, BNGNS99, BFKV92, BL93, CI95, CM99, CG91, ECBD97, FS98, FKST93, JP99, KP94a, McC96, PSTW97]. scheme [AKTT97, CM99, Coh94, DKKP95, JSOS99, Kar95, KS92]. schemes [AKK95, ARR98, AK99, GIKM98, JP99, MHSR94]. Search [Odl94, AM92a, BCE^{+}95, BK93, BOR99a, FG95, Gro96, Kle97, KOR98, PSY90, ST90]. Searching [AD91, Sho90, AHHP94, AHP95, CCGL99, Cha95, HS93b, VV96]. second [ACM90]. secret [BT94]. secret-ballot [BT94]. Secure [CO99, FY95, BMR90, BR95, BOCG93, CFGN96, FKN94, FY92, NY90, Rom90]. securely [DDFY94]. Security [DI99, Yao95, Mau91]. Security-preserving [DI99]. seems [KL98a]. segment [AGR95]. Segmentation [KPR98]. segments [CKR^{+}99]. selected [Ang92]. selection [ALM90, BFU97, Fre90]. selective [AS97b]. Self [BLR90, GLR^{+}91, MOOY92, Pip93, AKM^{+}93, CMS93, DLN96]. self-enforcing [DLN96]. Self-routing [Pip93]. self-simulation [CMS93]. Self-stabilizing [MOOY92, AKM^{+}93]. Self-testing [BLR90, GLR^{+}91]. Self-testing/correcting [BLR90, GLR^{+}91]. Semi [BKRV98, Bas96, BPR96, cDF97, KTU90]. semi-algebraic [Bas96, BPR96]. Semi-definite [BKRV98]. semi-local [cDF97]. semi-unification [KTU90]. semialgebraic [Vor98]. semidefinite [KL96, Zwi99b]. sensitive [Mul90]. sensitivity [GR99]. Separating [Sha91, Tha98]. separation [Raz99]. Separator [EGIS93, AST90]. Separators [MT90, CY94, CY97, Goo92, Ree92]. sequences [JLW94, ST90]. sequential [Pan95]. server [BIKM99, Gro91, KRR91, KP94b]. session [BR95]. set [cDF97, Fei96, GKS90, GHPT99, HR98, Kel92, LLSZ93, SZ93, Sla96, Vor98]. set-maxima [GKS90]. sets [ALS97, Bas96, BPR96, CK92, CEG^{+}93, FW93, HR94, OW90, SS92, ST90]. seventh [ACM95]. Shallow [PZ92]. share [DDFY94]. shared [DHW93]. sharing [BGR93, FGY96]. shooting [AM92a, BOW92]. shop [FS98]. shops [JSOS99]. Short [BF93, BSW99, GY95, ADM^{+}95, BS99, FK97, RU95]. shortest [Ajt98, BJL^{+}91, Coh94, HS93b, Kap99, KS92, KRRS94, Sei92, Zwi99a]. shortest-path [KRRS94]. shredders [CT96]. sieve [LLMP90]. sigmoidal [KM95, MS93a]. signatures [Rom90]. signed [HP95]. signets [DLN96]. similar [LMW99]. Simple [DW92, LY94, MS92, ORV94, AYZ94, BSW99, HS94]. simplex [AGR95, Kal92]. simplices [Pel94]. simplicial [DG96]. simplified [HH99]. simply [CK93]. Simulating [ARV94, AL94a, GK93]. simulation [CMS93, KLadH92, NZ93]. simulations [Cle90]. single [AGL98, BIKM99, KTW96]. single-server [BIKM99]. singly [BAP99]. singular [Adl91]. singularities [BD92]. sixth [ACM94]. Size [IPS93, Gál98, GW94b, Gro94, Ira97, KW93, KvM99, KOR96, Ma96, PS94, PR93]. Size-depth [IPS93]. slow [KT95a]. slow-down [KT95a]. Small [CK99a, KP92, NN90, AYZ94, BS92, Bod93, BPR95, DFK^{+}99, KRY94, KM93, LLSZ93, Rao92]. Small-bias [NN90]. Small-depth [KP92]. smarter [LH97]. SNP [FV93]. solution [Pan97, Rei95]. solutions [BBF98, Zwi99b]. Solving [AHL90, MP98]. Some [CK93, Hås97, Kar93]. sorted [AHHP94, AHP95, KT96]. Sorting [AHNR95, AFGV97, CP90, GKS90, Goo96, Hag91, Hag94, HP95, KLM^{+}95, KK92, Ma96, MV97, Pla92]. sources [TS96]. Space [BL93, AMS96, BR91, Blu90, DG96, Edm93, EP95, Gol93, Ind99b, KD95, KD96, KOR95, MOOY92, Nis90, RR99]. space-bounded [Nis90]. Space-efficient [BL93, KD95, KD96]. spaces [BOR99b, Cla97, Gup99, IMRV97, KM93, KOR98, NN90, Sch92]. span [BGK^{+}96, Gál98]. spanners [ADM^{+}95, LNS98, RS98]. spanning [GH94, HdLT98, KRY94, KT94, Wil96]. sparse [BT90, CT91, HR94, OW90]. sparsification [EGIS93]. Sparsity [KS96]. specialize [FSSW97]. specific [Kha93]. specifications [MHSR94]. Spectral [LR97a, AK94]. spectrum [BT95, KM91]. speed [CLP99, DKKP95]. speeding [FM91]. speeding-up [FM91]. splay [Col90]. split [La 90, Mil94]. split-find [La 90]. splitting [Gab94, NI96]. Spot [EKK^{+}98]. Spot-checkers [EKK^{+}98]. spots [KLL^{+}97]. st [Edm93]. st-connectivity [Edm93]. Stability [AZ98, Gam99]. stabilizing [AKM^{+}93, MOOY92]. stall [AGL98]. standard [MP94]. state [YL91]. states [AKN98, CHPW94, RR99]. Static [BFU97, PU99, SV99a]. statistical [ABV95, BMO90b, BFJ^{+}94, GSV98, Kea93, Oka96]. statistics [NW99]. status [Sip92]. steering [ART95]. steering-constrained [ART95]. Steiner [AKR91, BC97, BD95, CCGG98, Pul95, WGMV93]. Steinitz [KMY90]. Stochastic [RU95]. stone [RMR^{+}93]. stopping [LW95]. storage [OS97]. store [CadHSV96, SV99a]. store-and-forward [CadHSV96, SV99a]. strategies [CKLW99, KMvS94, LY94]. String [FT95, BG91, CGG^{+}95, FMdB99, Gal92, Hag94, JL93]. string-matching [Gal92, JL93]. stringology [MP94]. strings [AHHP94, AFGV97, FT95, KD95, KD96, LMW99]. strong [Saf92]. strong-fairness [Saf92]. strongly [LSW98, Spi96]. structural [Luk99]. Structure [BS92, CDKT93, FW93, FG95]. structures [Mic97, MNSW95, Rau94]. study [CCG^{+}91]. sub [RS97]. sub-constant [RS97]. subexponential [Kal92, KvM99]. subgraphs [AYZ94, DR91]. Sublinear [Ind99b, BR91, GR98]. Subquadratic [BOR99b, KS95]. Subquadratic-time [KS95]. subset [CH97, DV94]. substring [FG95]. succinct [MHSR94]. sufficient [Rom90]. suffix [Har94, Kos94, KD95, KD96, SV94]. sum [LY94, SW99]. superconcentrators [Pip93]. superpolynomial [BGK^{+}96]. superstrings [BJL^{+}91]. surface [Moh96, Mul91]. survey [Ang92]. Swendsen [GJ97]. switching [OR97, RT96]. Symmetric [NTS95, Bea97a, Ind99a, Pat92, Sze90]. Symmetry [BBTN92, SV94, ACS94, DMR99, MOOY92]. Symposium [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97, ACM98, ACM99]. synchronization [AAT94, AKM^{+}93, PSR94]. synchronizer [ER90]. syntactic [KM96b]. system [DN95]. systems [AGL98, ABO96, ARV94, BBBT97, Bar92, CS92, CFLS93, DW92, FL92a, FK95a, Ko91, LY92, MR97, MP98, Rei95].

table [OW90]. tapes [JSV94]. Target [BOW92]. targeting [CKR^{+}99]. task [BBBT97]. tasks [HS93a, HR97]. TCP [DGS98]. team [AFS94]. technique [AK94, CMS93, Vis90, Zuc90]. techniques [Hub98, KLP^{+}97, LR97a]. tensors [PR93]. tentative [KPRS91]. term [BR92]. terminal [Kar95]. terms [KOR96]. terrain [BRS91b]. test [RS97]. tester [GR98]. Testing [Erg95, Hås96, YL91, AS97a, Bar92, GIS92, GR97, GKV94, KMS99, La 94, Spi96, ST90]. testing/correcting [BLR90, GLR^{+}91]. tests [ACY95, Tre98]. Texas [ACM97, ACM98]. their [FL92a, GVY93, vzGKS93]. theorem [AST90, BDBK97, Bsh98, GR99, GY95, HS93a, HS94, Kil91, RS94, Raz95]. theorems [CNS99, GVY93, KMY90]. theoretic [BC92, CBDFS96, FW90, Fre98a, IK99]. Theory [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97, ACM98, ACM99, Ang92, Bab97, Bak93, BV93, BKR^{+}96, DGS98, Fre90, GM95, LY94, PSR94, Sch99a, Var97]. there [Mul97]. Thermodynamics [BGL^{+}93]. thin [ADM^{+}95]. third [ACM91]. thirtieth [ACM98]. thirty [ACM99]. thirty-first [ACM99]. three [BR95, BKL96, HI95, MS93b, MT90, MTTW95]. three-dimensional [BKL96, MS93b]. three-eights [HI95]. threshold [Fei96, GK93, IPS93, KP94c, RS95]. throughput [BNGNS99, MNV94]. Tight [GLM^{+}95, AHP95, Fre90, Sla96]. Time [AAT94, AKM^{+}93, Edm93, YA94, AM92b, Ajt99, AGL98, AHV93, AHNR95, AKK95, AKTT97, ADLS91, AAFL96, AALR99, BFLS91, BNGNS99, BR91, BS92, Ben94, BK96, Bod93, BKRW98, CNS99, CM99, Coh94, CH97, CP90, CGG^{+}95, EP95, FN91, FS97, Gal92, GadHW90, Hag91, HK95, Ind99b, JSOS99, Jon93, KS95, Kar95, Kar96a, KTW96, KT94, KvM99, Ko91, Kos94, LR97b, MNV94, MV91, Moh96, NI96, OW90, PSTW97, Spi95, VY94, Wil96, Zuc90]. Time-adaptive [AAT94]. time-critical [PSTW97]. Time-space [Edm93, EP95]. time/space [FN91]. times [SW99, Tha98]. timestamp [DW92]. timestamping [DW92]. timing [ADLS91]. Tiny [GW94b]. Toeplitz [Rei95]. together [BS96]. tolerance [KNT94]. tolerant [ABO97, BGM^{+}96, CBDFS96, CS92, KP94a, Kea93, LNS98, Lin92a, Ma96]. tool [Zwi99b]. top [KM96a]. top-down [KM96a]. topological [AFS94, BLY92]. topology [SZ93]. toss [CKK^{+}95]. total [KTW96, LR97b]. tours [AKTT97]. Trade [MNV94, AKW96, Edm93, GW94b, Gro94, IPS93]. trade-off [AKW96, GW94b, Gro94]. Trade-offs [MNV94, Edm93, IPS93]. tradeoffs [FN91]. traffic [AKOR98, CK99b, RS93]. transfer [NP99, VS90]. transfers [GHPT99]. transformations [KPRR92, SV99a]. Transforming [HP95]. transforms [Bea97a, BSS97]. transition [LY92]. transitive [Bab91, KK90, KS99]. transitive-closure [KK90]. Translational [Mil96]. transversal [RWY97]. Tree [CH97, Bar98, BC97, Bod93, GH94, Gup99, Har94, HdLT98, JLW94, KM96a, KV98, Lin92b, Pul95, Ree92, SV94, ST90, Yao94]. tree-decompositions [Bod93]. Trees [LMS98, AKR91, AABV95, AS97b, BLY92, BPWY96, CCGG98, Col90, EP96, FKW93, Fle93, GT91, GKV94, GKadHS96, KLP^{+}97, KLL^{+}97, KRY94, KT94, KMvS94, Kos94, KD95, KD96, KM91, Wil96, FW90]. treewidth [Bod93]. Trevisan [RRV99]. triangular [DV93]. triangulation [Goo92]. true [BMO90b]. truly [CO99]. truth [OW90]. truth-table [OW90]. TSP [Tre97]. Turing [CKK^{+}95]. turnip [HP95]. twenty [ACM90, ACM91, ACM92, ACM93, ACM94, ACM95, ACM96, ACM97]. twenty-eighth [ACM96]. twenty-fifth [ACM93]. twenty-fourth [ACM92]. twenty-ninth [ACM97]. twenty-second [ACM90]. twenty-seventh [ACM95]. twenty-sixth [ACM94]. Two [FL92a, FK94, JSV94, Kle97, Pul95, ABF92, CK93, CG91, CM91, FK95a, GLM^{+}95, Kil91, MT90, RS95, TW98]. two-dimensional [CK93]. two-processor [CG91]. Two-prover [FL92a, FK95a]. type [Ind99a]. typical [FKR^{+}93].

uncapacitated [Kar97]. uncertainty [ADLS91, AAFL96]. Undecidability [AI99, Fre98b, KTU90]. undirected [BR91, Ben94, Coh94, Edm93, Kar97, KL98a, NI96]. unfamiliar [BRS91b]. unification [KTU90]. uniform [Hås90, KR95, KST90, Sch92]. union [ABAR99, AW91, La 90, Mil94]. union-find [ABAR99, AW91, La 90]. union-split-find [Mil94]. Unique [GKT99, ST90, BBF98, CRS93]. uniqueness [PL98]. univariate [BT90]. Universal [CadHSV96, OR97, CK99a, MNT90]. unknown [AH97]. unless [KvM99]. unrelated [JP99]. unreliable [FPRU90]. unrooted [KLP^{+}97]. unsatisfiability [BKPS98, CEI96]. unsplittable [Kle98]. updates [KW95]. upper [IK99]. use [CBFH^{+}93, ER90]. Using [CEI96, FSSW97, Kar97, Adl91, BR91, BFJ^{+}94, BSS97, KM91, LMSS98, Mil96, Tre99].

valued [BD92]. variables

[BOW92, CM91, KM94b, TW98]. various [SS92]. VC [KM95]. vector [Ajt98]. vectors [BS99]. Vegas [ACM95]. vehicle [CKR98b]. verification [BKRW98, Kur94]. verifier [GSV98]. version [AK94, ARV94, BU96, BFU97, CK92, CMR98, CGH98, CY94, CHPW94, DG96, DLN96, Fei96, GW94b, Goo92, Goo96, KPRR92, Kos94, La 94, MHSR94, Pat92]. versions [GY95]. versus [Ajt99, BL92, Dam98, KW95, Sip92]. vertex [Bab91, BKRV98, CT91, DV94]. vertex-ordering [BKRV98]. vertex-transitive [Bab91]. very [KPRS91, SS90b]. via [ACG93, CRS94, CCGG98, CK97, Fei98, Fil97, Fri96, JLW94, KT95a, KPRR92, LK99, PSTW97, RS98]. Victoria [ACM92]. view [Mul91, RRS95]. violation [LV92]. virtual [AAF^{+}93]. virus [BGG95]. VLSI [AKW96]. volume [BLY92, Fei98]. Voronoi [AHL90, Mul90]. voting [ABFR91]. vs [BCW98, CG91, NS96, ST97].

Wait [ADT95, AW91, SZ93, HS94, LH97]. Wait-free [ADT95, AW91, SZ93, HS94, LH97]. wakeup [FMRT90]. walk [BFU97, Odl94]. walks [ABK^{+}92, BF93, CDRS90, FKR^{+}93].

Wang [GJ97]. way [ANTSV99, BIKM99, CMR98, DI99, JL93, Rom90]. Weak [KV98, CEG^{+}93, TS96]. Weakly [BFJ^{+}94]. Web [CK99b, Ira97, KLL^{+}97]. weight [AABV95, Gro94, KRY94]. weight-size [Gro94]. weighted [CDRS90, SW99]. wheel [CDKT93]. which [Sha93]. while [AL94a]. whose [VY94]. Wide [KLL^{+}97]. width [BFH94, Cle90, Ree92]. Williamson [Kar96b]. winner [AAFL96]. wiring [BCL94]. wise [BCFM98, KM94b, PZ92]. within [AYZ94, HI95]. without [AALR99, PY93, STV99]. Witness [FS90, FGY96]. Witness-based [FGY96]. won’t [AK91a]. word [Pla99, Wan95]. words [TW98]. Work [CGG^{+}95, Rei95, Coh94, Rei97]. Work-time-optimal [CGG^{+}95]. World [KLL^{+}97]. wormhole [CadHSV96]. Worst [ABAR99, AD97, JRS94]. Worst-case [ABAR99, AD97]. worst-case/ average-case [AD97].

XOR [IW97, STV99].

yields [ER90].

zero [Bea96a, BMO90a, BMO90b, BP92, CD97, DNS98, GSV98, Kil92, KPT97, Lak90, LY94, Oka96]. zero-knowledge [BMO90a, BP92, CD97, DNS98, GSV98, Kil92, Oka96]. zero-sum [LY94]. zeros [Pan95]. Ziv [FT95].

References

Awerbuch:1995:IAG

[AABV95] Baruch Awerbuch, Yossi Azar, Avrim Blum, and Santosh Vempala. Improved approximation guarantees for minimum-weight k-trees and prize-collecting salesmen. In ACM [ACM95], pages 277– 283. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p277-awerbuch/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p277-awerbuch/p277-
awerbuch.pdf. ACM order no.

508950.

Aspnes:1993:LLB

[AAF^{+}93] James Aspnes, Yossi Azar, Amos Fiat, Serge Plotkin, and Orli Waarts. On-line load balancing with applications to machine scheduling and virtual circuit routing. In ACM [ACM93], pages 623–631. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Awerbuch:1996:MCF

[AAFL96] Baruch Awerbuch, Yossi Azar, Amos Fiat, and Tom Leighton. Making commitments in the face of uncertainty: how to pick a winner almost every time (extended abstract). In ACM [ACM96], pages 519– 530. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p519-awerbuch/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p519-awerbuch/p519-
awerbuch.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Agrawal:1997:RCR

[AAI^{+}97] Manindra Agrawal, Eric Allender, Russell Impagliazzo, Toniann Pitassi, and Steven Rudich. Reducing the complexity of reductions. In ACM [ACM97], pages 730– 738. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p730-agrawal/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p730-agrawal/p730-

agrawal.pdf. ACM order no. 508970.

Awerbuch:1999:MFT

[AALR99] Baruch Awerbuch, Yossi Azar, Stefano Leonardi, and Oded Regev. Minimizing the flow time without migration. In ACM [ACM99], pages 198– 205. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p198-awerbuch/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p198-awerbuch/p198-
awerbuch.pdf. ACM order

number 508990.

Aiello:1993:ALB

[AAMR93] William Aiello, Baruch Awerbuch, Bruce Maggs, and Satish Rao. Approximate load balancing on dynamic and asynchronous networks. In ACM [ACM93], pages 632–641. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Alur:1994:TAA

[AAT94] Rajeev Alur, Hagit Attiya, and Gadi Taubenfeld. Timeadaptive algorithms for synchronization. In ACM [ACM94], pages 800–809. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p800-alur/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p800-
alur/p800-alur.pdf. ACM or-

der no. 508930.

Adler:1992:PAL

[AB92] Ilan Adler and Peter A. Beling. Polynomial algorithms for linear programming over the algebraic numbers. In ACM [ACM92], pages 483–494. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p483-adler/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p483-
adler/p483-adler.pdf. ACM

order no. 508920.

Alstrup:1999:WCA

[ABAR99] Stephen Alstrup, Amir M. Ben-Amram, and Theis Rauhe. Worst-case and amortised optimality in union-find (extended abstract). In ACM [ACM99], pages 499–506. ISBN 1-58113-067-8. LCCN QA75.5 .A14

1999. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/301250/p499-alstrup/;

http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p499-alstrup/p499-

alstrup.pdf. ACM order number 508990.

Ar:1993:CAC

[ABCG93] S. Ar, M. Blum, B. Codenotti, and P. Gemmell. Checking approximate computations over the reals. In ACM [ACM93], pages 786–795. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Amir:1992:AIT

[ABF92] Amihood Amir, Gary Benson, and Martin Farach. Alphabet independent two dimensional matching. In ACM [ACM92], pages 59–68. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p59-amir/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p59-
amir/p59-amir.pdf. ACM or-

der no. 508920.

Awerbuch:1993:CDF

[ABF93] Baruch Awerbuch, Yair Bartal, and Amos Fiat. Competitive distributed file allocation. In ACM [ACM93], pages 164–173. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p164-awerbuch/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p164-awerbuch/p164-

awerbuch.pdf. ACM order no.

508930.

Aspnes:1991:EPV

[ABFR91] James Aspnes, Richard Beigel, Merrick Furst, and Steven Rudich. The expressive power of voting polynomials. In ACM [ACM91], pages 402–

409. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p402-aspnes/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p402-aspnes/p402-aspnes.

pdf. ACM order no. 508910.

Azar:1992:BRW

[ABK^{+}92] Yossi Azar, Andrei Z. Broder, Anna R. Karlin, Nathan Linial, and Steven Phillips. Biased random walks. In ACM [ACM92], pages 1–9. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p1-azar/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p1-
azar/p1-azar.pdf. ACM order

no. 508920.

Azar:1994:BAE

[ABKU94] Yossi Azar, Andrei Z. Broder, Anna R. Karlin, and Eli Upfal. Balanced allocations (extended abstract). In ACM [ACM94], pages 593–602. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p593-azar/;

http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p593-
azar/p593-azar.pdf. ACM or-

der no. 508930.

Aumann:1992:CFA

[ABO92] Yonatan Aumann and Michael Ben-Or. Computing with faulty arrays. In ACM [ACM92], pages 162–169. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p162-aumann/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p162-aumann/p162-aumann.

pdf. ACM order no. 508920.

Allender:1996:CMR

[ABO96] Eric Allender, Robert Beals, and Mitsunori Ogihara. The complexity of matrix rank and feasible systems of linear equations (extended abstract). In ACM [ACM96], pages 161–

167. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p161-allender/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p161-allender/p161-
allender.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Aharonov:1997:FTQ

[ABO97] D. Aharonov and M. Ben-Or. Fault-tolerant quantum computation with constant error. In ACM [ACM97], pages 176–188. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p176-aharonov/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p176-aharonov/p176-
                                        aharonov.pdf. ACM order no.

508970.

Aiello:1995:KAP

[ABV95] William Aiello, Mihir Bellare, and Ramarathnam Venkatesan. Knowledge on the average — perfect, statistical and logarithmic. In ACM [ACM95], pages 469–478. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p469-aiello/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p469-aiello/p469-aiello.

pdf. ACM order no. 508950.

Alon:1993:RPG

[ACG93] N. Alon, F. R. K. Chung, and R. L. Graham. Routing permutations on graphs via matchings. In ACM [ACM93], pages 583–591. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

ACM:1990:PTS

[ACM90] ACM, editor. Proceedings of the twenty-second annual ACM Symposium on Theory of Computing, Baltimore, Maryland, May 14–16, 1990. ACM Press, New York, NY, USA, 1990. ISBN 0-89791-361-2. LCCN QA76.A15 1990. ACM order no. 508900.

ACM:1991:PTT

[ACM91] ACM, editor. Proceedings of the twenty third annual ACM Symposium on Theory of Computing, New Orleans, Louisiana, May 6–8, 1991. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991. ACM order no. 508910.

ACM:1992:PTF

[ACM92]  ACM, editor.  Proceedings

of the twenty-fourth annual ACM Symposium on Theory of Computing, Victoria, British Columbia, Canada, May 4–6, 1992. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-511-9. LCCN QA76.A15 1992. ACM order no. 508920.

ACM:1993:PTF

[ACM93] ACM, editor. Proceedings of the twenty-fifth annual ACM Symposium on the Theory of Computing, San Diego, California, May 16–18, 1993. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

ACM:1994:PTS

[ACM94] ACM, editor. Proceedings of the twenty-sixth annual ACM Symposium on the Theory of Computing: Montreal, Quebec, Canada, May 23–25, 1994. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-663-8. LCCN QA76 .A15 1994. ACM order no. 508930.

ACM:1995:PTS

[ACM95] ACM, editor. Proceedings of the twenty-seventh annual ACM Symposium on Theory of Computing: Las Vegas, Nevada, May 29–June 1, 1995. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995. ACM order no. 508950.

ACM:1996:PTE

[ACM96]  ACM, editor.  Proceedings

of the twenty-eighth annual ACM Symposium on the Theory of Computing, Philadelphia, Pennsylvania, May 22– 24, 1996. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

ACM:1997:PTN

[ACM97] ACM, editor. Proceedings of the twenty-ninth annual ACM Symposium on the Theory of Computing: El Paso, Texas, May 4– 6, 1997. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997. ACM order no. 508970.

ACM:1998:PTA

[ACM98] ACM, editor. Proceedings of the thirtieth annual ACM Symposium on Theory of Computing: Dallas, Texas, May 23– 26, 1998. ACM Press, New York, NY, USA, 1998. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998. ACM order number 508980.

ACM:1999:PTF

[ACM99] ACM, editor. Proceedings of the thirty-first annual ACM Symposium on Theory of Computing: Atlanta, Georgia, May 1–4, 1999. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. ACM order number 508990.

Adler:1995:PRL

[ACMR95] Micah Adler, Soumen Chakrabarti, Michael Mitzenmacher, and Lars Rasmussen. Parallel randomized load balancing. In ACM [ACM95], pages 238–

247. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p238-adler/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p238-
adler/p238-adler.pdf. ACM

order no. 508950.

Awerbuch:1994:EAD

[ACS94] Baruch Awerbuch, Lenore J.

Cowen, and Mark A. Smith. Efficient asynchronous distributed symmetry breaking. In ACM [ACM94], pages 214–223. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p214-awerbuch/

                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p214-awerbuch/p214-
                                        awerbuch.pdf. ACM order no.

508930.

Alur:1995:DTN

[ACY95] Rajeev Alur, Costas Courcoubetis, and Mihalis Yannakakis. Distinguishing tests for nondeterministic and probabilistic machines. In ACM [ACM95], pages 363–372. ISBN 0-89791-718-9. LCCN QA 76.6 A13

                                        1995. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/225058/p363-alur/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p363-
                                        alur/p363-alur.pdf. ACM or-

der no. 508950.

Aslam:1991:SPL

[AD91] Javed A. Aslam and Aditi Dhagat. Searching in the presence of linearly bounded errors. In ACM [ACM91], pages 486–493. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p486-aslam/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p486-
                                        aslam/p486-aslam.pdf. ACM

order no. 508910.

Ajtai:1997:PKC

[AD97] Miklós Ajtai and Cynthia Dwork. A public-key cryptosystem with worst-case/ average-case equivalence. In ACM [ACM97], pages 284–

293. ISBN 0-89791-888-6.

LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p284-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p284-
ajtai/p284-ajtai.pdf. ACM

order no. 508970.

Adleman:1991:FNU

[Adl91] Leonard M. Adleman. Factoring numbers using singular integers. In ACM [ACM91], pages 64–71. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p64-adleman/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p64-adleman/p64-adleman.

pdf. ACM order no. 508910.

Attiya:1991:BTR

[ADLS91] Hagit Attiya, Cynthia Dwork, Nancy Lynch, and Larry Stockmeyer. Bounds on the time to reach agreement in the presence of timing uncertainty. In ACM [ACM91], pages 359–

369. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p359-attiya/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p359-attiya/p359-attiya.

pdf. ACM order no. 508910.

Arya:1995:ESS

[ADM^{+}95] Sunil Arya, Gautam Das, David M. Mount, Jeffrey S. Salowe, and Michiel Smid. Euclidean spanners: short, thin, and lanky. In ACM [ACM95], pages 489–498. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p489-arya/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p489-
arya/p489-arya.pdf. ACM or-

der no. 508950.

Alon:1997:LHG

[ADM^{+}97] Noga Alon, Martin Dietzfelbinger, Peter Bro Miltersen, Erez Petrank, and Gábor Tardos. Is linear hashing good? In ACM [ACM97], pages 465–

474. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p465-alon/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p465-
alon/p465-alon.pdf. ACM or-

der no. 508970.

Afek:1995:WFM

[ADT95] Yehuda Afek, Dalia Dauber, and Dan Touitou. Wait-free made fast. In ACM [ACM95], pages 538–547. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p538-afek/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p538-
afek/p538-afek.pdf. ACM or-

der no. 508950.

Attiya:1992:CCH

[AF92] Hagit Attiya and Roy Friedman. A correctness condition for high-performance multiprocessors (extended abstract). In ACM [ACM92], pages 679–690. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p679-attiya/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p679-attiya/p679-attiya.

pdf. ACM order no. 508920.

Arge:1997:SSE

[AFGV97] Lars Arge, Paolo Ferragina, Roberto Grossi, and Jeffrey Scott Vitter. On sorting strings in external memory (extended abstract). In ACM [ACM97], pages 540–

548. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p540-arge/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p540-
arge/p540-arge.pdf. ACM or-

der no. 508970.

Apsitis:1994:CLT

[AFS94] Kalvis Apsītis, R. Ūsiņš Freivalds, and Carl H. Smith. Choosing a learning team: a topological approach. In ACM [ACM94], pages 283–289. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/195058/p283-apsitis/

                                        p283-apsitis.pdf;   http:/
                                        /www.acm.org/pubs/citations/
                                        proceedings/stoc/195058/p283-

apstis/. ACM order no.

508930.

Ajtai:1998:CMN

[AFS98] Miklós Ajtai, Ronald Fagin, and Larry Stockmeyer. The closure of Monadic NP (extended abstract). In ACM [ACM98], pages 309–

318. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p309-ajtai/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p309-
                                        ajtai/p309-ajtai.pdf. ACM

order number 508980.

Albers:1998:MST

[AGL98] Susanne Albers, Naveen Garg, and Stefano Leonardi. Minimizing stall time in single and parallel disk systems. In ACM [ACM98], pages 454–

462. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p454-albers/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p454-albers/p454-albers.

pdf. ACM order number

508980.

Amato:1995:CFS

[AGR95] Nancy M. Amato, Michael T. Goodrich, and Edgar A. Ramos. Computing faces in segment and simplex arrangements. In ACM [ACM95], pages 672–

682. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p672-amato/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p672-
amato/p672-amato.pdf. ACM

order no. 508950.

Albers:1997:EUE

[AH97] Susanne Albers and Monika R.

Henzinger. Exploring unknown environments. In ACM [ACM97], pages 416–

425. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p416-albers/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p416-albers/p416-albers.

pdf. ACM order no. 508970.

Andersson:1994:CSS

[AHHP94] Arne Andersson, Torben Hagerup, Johan Håstad, and Ola Petersson. The complexity of searching a sorted array of strings. In ACM [ACM94], pages 317–325. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p317-andersson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p317-andersson/p317-
andersson.pdf. ACM order no.

508930.

Aggarwal:1990:SQR

[AHL90] A. Aggarwal, M. Hansen, and T. Leighton. Solving queryretrieval problems by compacting Voronoi diagrams. In ACM [ACM90], pages 331– 340. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p331-aggarwal/

. ACM order no. 508900.

Andersson:1995:SLT

[AHNR95] Arne Andersson, Torben Hagerup, Stefan Nilsson, and Rajeev Raman. Sorting in linear time? In ACM [ACM95], pages 427–436. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p427-andersson/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p427-andersson/p427-
                                        andersson.pdf. ACM order no.

508950.

Andersson:1995:TLB

[AHP95] Arne Andersson, Johan Håstad, and Ola Petersson. A tight lower bound for searching a sorted array. In ACM [ACM95], pages 417– 426. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p417-andersson/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p417-andersson/p417-

andersson.pdf. ACM order no.

508950.

Aspnes:1991:CNM

[AHS91] James Aspnes, Maurice Herlihy, and Nir Shavit. Counting networks and multi-processor coordination. In ACM [ACM91], pages 348–358. ISBN 0-89791-397-3. LCCN QA 76.6 A13

1991. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/103418/p348-aspnes/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p348-aspnes/p348-aspnes.

pdf. ACM order no. 508910.

Alur:1993:PRT

[AHV93] Rajeev Alur, Thomas A. Henzinger, and Moshe Y. Vardi. Parametric real-time reasoning. In ACM [ACM93], pages 592–601. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Amano:1999:UQF

[AI99] Masami Amano and Kazuo Iwama. Undecidability on quantum finite automata. In ACM [ACM99], pages 368–

375. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p368-amano/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p368-
amano/p368-amano.pdf. ACM

order number 508990.

Allender:1993:DRN

[AJ93] Eric Allender and Jia Jiao. Depth reduction for noncommutative arithmetic circuits. In ACM [ACM93], pages 515–522. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Ajtai:1994:IMC

[Ajt94] Miklos Ajtai. The independence of the modulo p counting principles. In ACM [ACM94], pages 402–411. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p402-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p402-
ajtai/p402-ajtai.pdf. ACM

order no. 508930.

Ajtai:1996:GHI

[Ajt96] M. Ajtai. Generating hard instances of lattice problems (extended abstract). In ACM [ACM96], pages 99–

108. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p99-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p99-
ajtai/p99-ajtai.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Ajtai:1998:SVP

[Ajt98]  Miklós Ajtai.  The short-

est vector problem in L_{2} is NP-hard for randomized reductions (extended abstract). In ACM [ACM98], pages 10–

19. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p10-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p10-
ajtai/p10-ajtai.pdf. ACM

order number 508980.

Ajtai:1999:DVN

[Ajt99] Miklós Ajtai. Determinism versus non-determinism for linear time RAMs (extended abstract). In ACM [ACM99], pages 632–641. ISBN 1-58113-067-8. LCCN QA75.5 .A14

1999. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/301250/p632-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p632-
ajtai/p632-ajtai.pdf. ACM

order number 508990.

Angluin:1991:WWM

[AK91a] Dana Angluin and Michael Kharitonov. When won’t membership queries help? In ACM [ACM91], pages 444–

454. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p444-angluin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p444-angluin/p444-

angluin.pdf. ACM order no. 508910.

Applegate:1991:SIN

[AK91b] David Applegate and Ravi Kannan. Sampling and integration of near log-concave functions. In ACM [ACM91], pages 156–163. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p156-applegate/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p156-applegate/p156-
                                        applegate.pdf. ACM order no.

508910.

Alon:1994:STC

[AK94] Noga Alon and Nabil Kahale. A spectral technique for coloring random 3-colorable graphs (preliminary version). In ACM [ACM94], pages 346–355. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p346-alon/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p346-
                                        alon/p346-alon.pdf. ACM or-

der no. 508930.

Arora:1999:ASM

[AK99] Sanjeev Arora and George Karakostas. Approximation schemes for minimum latency problems. In ACM [ACM99], pages 688–693. ISBN 1-58113-067-8. LCCN QA75.5 .A14

                                        1999. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/301250/p688-arora/;
                                        http://www.acm.org/pubs/articles/

proceedings/stoc/301250/p688-
arora/p688-arora.pdf. ACM

order number 508990.

Arora:1995:PTA

[AKK95] Sanjeev Arora, David Karger, and Marek Karpinski. Polynomial time approximation schemes for dense instances of NP-hard problems. In ACM [ACM95], pages 284–

293. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p284-arora/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p284-
arora/p284-arora.pdf. ACM

order no. 508950.

Awerbuch:1993:TOS

[AKM^{+}93] Baruch Awerbuch, Shay Kutten, Yishay Mansour, Boaz Patt-Shamir, and George Varghese. Time optimal selfstabilizing synchronization. In ACM [ACM93], pages 652–661. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Aharonov:1998:QCM

[AKN98] Dorit Aharonov, Alexei Kitaev, and Noam Nisan. Quantum circuits with mixed states. In ACM [ACM98], pages 20–

30. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p20-aharonov/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        276698/p20-aharonov/p20-aharonov.

pdf. ACM order number

508980.

Aiello:1998:APR

[AKOR98] William Aiello, Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Adaptive packet routing for bursty adversarial traffic. In ACM [ACM98], pages 359–368. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p359-aiello/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p359-aiello/p359-aiello.

pdf. ACM order number

508980.

Awerbuch:1992:CDJ

[AKP92] Baruch Awerbuch, Shay Kutten, and David Peleg. Competitive distributed job scheduling (extended abstract). In ACM [ACM92], pages 571–

580. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p571-awerbuch/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p571-awerbuch/p571-
                                        awerbuch.pdf. ACM order no.

508920.

Agrawal:1991:WTC

[AKR91] Ajit Agrawal, Philip Klein, and R. Ravi. When trees collide: an approximation algorithm for the generalized Steiner problem on networks.

In ACM [ACM91], pages 134–

144. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p134-agrawal/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p134-agrawal/p134-

agrawal.pdf. ACM order no. 508910.

Asano:1997:CPP

[AKTT97] Tetsuo Asano, Naoki Katoh, Hisao Tamaki, and Takeshi Tokuyama. Covering points in the plane by k-tours: towards a polynomial time approximation scheme for general k. In ACM [ACM97], pages 275–283. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p275-asano/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p275-
asano/p275-asano.pdf. ACM

order no. 508970.

Aggarwal:1996:NDP

[AKW96] Alok Aggarwal, Jon Kleinberg, and David P. Williamson. Nodedisjoint paths on the mesh and a new trade-off in VLSI layout. In ACM [ACM96], pages 585–594. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p585-aggarwal/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p585-aggarwal/p585-

                                        aggarwal.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Auer:1994:SAH

[AL94a] Peter Auer and Philip M. Long. Simulating access to hidden information while learning. In ACM [ACM94], pages 263–272. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p263-auer/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p263-
                                        auer/p263-auer.pdf. ACM or-

der no. 508930.

Awerbuch:1994:IAA

[AL94b] Baruch Awerbuch and Tom Leighton. Improved approximation algorithms for the multi-commodity flow problem and local competitive routing in dynamic networks. In ACM [ACM94], pages 487–

496. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p487-awerbuch/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p487-awerbuch/p487-
                                        awerbuch.pdf. ACM order no.

508930.

Albers:1997:BBO

[Alb97] Susanne Albers. Better bounds for online scheduling. In ACM [ACM97], pages 130–

139. ISBN 0-89791-888-6.

LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p130-albers/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p130-albers/p130-albers.

pdf. ACM order no. 508970.

Arora:1990:LAP

[ALM90] S. Arora, T. Leighton, and B. Maggs. On-line algorithms for path selection in a nonblocking network. In ACM [ACM90], pages 149–158. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p149-arora/. ACM

order no. 508900.

Andrews:1996:AMH

[ALMZ96] Matthew Andrews, Tom Leighton, P. Takis Metaxas, and Lisa Zhang. Automatic methods for hiding latency in high bandwidth networks (extended abstract). In ACM [ACM96], pages 257–265. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p257-andrews/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p257-andrews/p257-

andrews.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Auer:1997:AHR

[ALS97] Peter Auer, Philip M. Long, and Aravind Srinivasan. Approximating hyper-rectangles: learning and pseudo-random sets. In ACM [ACM97], pages 314–323. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p314-auer/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p314-
auer/p314-auer.pdf. ACM or-

der no. 508970.

Agarwal:1992:RSP

[AM92a] Pankaj K. Agarwal and Jiří Matoušek. Ray shooting and parametric search. In ACM [ACM92], pages 517–

526. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p517-agarwal/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p517-agarwal/p517-

agarwal.pdf. ACM order no. 508920.

Ajtai:1992:DPT

[AM92b] Miklos Ajtai and Nimrod Megiddo. A deterministic poly(log log N)-time Nprocessor algorithm for linear programming in fixed dimension. In ACM [ACM92], pages 327–338. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p327-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p327-

ajtai/p327-ajtai.pdf. ACM

order no. 508920.

Afek:1996:CCO

[AMO96] Yehuda Afek, Yishay Mansour, and Zvi Ostfeld. Convergence complexity of optimistic rate based flow control algorithms (extended abstract). In ACM [ACM96], pages 89–98. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p89-afek/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/237814/p89-
afek/p89-afek.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Alon:1996:SCA

[AMS96] Noga Alon, Yossi Matias, and Mario Szegedy. The space complexity of approximating the frequency moments. In ACM [ACM96], pages 20– 29. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p20-alon/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/237814/p20-
alon/p20-alon.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Angluin:1992:CLT

[Ang92]  Dana Angluin.  Computa-

tional learning theory: survey and selected bibliography. In ACM [ACM92], pages 351–369. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p351-angluin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p351-angluin/p351-

angluin.pdf. ACM order no. 508920.

Ambainis:1999:DQC

[ANTSV99] Andris Ambainis, Ashwin Nayak, Ammon Ta-Shma, and Umesh Vazirani. Dense quantum coding and a lower bound for 1-way quantum automata. In ACM [ACM99], pages 376– 383. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p376-ambainis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p376-ambainis/p376-
ambainis.pdf. ACM order

number 508990.

Awerbuch:1992:AAD

[APSPS92] Baruch Awerbuch, Boaz Patt-Shamir, David Peleg, and Michael Saks. Adapting to asynchronous dynamic networks (extended abstract). In ACM [ACM92], pages 557– 570. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p557-awerbuch/
;   http://www.acm.org/pubs/

articles/proceedings/stoc/
129712/p557-awerbuch/p557-
awerbuch.pdf. ACM order no.

508920.

Arora:1998:ANH

[Aro98] Sanjeev Arora. The approximability of NP-hard problems. In ACM [ACM98], pages 337–348. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p337-arora/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p337-
arora/p337-arora.pdf. ACM

order number 508980.

Arora:1998:ASE

[ARR98] Sanjeev Arora, Prabhakar Raghavan, and Satish Rao. Approximation schemes for Euclidean k-medians and related problems. In ACM [ACM98], pages 106–113. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p106-arora/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p106-
arora/p106-arora.pdf. ACM

order number 508980.

Agarwal:1995:MPS

[ART95] Pankaj K. Agarwal, Prabhakar Raghavan, and Hisao Tamaki. Motion planning for a steering-constrained robot through moderate obstacles. In ACM [ACM95], pages 343–

352. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p343-agarwal/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p343-agarwal/p343-

agarwal.pdf. ACM order no.

508950.

Arora:1994:SQD

[ARV94] Sanjeev Arora, Yuval Rabani, and Umesh Vazirani. Simulating quadratic dynamical systems is PSPACE-complete (preliminary version). In ACM [ACM94], pages 459–

467. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p459-arora/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p459-
                                        arora/p459-arora.pdf. ACM

order no. 508930.

Arora:1997:ILD

[AS97a] Sanjeev Arora and Madhu Sudan. Improved low-degree testing and its applications. In ACM [ACM97], pages 485–

495. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p485-arora/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/258533/p485-
                                        arora/p485-arora.pdf. ACM

order no. 508970.

Awerbuch:1997:OAS

[AS97b] Baruch Awerbuch and Tripurari Singh. Online algorithms for selective multicast and maximal dense trees. In ACM [ACM97], pages 354–

362. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p354-awerbuch/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p354-awerbuch/p354-
awerbuch.pdf. ACM order no.

508970.

Aspnes:1997:LBD

[Asp97] James Aspnes. Lower bounds for distributed coin-flipping and randomized consensus. In ACM [ACM97], pages 559–

568. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p559-aspnes/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p559-aspnes/p559-aspnes.

pdf. ACM order no. 508970.

Alon:1990:STG

[AST90] N. Alon, P. Seymour, and R. Thomas. A separator theorem for graphs with an excluded minor and its applications. In ACM [ACM90], pages 293–299. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p293-alon/. ACM

order no. 508900.

Armoni:1997:XXX

[ATSWZ97] Roy Armoni, Amnon Ta-Shma, Avi Wigderson, and Shiyu Zhou. SL ⊆ L^{4/3}. In ACM [ACM97], pages 230–

239. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p230-armoni/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p230-armoni/p230-armoni.

pdf. ACM order no. 508970.

Abiteboul:1991:GCC

[AV91] Serge Abiteboul and Victor Vianu. Generic computation and its complexity. In ACM [ACM91], pages 209–

219. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p209-abiteboul/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p209-abiteboul/p209-
abiteboul.pdf. ACM order no.

508910.

Anderson:1991:WFP

[AW91] Richard J. Anderson and Heather Woll. Wait-free parallel algorithms for the union-find problem. In ACM [ACM91], pages 370–380. ISBN 0-89791-397-3. LCCN QA 76.6 A13

1991. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/103418/p370-anderson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p370-anderson/p370-
anderson.pdf. ACM order no.

508910.

Aspnes:1996:MCD

[AW96] James Aspnes and Orli Waarts.

Modular competitiveness for distributed algorithms. In ACM [ACM96], pages 237–

246. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p237-aspnes/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p237-aspnes/p237-aspnes.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Angluin:1996:RNR

[AWZ96] Dana Angluin, Jeffery Westbrook, and Wenhong Zhu. Robot navigation with range queries. In ACM [ACM96], pages 469–478. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p469-angluin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p469-angluin/p469-

angluin.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Alon:1994:CCN

[AYZ94] Noga Alon, Raphy Yuster, and Uri Zwick. Color-coding: a new method for finding simple paths, cycles and other small subgraphs within large graphs. In ACM [ACM94], pages 326–335. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p326-alon/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p326-
                                        alon/p326-alon.pdf. ACM or-

der no. 508930.

Andrews:1998:SRN

[AZ98] Matthew Andrews and Lisa Zhang. Stability results for networks with input and output blocking. In ACM [ACM98], pages 369–377. ISBN 0-89791-962-9. LCCN QA75.5 .A14

                                        1998. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/276698/p369-andrews/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p369-andrews/p369-

andrews.pdf. ACM order num-

ber 508980.

Andrews:1999:PRA

[AZ99] Matthew Andrews and Lisa Zhang. Packet routing with arbitrary end-to-end delay requirements. In ACM [ACM99], pages 557–565. ISBN 1-58113-067-8. LCCN QA75.5 .A14

                                        1999. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/301250/p557-andrews/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p557-andrews/p557-

andrews.pdf. ACM order num-

ber 508990.

Babai:1991:LEV

[Bab91]  László Babai.  Local ex-

pansion of vertex-transitive graphs and random generation in finite groups. In ACM [ACM91], pages 164–

174. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p164-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p164-
babai/p164-babai.pdf. ACM

order no. 508910.

Babai:1997:PEH

[Bab97]  László Babai.  Paul Erdős

(1913–1996): his influence on the theory of computing. In ACM [ACM97], pages 383–

401. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p383-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p383-
babai/p383-babai.pdf. ACM

order no. 508970.

Baker:1993:TPP

[Bak93] Brenda S. Baker. A theory of parameterized pattern matching: algorithms and applications. In ACM [ACM93], pages 71–80. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p71-baker/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p71-
baker/p71-baker.pdf. ACM

order no. 508930.

Ben-Amram:1999:BSL

[BAP99] Amir M. Ben-Amram and Holger Petersen. Backing up in singly linked lists. In ACM [ACM99], pages 780– 786. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p780-ben-amram/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p780-ben-amram/p780-
                                        ben-amram.pdf. ACM order

number 508990.

Barvinok:1992:FTS

[Bar92] Alexander I. Barvinok. Feasibility testing for systems of real quadratic equations. In ACM [ACM92], pages 126– 132. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p126-barvinok/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p126-barvinok/p126-
                                        barvinok.pdf. ACM order no.

508920.

Bartal:1998:AAM

[Bar98] Yair Bartal. On approximating arbitrary metrices by tree metrics. In ACM [ACM98], pages 161–168. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p161-bartal/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/

276698/p161-bartal/p161-bartal.

pdf. ACM order number 508980.

Basu:1996:BBN

[Bas96] Saugata Basu. On bounding the Betti numbers and computing the Euler characteristic of semi-algebraic sets. In ACM [ACM96], pages 408–

417. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p408-basu/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p408-
basu/p408-basu.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Bartal:1997:PCA

[BBBT97] Yair Bartal, Avrim Blum, Carl Burch, and Andrew Tomkins. A polylog(n)-competitive algorithm for metrical task systems. In ACM [ACM97], pages 711–719. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p711-bartal/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p711-bartal/p711-bartal.

pdf. ACM order no. 508970.

Beigel:1998:NME

[BBF98] Richard Beigel, Harry Buhrman, and Lance Fortnow. NP might not be as easy as detecting unique solutions. In ACM [ACM98], pages 203–

208. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p203-beigel/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p203-beigel/p203-beigel.

pdf. ACM order number 508980.

Barrington:1992:RBF

[BBR92] David A. Mix Barrington, Richard Beigel, and Steven Rudich. Representing Boolean functions as polynomials modulo composite numbers. In ACM [ACM92], pages 455–

461. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p455-mix_barrington/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p455-barrington/p455-
barrington.pdf. ACM order

no. 508920.

Babai:1992:SC

[BBTN92] László Babai, Robert Beals, and Pál Takácsi-Nagy. Symmetry and complexity. In ACM [ACM92], pages 438–

449. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p438-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p438-
babai/p438-babai.pdf. ACM

order no. 508920.

Bellantoni:1992:NRT

[BC92] Stephen Bellantoni and Stephen Cook. A new recursiontheoretic characterization of the polytime functions (extended abstract). In ACM [ACM92], pages 283–293. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p283-bellantoni/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p283-bellantoni/p283-
bellantoni.pdf. ACM order

no. 508920.

Berman:1997:LAS

[BC97] Piotr Berman and Chris Coulston. On-line algorithms for Steiner tree problems (extended abstract). In ACM [ACM97], pages 344–

353. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p344-berman/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p344-berman/p344-berman.

pdf. ACM order no. 508970.

Blum:1994:MLP

[BCC^{+}94] Avrim Blum, Prasad Chalasani, Don Coppersmith, Bill Pulleyblank, Prabhakar Raghavan, and Madhu Sudan. The minimum latency problem. In ACM [ACM94], pages 163–

171. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/

pubs/citations/proceedings/
stoc/195058/p163-blum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p163-
blum/p163-blum.pdf. ACM or-

der no. 508930.

Beame:1995:RCN

[BCE^{+}95] Paul Beame, Stephen Cook, Jeff Edmonds, Russell Impagliazzo, and Toniann Pitassi. The relative complexity of NP search problems. In ACM [ACM95], pages 303–

314. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p303-beame/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p303-
beame/p303-beame.pdf. ACM

order no. 508950.

Babai:1991:FMC

[BCF^{+}91] László Babai, Gene Cooperman, Larry Finkelstein, Eugene Luks, and Ákos Seress. Fast Monte Carlo algorithms for permutation groups. In ACM [ACM91], pages 90–

100. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p90-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p90-
babai/p90-babai.pdf. ACM

order no. 508910.

Broder:1998:MWI

[BCFM98] Andrei Z. Broder, Moses Charikar, Alan M. Frieze, and Michael Mitzenmacher. Minwise independent permutations (extended abstract). In ACM [ACM98], pages 327–

336. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p327-broder/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p327-broder/p327-broder.

pdf. ACM order number 508980.

Bellare:1998:MAD

[BCK98] Mihir Bellare, Ran Canetti, and Hugo Krawczyk. A modular approach to the design and analysis of authentication and key exchange protocols (extended abstract). In ACM [ACM98], pages 419–

428. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p419-bellare/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p419-bellare/p419-

bellare.pdf. ACM order number 508980.

Brewer:1994:SEE

[BCL94] Eric A. Brewer, Frederic T.

Chong, and Tom Leighton. Scalable expanders: exploiting hierarchical random wiring. In ACM [ACM94], pages 144–152. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/195058/p144-brewer/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p144-brewer/p144-brewer.

pdf. ACM order no. 508930.

Baruah:1993:PPN

[BCPV93] S. K. Baruah, N. K. Cohen, C. G. Plaxton, and D. A. Varvel. Proportionate progress: a notion of fairness in resource allocation. In ACM [ACM93], pages 345–

354. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p345-baruah/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p345-baruah/p345-baruah.

pdf. ACM order no. 508930.

Blum:1995:CFA

[BCV95] Avrim Blum, Prasad Chalasani, and Santosh Vempala. A constant-factor approximation for the k-MST problem in the plane. In ACM [ACM95], pages 294–302. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p294-blum/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p294-
                                        blum/p294-blum.pdf. ACM or-

der no. 508950.

Bergadano:1996:LSD

[BCV96] F. Bergadano, D. Catalano, and S. Varricchio. Learning Sat-k-DNF formulas from membership queries. In ACM [ACM96], pages 126–

130. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p126-bergadano/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p126-bergadano/p126-
bergadano.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Buhrman:1998:QVC

[BCW98] Harry Buhrman, Richard Cleve, and Avi Wigderson. Quantum vs. classical communication and computation. In ACM [ACM98], pages 63–

68. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p63-buhrman/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p63-buhrman/p63-buhrman.

pdf. ACM order number 508980.

Ben-David:1992:CFS

[BD92]  Shai Ben-David.  Can fi-

nite samples detect singularities of real-valued functions? In ACM [ACM92], pages 390–399. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p390-ben-david/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        129712/p390-ben-david/p390-
                                        ben-david.pdf. ACM order no.

508920.

Borchers:1995:SRG

[BD95] Al Borchers and Ding-Zhu Du. The k-Steiner ratio in graphs. In ACM [ACM95], pages 641– 649. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p641-borchers/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p641-borchers/p641-
                                        borchers.pdf. ACM order no.

508950.

Ben-David:1990:PRO

[BDBK^{+}90] S. Ben-David, A. Borodin, R. Karp, G. Tardos, and A. Wigderson. On the power of randomization in online algorithms. In ACM [ACM90], pages 379–386. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p379-ben-david/

. ACM order no. 508900.

Ben-David:1997:CTL

[BDBK97] Shai Ben-David, Nader H. Bshouty, and Eyal Kushilevitz. A composition theorem for learning algorithms with applications to geometric concept classes. In ACM [ACM97], pages 324– 333. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/258533/p324-ben-david/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p324-ben-david/p324-
ben-david.pdf. ACM order no.

508970.

Beaver:1996:AZK

[Bea96a]  Donald Beaver.  Adaptive

zero knowledge and computational equivocation (extended abstract). In ACM [ACM96], pages 629–638. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p629-beaver/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p629-beaver/p629-beaver.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Beaver:1996:CPC

[Bea96b] Donald Beaver. Correlated pseudorandomness and the complexity of private computations. In ACM [ACM96], pages 479–488. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p479-beaver/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p479-beaver/p479-beaver.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Beals:1997:QCF

[Bea97a] Robert Beals. Quantum computation of Fourier transforms over symmetric groups. In ACM [ACM97], pages 48–

53. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p48-beals/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/258533/p48-
                                        beals/p48-beals.pdf. ACM

order no. 508970.

Beaver:1997:CBC

[Bea97b] Donald Beaver. Commoditybased cryptography (extended abstract). In ACM [ACM97], pages 446–455. ISBN 0-89791-888-6. LCCN QA76.5 .A849

                                        1997. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/258533/p446-beaver/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p446-beaver/p446-beaver.

pdf. ACM order no. 508970.

Beigel:1992:WDE

[Bei92] Richard Beigel. When do extra majority gates help? In ACM [ACM92], pages 450–

454. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p450-beigel/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p450-beigel/p450-beigel.

pdf. ACM order no. 508920.

Benczur:1994:AUC

[Ben94] András A. Benczúr. Augmenting undirected connectivity in RNC and in randomized Õ(n^{3}) time. In ACM [ACM94], pages 658–667. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p658-benczur/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p658-benczur/p658-

benczur.pdf. ACM order no. 508930.

Barnes:1993:SRW

[BF93] Greg Barnes and Uriel Feige.

Short random walks on graphs. In ACM [ACM93], pages 728–

737. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Beame:1999:OBP

[BF99] Paul Beame and Faith E.

Fich. Optimal bounds for the predecessor problem. In ACM [ACM99], pages 295–

304. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p295-beame/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p295-
beame/p295-beame.pdf. ACM

order number 508990.

Bodlaender:1994:BNC

[BFH94] Hans L. Bodlaender, Michael R.

Fellows, and Michael T. Hallett. Beyond NP-completeness for problems of bounded width (extended abstract): hardness for the W hierarchy. In ACM [ACM94], pages 449–

458. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p449-bodlaender/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p449-bodlaender/p449-
bodlaender.pdf. ACM order

no. 508930.

Blum:1994:WLD

[BFJ^{+}94] Avrim Blum, Merrick Furst, Jeffrey Jackson, Michael Kearns, Yishay Mansour, and Steven Rudich. Weakly learning DNF and characterizing statistical query learning using Fourier analysis. In ACM [ACM94], pages 253–262. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p253-blum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p253-
blum/p253-blum.pdf. ACM or-

der no. 508930.

Bartal:1992:NAA

[BFKV92] Yair Bartal, Amos Fiat, Howard Karloff, and Rakesh Vohra. New algorithms for an ancient scheduling problem. In ACM [ACM92], pages 51–58. ISBN 0-89791-511-9. LCCN QA76.A15

1992. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/129712/p51-bartal/;
http://www.acm.org/pubs/articles/

proceedings/stoc/129712/p51-
bartal/p51-bartal.pdf. ACM

order no. 508920.

Bartal:1996:LBL

[BFL96] Yair Bartal, Amos Fiat, and Stefano Leonardi. Lower bounds for on-line graph problems with application to online circuit and optical routing. In ACM [ACM96], pages 531–540. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p531-bartal/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p531-bartal/p531-bartal.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Babai:1991:CCP

[BFLS91] László Babai, Lance Fortnow, Leonid A. Levin, and Mario Szegedy. Checking computations in polylogarithmic time. In ACM [ACM91], pages 21–32. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p21-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p21-
babai/p21-babai.pdf. ACM

order no. 508910.

Bartal:1992:CAD

[BFR92] Yair Bartal, Amos Fiat, and Yuval Rabani. Competitive algorithms for distributed data management (extended abstract). In ACM [ACM92], pages 39–50. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p39-bartal/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p39-
bartal/p39-bartal.pdf. ACM

order no. 508920.

Bender:1998:PPE

[BFR^{+}98] Michael A. Bender, Antonio Fernández, Dana Ron, Amit Sahai, and Salil Vadhan. The power of a pebble: exploring and mapping directed graphs. In ACM [ACM98], pages 269–

278. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p269-bender/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p269-bender/p269-bender.

pdf. ACM order number 508980.

Broder:1991:FHH

[BFS91] Andrei Z. Broder, Alan M. Frieze, and Eli Shamir. Finding hidden Hamiltonian cycles. In ACM [ACM91], pages 182–189. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p182-broder/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p182-broder/p182-broder.

pdf. ACM order no. 508910.

Broder:1992:ECE

[BFU92] Andrei Z. Broder, Alan M.

Frieze, and Eli Upfal. Existence and construction of edge disjoint paths on expander graphs. In ACM [ACM92], pages 140–149. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p140-broder/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p140-broder/p140-broder.

pdf. ACM order no. 508920.

Broder:1997:SDP

[BFU97] Andrei Z. Broder, Alan M.

Frieze, and Eli Upfal. Static and dynamic path selection on expander graphs (preliminary version): a random walk approach. In ACM [ACM97], pages 531–

539. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p531-broder/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p531-broder/p531-broder.

pdf. ACM order no. 508970.

Breslauer:1991:LBP

[BG91] Dany Breslauer and Zvi Galil.

A lower bound for parallel string matching. In ACM [ACM91], pages 439–

443. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p439-breslauer/

                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p439-breslauer/p439-
                                        breslauer.pdf. ACM order no.

508910.

Borowsky:1993:GFI

[BG93] Elizabeth Borowsky and Eli Gafni. Generalized FLP impossibility result for t-resilient asynchronous computations. In ACM [ACM93], pages 91–

100. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p91-borowsky/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p91-borowsky/p91-borowsky.

pdf. ACM order no. 508930.

Bellare:1995:ICA

[BGG95] Mihir Bellare, Oded Goldreich, and Shafi Goldwasser. Incremental cryptography and application to virus protection. In ACM [ACM95], pages 45–56. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p45-bellare/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p45-bellare/p45-bellare.

pdf. ACM order no. 508950.

Buss:1999:LGB

[BGIP99] Sam Buss, Dima Grigoriev, Russell Impagliazzo, and Toniann Pitassi. Linear gaps between degrees for the polynomial calculus modulo distinct primes. In ACM [ACM99], pages 547–556. ISBN 1-58113-067-8. LCCN QA75.5 .A14

1999. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/301250/p547-buss/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p547-
buss/p547-buss.pdf. ACM or-

der number 508990.

Babai:1996:EBG

[BGK^{+}96] László Babai, Anna Gál, János Kollár, Lajos Rónyai, Tibor Szabó, and Avi Wigderson. Extremal bipartite graphs and superpolynomial lower bounds for monotone span programs. In ACM [ACM96], pages 603–

611. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p603-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p603-
babai/p603-babai.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Bennett:1993:TCI

[BGL^{+}93] Charles H. Bennett, Péter Gács, Ming Li, Paul M. B. Vitányi, and Wojciech H. Zurek. Thermodynamics of computation and information distance. In ACM [ACM93], pages 21–30. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p21-bennett/;

                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p21-bennett/p21-bennett.

pdf. ACM order no. 508930.

Bellare:1993:EPC

[BGLR93] M. Bellare, S. Goldwasser, C. Lund, and A. Russeli. Efficient probabilistically checkable proofs and applications to approximations. In ACM [ACM93], pages 294–304. ISBN 0-89791-591-7. LCCN QA 76.6

                                        A13 1993. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/167088/p294-bellare/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p294-bellare/p294-

bellare.pdf. ACM order no.

508930.

Bellare:1994:EPC

[BGLR94] M. Bellare, S. Goldwasser, C. Lund, and A. Russell. Efficient probabilistic checkable proofs and applications to approximation. In ACM [ACM94], page 820. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p820-bellare/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p820-bellare/p820-

bellare.pdf. ACM order no.

508930.

Bshouty:1996:NTD

[BGM^{+}96] Nader H. Bshouty, Sally A. Goldman, H. David Mathias, Subhash Suri, and Hisao Tamaki. Noise-tolerant distributionfree learning of general geometric concepts. In ACM [ACM96], pages 151–160. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p151-bshouty/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p151-bshouty/p151-

bshouty.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Bern:1993:LAC

[BGR93] Marshall Bern, Daniel Greene, and Arvind Raghunathan. Online algorithms for cache sharing. In ACM [ACM93], pages 422–430. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Bern:1990:OAL

[BGRS90] M. Bern, D. H. Greene, A. Raghunathan, and M. Sudan. Online algorithms for locating checkpoints. In ACM [ACM90], pages 359–

368. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p359-bern/. ACM

order no. 508900.

Beigel:1998:OHB

[BH98] Richard Beigel and Tirza Hirst.

One help-bit doesn’t help. In ACM [ACM98], pages 124–

130. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p124-beigel/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p124-beigel/p124-beigel.

pdf. ACM order number

508980.

Bshouty:1992:LAR

[BHH92] Nader H. Bshouty, Thomas R. Hancock, and Lisa Hellerstein. Learning arithmetic read-once formulas. In ACM [ACM92], pages 370–381. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p370-bshouty/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p370-bshouty/p370-

bshouty.pdf. ACM order no.

508920.

Babai:1998:CMB

[BHK98] László Babai, Thomas P. Hayes, and Peter G. Kimmel. The cost of the missing bit: communication complexity with help. In ACM [ACM98], pages 673–682. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p673-babai/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p673-
                                        babai/p673-babai.pdf. ACM

order number 508980.

Beame:1992:ELB

[BIK^{+}92] Paul Beame, Russell Impagliazzo, Jan Krajíček, Toniann Pitassi, Pavel Pudlák, and Alan Woods. Exponential lower bounds for the pigeonhole principle. In ACM [ACM92], pages 200–220. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p200-beame/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p200-
beame/p200-beame.pdf. ACM

order no. 508920.

Beimel:1999:OWF

[BIKM99] Amos Beimel, Yuval Ishai, Eyal Kushilevitz, and Tal Malkin. One-way functions are essential for single-server private information retrieval. In ACM [ACM99], pages 89–

98. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p89-beimel/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p89-
beimel/p89-beimel.pdf. ACM

order number 508990.

Blum:1991:LAS

[BJL^{+}91] Avrim Blum, Tao Jiang, Ming Li, John Tromp, and Mihalis Yannakakis. Linear approximation of shortest superstrings. In ACM [ACM91], pages 328–

336. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p328-blum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p328-

                                        blum/p328-blum.pdf. ACM or-

der no. 508910.

Borgstrom:1993:CBS

[BK93] Ryan S. Borgstrom and S. Rao Kosaraju. Comparison-based search in the presence of errors. In ACM [ACM93], pages 130–136. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p130-borgstrom/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p130-borgstrom/p130-
                                        borgstrom.pdf. ACM order no.

508930.

Benczur:1996:AMC

[BK96] András A. Benczúr and David R. Karger. Approximating st minimum cuts in Õ(n^{2}) time. In ACM [ACM96], pages 47–55. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p47-benczur/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p47-benczur/p47-benczur.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Berger:1996:RTD

[BKL96] Bonnie Berger, Jon Kleinberg, and Tom Leighton. Reconstructing a three-dimensional model with arbitrary errors. In ACM [ACM96], pages 449–

458. ISBN 0-89791-785-5.

LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p449-berger/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p449-berger/p449-berger.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Beame:1998:CUP

[BKPS98] Paul Beame, Richard Karp, Toniann Pitassi, and Michael Saks. On the complexity of unsatisfiability proofs for random k-CNF formulas. In ACM [ACM98], pages 561–

571. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p561-beame/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p561-
beame/p561-beame.pdf. ACM

order number 508980.

Borodin:1996:AQT

[BKR^{+}96] Allan Borodin, Jon Kleinberg, Prabhakar Raghavan, Madhu Sudan, and David P. Williamson. Adversarial queueing theory. In ACM [ACM96], pages 376–385. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p376-borodin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p376-borodin/p376-

borodin.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Blum:1998:SDR

[BKRV98] Avrim Blum, Goran Konjevod, R. Ravi, and Santosh Vempala. Semi-definite relaxations for minimum bandwidth and other vertex-ordering problems. In ACM [ACM98], pages 100–105. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p100-blum/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p100-
                                        blum/p100-blum.pdf. ACM or-

der number 508980.

Buchsbaum:1998:LTP

[BKRW98] Adam L. Buchsbaum, Haim Kaplan, Anne Rogers, and Jeffery R. Westbrook. Linear-time pointer-machine algorithms for least common ancestors, MST verification, and dominators. In ACM [ACM98], pages 279–

288. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p279-buchsbaum/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p279-buchsbaum/p279-
                                        buchsbaum.pdf. ACM order

number 508980.

Beame:1992:RVN

[BL92] Paul Beame and Joan Lawry. Randomized versus nondeterministic communication complexity. In ACM [ACM92], pages 188–199. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p188-beame/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p188-
beame/p188-beame.pdf. ACM

order no. 508920.

Blumofe:1993:SES

[BL93] Robert D. Blumofe and Charles E.

Leiserson. Space-efficient scheduling of multithreaded computations. In ACM [ACM93], pages 362–371. ISBN 0-89791-591-7. LCCN QA 76.6

A13 1993. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/167088/p362-blumofe/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p362-blumofe/p362-

blumofe.pdf. ACM order no. 508930.

Blum:1990:STC

[BLR90] M. Blum, M. Luby, and R. Rubinfeld. Self-testing/correcting with applications to numerical problems. In ACM [ACM90], pages 73–83. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p73-blum/. ACM

order no. 508900.

Blum:1990:LBF

[Blu90] A. Blum. Learning boolean functions in an infinite attribute space. In ACM [ACM90], pages 64–72. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p64-blum/. ACM

order no. 508900.

Bjorner:1992:LDT

[BLY92] Anders Björner, László Lovász, and Andrew C. C. Yao. Linear decision trees: volume estimates and topological bounds. In ACM [ACM92], pages 170–177. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p170-bjorner/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p170-bjorner/p170-

bjorner.pdf. ACM order no. 508920.

Bellare:1990:PZK

[BMO90a] M. Bellare, S. Micali, and R. Ostrovsky. Perfect zeroknowledge in constant rounds. In ACM [ACM90], pages 482–493. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p482-bellare/.

ACM order no. 508900.

Bellare:1990:TCS

[BMO90b] M. Bellare, S. Micali, and R. Ostrovsky. The (true) complexity of statistical zero knowledge. In ACM [ACM90], pages 494–502. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/100216/p494-bellare/.

ACM order no. 508900.

Beaver:1990:RCS

[BMR90] D. Beaver, S. Micali, and P. Rogaway. The round complexity of secure protocols. In ACM [ACM90], pages 503–513. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p503-beaver/.

ACM order no. 508900.

Boppana:1993:BCP

[BN93] Ravi B. Boppana and Babu O.

Narayanan. The biased coin problem. In ACM [ACM93], pages 252–257. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p252-boppana/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p252-boppana/p252-

boppana.pdf. ACM order no. 508930.

Bar-Noy:1995:BAP

[BNCK^{+}95] Amotz Bar-Noy, Ran Canetti, Shay Kutten, Yishay Mansour, and Baruch Schieber. Bandwidth allocation with preemption. In ACM [ACM95], pages 616–625. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p616-bar-noy/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p616-bar-noy/p616-

bar-noy.pdf. ACM order no. 508950.

Bar-Noy:1998:MHN

[BNGNS98] Amotz Bar-Noy, Sudipto Guha, Joseph (Seffi) Naor, and Baruch Schieber. Multicasting in heterogeneous networks. In ACM [ACM98], pages 448– 453. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p448-bar-noy/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p448-bar-noy/p448-

bar-noy.pdf. ACM order number 508980.

Bar-Noy:1999:ATM

[BNGNS99] Amotz Bar-Noy, Sudipto Guha, Joseph (Seffi) Naor, and Baruch Schieber. Approximating the throughput of multiple machines under real-time scheduling. In ACM [ACM99], pages 622–631. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p622-bar-noy/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p622-bar-noy/p622-

bar-noy.pdf. ACM order number 508990.

Beals:1995:MCN

[BNT95] Robert Beals, Tetsuro Nishino, and Keisuke Tanaka. More on the complexity of negationlimited circuits. In ACM [ACM95], pages 585–595. ISBN 0-89791-718-9. LCCN QA 76.6

A13 1995. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/225058/p585-beals/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p585-
beals/p585-beals.pdf. ACM

order no. 508950.

Ben-Or:1993:ASC

[BOCG93] Michael Ben-Or, Ran Canetti, and Oded Goldreich. Asynchronous secure computation. In ACM [ACM93], pages 52–

61. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p52-ben-or/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p52-
ben-or/p52-ben-or.pdf. ACM

order no. 508930.

Bodlaender:1993:LTA

[Bod93] Hans L. Bodlaender. A linear time algorithm for finding tree-decompositions of small treewidth. In ACM [ACM93], pages 226–234. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p226-bodlaender/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p226-bodlaender/p226-
bodlaender.pdf. ACM order

no. 508930.

Borodin:1999:LBH

[BOR99a] Allan Borodin, Rafail Ostrovsky, and Yuval Rabani. Lower bounds for high dimensional nearest neighbor search and related problems. In ACM [ACM99], pages 312– 321. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p312-borodin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p312-borodin/p312-

borodin.pdf. ACM order number 508990.

Borodin:1999:SAA

[BOR99b] Allan Borodin, Rafail Ostrovsky, and Yuval Rabani. Subquadratic approximation algorithms for clustering problems in high dimensional spaces. In ACM [ACM99], pages 435– 444. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p435-borodin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p435-borodin/p435-

borodin.pdf. ACM order number 508990.

Brightwell:1992:TSP

[BOW92] Graham Brightwell, Teunis J. Ott, and Peter Winkler. Target shooting with programmed random variables. In ACM [ACM92], pages 691– 698. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p691-brightwell/
;   http://www.acm.org/pubs/

articles/proceedings/stoc/
129712/p691-brightwell/p691-
brightwell.pdf. ACM order

no. 508920.

Board:1990:NOA

[BP90] R. Board and L. Pitt. On the necessity of Occam algorithms. In ACM [ACM90], pages 54–63. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p54-board/. ACM

order no. 508900.

Bellare:1992:MZK

[BP92] Mihir Bellare and Erez Petrank.

Making zero-knowledge provers efficient. In ACM [ACM92], pages 711–722. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p711-bellare/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p711-bellare/p711-

bellare.pdf. ACM order no. 508920.

Bonet:1995:LBC

[BPR95] Maria Bonet, Toniann Pitassi, and Ran Raz. Lower bounds for cutting planes proofs with small coefficients. In ACM [ACM95], pages 575–

584. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p575-bonet/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p575-

bonet/p575-bonet.pdf. ACM

order no. 508950.

Basu:1996:CRS

[BPR96] S. Basu, R. Pollack, and M.- F. Roy. Computing roadmaps of semi-algebraic sets (extended abstract). In ACM [ACM96], pages 168–173. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p168-basu/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p168-
basu/p168-basu.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Bonet:1996:CET

[BPWY96] Maria Bonet, Cynthia Phillips, Tandy J. Warnow, and Shibu Yooseph. Constructing evolutionary trees in the presence of polymorphic characters. In ACM [ACM96], pages 220–229. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p220-bonet/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p220-
bonet/p220-bonet.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Barnes:1991:DAU

[BR91] Greg Barnes and Walter L. Ruzzo. Deterministic algorithms for undirected s-t connectivity using polynomial time and sublinear space. In ACM [ACM91], pages 43–

53. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p43-barnes/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p43-
barnes/p43-barnes.pdf. ACM

order no. 508910.

Blum:1992:FLT

[BR92] Avrim Blum and Steven Rudich.

Fast learning of k-term DNF formulas with queries. In ACM [ACM92], pages 382–

389. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p382-blum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p382-
blum/p382-blum.pdf. ACM or-

der no. 508920.

Bellare:1995:PSS

[BR95] Mihir Bellare and Phillip Rogaway. Provably secure session key distribution: the three party case. In ACM [ACM95], pages 57–66. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p57-bellare/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p57-bellare/p57-bellare.

pdf. ACM order no. 508950.

Borodin:1991:CPL

[BRIS91] Allan Borodin, Prabhakar Raghavan, Sandy Irani, and Baruch Schieber. Competitive paging with locality of reference. In ACM [ACM91], pages 249–259. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p249-borodin/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p249-borodin/p249-

borodin.pdf. ACM order no. 508910.

Beigel:1991:PCU

[BRS91a] Richard Beigel, Nick Reingold, and Daniel Spielman. PP is closed under intersection. In ACM [ACM91], pages 1–9. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p1-beigel/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p1-
beigel/p1-beigel.pdf. ACM

order no. 508910.

Blum:1991:NUG

[BRS91b] Avrim Blum, Prabhakar Raghavan, and Baruch Schieber. Navigating in unfamiliar geometric terrain. In ACM [ACM91], pages 494–504. ISBN 0-89791-397-3. LCCN QA 76.6 A13

1991. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/103418/p494-blum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p494-
blum/p494-blum.pdf. ACM or-

der no. 508910.

Borodin:1993:HMC

[BRSU93] Allan Borodin, Prabhakar Raghavan, Baruch Scheiber, and Eli Upfal. How much can hardware help routing? In ACM [ACM93], pages 573–582. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Blum:1996:CFA

[BRV96] Avrim Blum, R. Ravi, and Santosh Vempala. A constantfactor approximation algorithm for the k MST problem (extended abstract). In ACM [ACM96], pages 442–

448. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p442-blum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p442-
blum/p442-blum.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Buchmann:1991:CNF

[BS91] Johannes Buchmann and Victor Shoup. Constructing nonresidues in finite fields and the extended Riemann hypothesis. In ACM [ACM91], pages 72–79. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p72-buchmann/;
http://www.acm.org/pubs/
articles/proceedings/stoc/

103418/p72-buchmann/p72-buchmann.

pdf. ACM order no. 508910.

Beals:1992:SFC

[BS92] Robert Beals and Ákos Seress. Structure forest and composition factors for small base groups in nearly linear time. In ACM [ACM92], pages 116–125. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p116-beals/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p116-
beals/p116-beals.pdf. ACM

order no. 508920.

Bellare:1994:INA

[BS94] Mihir Bellare and Madhu Sudan. Improved non-approximability results. In ACM [ACM94], pages 184–193. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p184-bellare/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p184-bellare/p184-

bellare.pdf. ACM order no. 508930.

Bern:1996:PDT

[BS96] Marshall Bern and Amit Sahai. Pushing disks together

— the continuous-motion case. In ACM [ACM96], pages 119–

125. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p119-bern/;

http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p119-
bern/p119-bern.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Blomer:1999:CCS

[BS99] Johannes Blömer and Jean-Pierre Seifert. On the complexity of computing short linearly independent vectors and short bases in a lattice. In ACM [ACM99], pages 711–720. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p711-blomer/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p711-blomer/p711-blomer.

pdf. ACM order number 508990.

Bshouty:1996:TLD

[Bsh96] Nader H. Bshouty. Towards the learnability of DNF formulae. In ACM [ACM96], pages 131–140. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p131-bshouty/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p131-bshouty/p131-

bshouty.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Bshouty:1998:NCT

[Bsh98] Nader H. Bshouty. A new composition theorem for learning algorithms. In ACM [ACM98], pages 583–589. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p583-bshouty/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p583-bshouty/p583-

bshouty.pdf. ACM order number 508980.

Buhler:1997:FPC

[BSS97] Joe Buhler, M. Amin Shokrollahi, and Volker Stemann. Fast and precise computations of discrete Fourier transforms using cyclotomic integers. In ACM [ACM97], pages 40–47. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p40-buhler/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p40-
buhler/p40-buhler.pdf. ACM

order no. 508970.

Ben-Sasson:1999:SPN

[BSW99] Eli Ben-Sasson and Avi Wigderson. Short proofs are narrow — resolution made simple. In ACM [ACM99], pages 517–526. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p517-ben-sasson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p517-ben-sasson/p517-

ben-sasson.pdf. ACM order

number 508990.

Borodin:1990:DSU

[BT90] A. Borodin and P. Tiwari. On the decidability of sparse univariate polynomial interpolation. In ACM [ACM90], pages 535–545. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p535-borodin/.

ACM order no. 508900.

Benaloh:1994:RFS

[BT94] Josh Benaloh and Dwight Tuinstra. Receipt-free secretballot elections (extended abstract). In ACM [ACM94], pages 544–553. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p544-benaloh/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p544-benaloh/p544-

benaloh.pdf. ACM order no. 508930.

Bshouty:1995:FSM

[BT95] Nader H. Bshouty and Christino Tamon. On the Fourier spectrum of monotone functions. In ACM [ACM95], pages 219– 228. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p219-bshouty/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p219-bshouty/p219-

bshouty.pdf. ACM order no.

508950.

Broder:1996:DDR

[BU96] Andrei Broder and Eli Upfal. Dynamic deflection routing on arrays (preliminary version). In ACM [ACM96], pages 348–355. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p348-broder/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p348-broder/p348-broder.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Bernstein:1993:QCT

[BV93] Ethan Bernstein and Umesh Vazirani. Quantum complexity theory. In ACM [ACM93], pages 11–20. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p11-bernstein/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p11-bernstein/p11-
                                        bernstein.pdf. ACM order no.

508930.

Beling:1997:CCC

[BV97] Peter A. Beling and Sushil Verma. Combinatorial complexity of the central curve. In ACM [ACM97], pages 250–

255. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p250-beling/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p250-beling/p250-beling.

pdf. ACM order no. 508970.

Brightwell:1991:CLE

[BW91] Graham Brightwell and Peter Winkler. Counting linear extensions is #p-complete. In ACM [ACM91], pages 175–

181. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p175-brightwell/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p175-brightwell/p175-
brightwell.pdf. ACM order

no. 508910.

Cypher:1996:UAS

[CadHSV96] Robert Cypher, Friedhelm Meyer auf der Heide, Christian Scheideler, and Berthold Vöcking. Universal algorithms for storeand-forward and wormhole routing. In ACM [ACM96], pages 356–365. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p356-cypher/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p356-cypher/p356-cypher.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Cesa-Bianchi:1996:NTL

[CBDFS96] Nicolò Cesa-Bianchi, Eli Dichterman, Paul Fischer, and Hans Ulrich Simon. Noise-tolerant learning near the informationtheoretic bound. In ACM [ACM96], pages 141–150. ISBN 0-89791-785-5. LCCN QA 76.6

                                        A13 1996. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/237814/p141-cesa-bianchi/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p141-cesa-bianchi/
                                        p141-cesa-bianchi.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Cesa-Bianchi:1993:HUE

[CBFH^{+}93] Nicolò Cesa-Bianchi, Yoav Freund, David P. Helmbold, David Haussler, Robert E. Schapire, and Manfred K. Warmuth. How to use expert advice. In ACM [ACM93], pages 382–

391. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p382-cesa-bianchi/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p382-cesa-bianchi/
                                        p382-cesa-bianchi.pdf. ACM

order no. 508930.

Charikar:1997:ICD

[CCFM97] Moses Charikar, Chandra Chekuri, Tomás Feder, and Rajeev Motwani. Incremental clustering and dynamic information retrieval. In ACM [ACM97], pages 626–635. ISBN 0-89791-888-6. LCCN QA76.5 .A849

1997. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/258533/p626-charikar/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p626-charikar/p626-
charikar.pdf. ACM order no.

508970.

Coffman:1991:FDB

[CCG^{+}91] E. G. Coffman, C. Courcoubetis, M. R. Garey, D. S. Johnson, L. A. McGeoch, P. W. Shor, R. R. Weber, and M. Yannakakis. Fundamental discrepancies between averagecase analyses under discrete and continuous distributions: a bin packing case study. In ACM [ACM91], pages 230–

240. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p230-coffman/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p230-coffman/p230-

coffman.pdf. ACM order no. 508910.

Charikar:1998:RTD

[CCGG98] Moses Charikar, Chandra Chekuri, Ashish Goel, and Sudipto Guha. Rounding via trees: deterministic approximation algorithms for group Steiner trees and k-median. In ACM [ACM98], pages 114–

123. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/

                                        pubs/citations/proceedings/
                                        stoc/276698/p114-charikar/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p114-charikar/p114-
                                        charikar.pdf. ACM order

number 508980.

Chakrabarti:1999:LBC

[CCGL99] Amit Chakrabarti, Bernard Chazelle, Benjamin Gum, and Alexey Lvov. A lower bound on the complexity of approximate nearest-neighbor searching on the Hamming cube. In ACM [ACM99], pages 305–

311. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p305-chakrabarti/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p305-chakrabarti/p305-
                                        chakrabarti.pdf. ACM order

number 508990.

Cramer:1997:LZK

[CD97] Ronald Cramer and Ivan Damgård. Linear zero-knowledge

— a note on efficient zeroknowledge proofs and arguments. In ACM [ACM97], pages 436–445. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p436-cramer/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p436-cramer/p436-cramer.

pdf. ACM order no. 508970.

Duh:1997:ASC

[cDF97] Rong chii Duh and Martin Fürer. Approximation of k-set cover by semi-local optimization. In ACM [ACM97], pages 256–264. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p256-duh/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/258533/p256-
duh/p256-duh.pdf. ACM order

no. 508970.

Cohen:1993:RWO

[CDKT93] Robert F. Cohen, Giuseppe Di Battista, Arkady Kanevsky, and Roberto Tamassia. Reinventing the wheel: an optimal data structure for connectivity queries. In ACM [ACM93], pages 194–200. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p194-cohen/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p194-
cohen/p194-cohen.pdf. ACM

order no. 508930.

Coppersmith:1990:RWW

[CDRS90] D. Coppersmith, P. Doyle, P. Raghavan, and M. Snir. Random walks on weighted graphs, and applications to on-line algorithms. In ACM [ACM90], pages 369–378. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/

stoc/100216/p369-coppersmith/

. ACM order no. 508900.

Chazelle:1993:IBW

[CEG^{+}93] Bernard Chazelle, Herbert Edelsbrunner, Michelangelo Grigni, Leonidas Guibas, Micha Sharir, and Emo Welzl. Improved bounds on weak ϵ-nets for convex sets. In ACM [ACM93], pages 495–504. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Clegg:1996:UGB

[CEI96] Matthew Clegg, Jeffery Edmonds, and Russell Impagliazzo. Using the Groebner basis algorithm to find proofs of unsatisfiability. In ACM [ACM96], pages 174–

183. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p174-clegg/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p174-
clegg/p174-clegg.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Canetti:1996:ASM

[CFGN96] Ran Canetti, Uri Feige, Oded Goldreich, and Moni Naor. Adaptively secure multi-party computation. In ACM [ACM96], pages 639–648. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p639-canetti/;

http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p639-canetti/p639-

canetti.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Condon:1993:PCD

[CFLS93] Anne Condon, Joan Feigenbaum, Carsten Lund, and Peter Shor. Probabilistically checkable debate systems and approximation algorithms for PSPACE-hard functions. In ACM [ACM93], pages 305–

314. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p305-condon/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p305-condon/p305-condon.

pdf. ACM order no. 508930.

Coffman:1991:PCP

[CG91] E. G. Coffman and M. R.

Garey. Proof of the 4/3 conjecture for preemptive vs. nonpreemptive two-processor scheduling. In ACM [ACM91], pages 241–248. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p241-coffman/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p241-coffman/p241-

coffman.pdf. ACM order no. 508910.

Chor:1997:CPI

[CG97] Benny Chor and Niv Gilboa. Computationally private information retrieval (extended abstract). In ACM [ACM97], pages 304–313. ISBN 0-89791-888-6. LCCN QA76.5 .A849

1997. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/258533/p304-chor/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p304-
chor/p304-chor.pdf. ACM or-

der no. 508970.

Czumaj:1995:WTO

[CGG^{+}95] Artur Czumaj, Zvi Galil, Leszek Gasieniec, Kunsoo Park, and Wojciech Plandowski. Worktime-optimal parallel algorithms for string problems. In ACM [ACM95], pages 713–

722. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p713-czumaj/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p713-czumaj/p713-czumaj.

pdf. ACM order no. 508950.

Canetti:1998:ROM

[CGH98] Ran Canetti, Oded Goldreich, and Shai Halevi. The random oracle methodology, revisited (preliminary version). In ACM [ACM98], pages 209–

218. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p209-canetti/;

http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p209-canetti/p209-

canetti.pdf. ACM order number 508980.

Chen:1998:PMG

[CGP98a] Zhi-Zhong Chen, Enory Grigni, and Christos H. Papadimitriou. Planar map graphs. In ACM [ACM98], pages 514–

523. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p514-chen/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p514-
chen/p514-chen.pdf. ACM or-

der number 508980.

Crescenzi:1998:CPF

[CGP^{+}98b] Pierluigi Crescenzi, Deborah Goldman, Christos Papadimitriou, Antonio Piccolboni, and Mihalis Yannakakis. On the complexity of protein folding (extended abstract). In ACM [ACM98], pages 597–

603. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p597-crescenzi/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p597-crescenzi/p597-
crescenzi.pdf. ACM order

number 508980.

Charikar:1999:CFA

[CGTS99] Moses Charikar, Sudipto Guha, Éva Tardos, and David B. Shmoys. A constant-factor approximation algorithm for the k-median problem (extended abstract). In ACM [ACM99], pages 1–10. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p1-charikar/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p1-charikar/p1-charikar.

pdf. ACM order number 508990.

Cole:1997:TPM

[CH97] Richard Cole and Ramesh Hariharan. Tree pattern matching and subset matching in randomized O(n log^{3} m) time. In ACM [ACM97], pages 66–

75. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p66-cole/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/258533/p66-
cole/p66-cole.pdf. ACM or-

der no. 508970.

Chazelle:1994:CGR

[Cha94] Bernard Chazelle. Computational geometry: a retrospective. In ACM [ACM94], pages 75–94. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p75-chazelle/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p75-chazelle/p75-chazelle.

pdf. ACM order no. 508930.

Chazelle:1995:LBL

[Cha95]  Bernard Chazelle.  Lower

bounds for off-line range searching. In ACM [ACM95], pages 733–740. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p733-chazelle/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p733-chazelle/p733-
chazelle.pdf. ACM order no.

508950.

Condon:1994:PFA

[CHPW94] Anne Condon, Lisa Hellerstein, Samuel Pottle, and Avi Wigderson. On the power of finite automata with both nondeterministic and probabilistic states (preliminary version). In ACM [ACM94], pages 676–685. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p676-condon/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p676-condon/p676-condon.

pdf. ACM order no. 508930.

Canetti:1995:BPP

[CI95] Ran Canetti and Sandy Irani.

Bounding the power of preemption in randomized scheduling. In ACM [ACM95], pages 606–615. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p606-canetti/;

                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p606-canetti/p606-

canetti.pdf. ACM order no. 508950.

Coffman:1993:MCC

[CJSW93] E. G. Coffman, D. S. Johnson, P. W. Shor, and R. R. Weber. Markov chains, computer proofs, and average-case analysis of best fit bin packing. In ACM [ACM93], pages 412–421. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p412-coffman/.

ACM order no. 508930.

Callahan:1992:DMD

[CK92] Paul B. Callahan and S. Rao Kosaraju. A decomposition of multi-dimensional point-sets with applications to k-nearest-neighbors and n-body potential fields (preliminary version). In ACM [ACM92], pages 546–556. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p546-callahan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p546-callahan/p546-
                                        callahan.pdf. ACM order no.

508920.

Chou:1993:SCI

[CK93] Arthur W. Chou and Ker-I. Ko. Some complexity issues on the simply connected regions of the two-dimensional plane. In ACM [ACM93], pages 1–10. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p1-chou/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/167088/p1-
chou/p1-chou.pdf. ACM order

no. 508930.

Chen:1997:RRI

[CK97] Zhi-Zhong Chen and Ming-Yang Kao. Reducing randomness via irrational numbers. In ACM [ACM97], pages 200–209. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p200-chen/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p200-
chen/p200-chen.pdf. ACM or-

der no. 508970.

Capalbo:1999:SUG

[CK99a] M. R. Capalbo and S. R.

Kosaraju. Small universal graphs. In ACM [ACM99], pages 741–749. ISBN 1-58113-067-8. LCCN QA75.5 .A14

1999. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/301250/p741-capalbo/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p741-capalbo/p741-

capalbo.pdf. ACM order number 508990.

Cohen:1999:ERW

[CK99b] Edith Cohen and Haim Kaplan. Exploiting regularities in Web traffic patterns for cache replacement. In ACM [ACM99], pages 109–

118. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p109-cohen/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p109-
cohen/p109-cohen.pdf. ACM

order number 508990.

Cucker:1995:RTM

[CKK^{+}95] Felipe Cucker, Marek Karpinski, Pascal Koiran, Thomas Lickteig, and Kai Werther. On real Turing machines that toss coins. In ACM [ACM95], pages 335–342. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p335-cucker/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p335-cucker/p335-cucker.

pdf. ACM order no. 508950.

Chen:1999:OBH

[CKLW99] Gen-Huey Chen, Ming-Yang Kao, Yuh-Dauh Lyuu, and Hsing-Kuo Wong. Optimal buyand-hold strategies for financial markets with bounded daily returns. In ACM [ACM99], pages 119–128. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p119-chen/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p119-

chen/p119-chen.pdf. ACM or-

der number 508990.

Calinescu:1998:IAA

[CKR98a] Gruia C. Călinescu, Howard Karloff, and Yuval Rabani. An improved approximation algorithm for multiway cut. In ACM [ACM98], pages 48–

52. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/276698/p48-calinescu/
p48-calinescu.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/276698/p48-

calinescu/. ACM order number 508980.

Charikar:1998:ACV

[CKR98b] Moses Charikar, Samir Khuller, and Balaji Raghavachari. Algorithms for capacitated vehicle routing. In ACM [ACM98], pages 349–358. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p349-charikar/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p349-charikar/p349-
charikar.pdf. ACM order

number 508980.

Charikar:1999:TMS

[CKR^{+}99] Moses Charikar, Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew Tomkins. On targeting Markov segments. In ACM [ACM99], pages 99–

108. ISBN 1-58113-067-8.

LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p99-charikar/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p99-charikar/p99-charikar.

pdf. ACM order number

508990.

Cohen:1999:CC

[CKZ99] Edith Cohen, Haim Kaplan, and Uri Zwick. Connection caching. In ACM [ACM99], pages 612–621. ISBN 1-58113-067-8. LCCN QA75.5 .A14

                                        1999. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/301250/p612-cohen/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p612-
                                        cohen/p612-cohen.pdf. ACM

order number 508990.

Cooper:1993:FPI

[CL93] Jason Cooper and Nathan Linial. Fast perfectioninformation leader-election protocol with linear immunity. In ACM [ACM93], pages 662–671. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Clarkson:1997:NNQ

[Cla97] Kenneth L. Clarkson. Nearest neighbor queries in metric spaces. In ACM [ACM97], pages 609–617. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p609-clarkson/
                                        ;   http://www.acm.org/pubs/

articles/proceedings/stoc/
258533/p609-clarkson/p609-
clarkson.pdf. ACM order no.

508970.

Cleve:1990:TOS

[Cle90]  R. Cleve.  Towards opti-

mal simulations of formulas by bounded-width programs. In ACM [ACM90], pages 271–277. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p271-cleve/. ACM

order no. 508900.

Chen:1999:LMC

[CLP99] Fang Chen, László Lovász, and Igor Pak. Lifting Markov chains to speed up mixing. In ACM [ACM99], pages 275–

281. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p275-chen/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p275-
chen/p275-chen.pdf. ACM or-

der number 508990.

Cohen:1991:IAL

[CM91] Edith Cohen and Nimrod Megiddo. Improved algorithms for linear inequalities with two variables per inequality. In ACM [ACM91], pages 145–155. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p145-cohen/;
http://www.acm.org/pubs/articles/

                                        proceedings/stoc/103418/p145-
                                        cohen/p145-cohen.pdf. ACM

order no. 508910.

Chen:1999:PTA

[CM99] Jianer Chen and Antonio Miranda. A polynomial time approximation scheme for general multiprocessor job scheduling (extended abstract). In ACM [ACM99], pages 418–

427. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p418-chen/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p418-
                                        chen/p418-chen.pdf. ACM or-

der number 508990.

Cole:1998:RPL

[CMadH^{+}98] Richard Cole, Bruce M. Maggs, Friedhelm Meyer auf der Heide, Michael Mitzenmacher, Andréa W. Richa, Klaus Schröder, Ramesh K. Sitaraman, and Berthold Vöcking. Randomized protocols for lowcongestion circuit routing in multistage interconnection networks. In ACM [ACM98], pages 378–388. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p378-cole/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p378-
                                        cole/p378-cole.pdf. ACM or-

der number 508980.

Canetti:1998:POW

[CMR98] Ran Canetti, Daniele Micciancio, and Omer Reingold. Perfectly one-way probabilistic hash functions (preliminary version). In ACM [ACM98], pages 131–140. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p131-canetti/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p131-canetti/p131-

canetti.pdf. ACM order number 508980.

Cole:1993:MSS

[CMS93] Richard Cole, Bruce Maggs, and Ramesh Sitaraman. Multiscale self-simulation: a technique for reconfiguring arrays with faults. In ACM [ACM93], pages 561–572. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Cai:1999:HHT

[CNS99] Jin-Yi Cai, Ajay Nerurkar, and D. Sivakumar. Hardness and hierarchy theorems for probabilistic quasi-polynomial time. In ACM [ACM99], pages 726–735. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p726-cai/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/301250/p726-
cai/p726-cai.pdf. ACM order

number 508990.

Canetti:1999:SCH

[CO99] Ran Canetti and Rafail Ostrovsky. Secure computation with honest-looking parties (extended abstract): what if nobody is truly honest? In ACM [ACM99], pages 255–

264. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p255-canetti/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p255-canetti/p255-

canetti.pdf. ACM order num-

ber 508990.

Cohen:1994:PTN

[Coh94] Edith Cohen. Polylog-time and near-linear work approximation scheme for undirected shortest paths. In ACM [ACM94], pages 16–26. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p16-cohen/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p16-
                                        cohen/p16-cohen.pdf. ACM

order no. 508930.

Cole:1990:DFC

[Col90] R. Cole. On the dynamic finger conjecture for splay trees. In ACM [ACM90], pages 8–17. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p8-cole/. ACM

order no. 508900.

Cypher:1990:DSN

[CP90] R. Cypher and C. G. Plaxton. Deterministic sorting in nearly logarithmic time on the hypercube and related computers. In ACM [ACM90], pages 193–203. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p193-cypher/.

ACM order no. 508900.

Canetti:1993:FAB

[CR93] Ran Canetti and Tal Rabin.

Fast asynchronous Byzantine agreement with optimal resilience. In ACM [ACM93], pages 42–51. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p42-canetti/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p42-canetti/p42-canetti.

pdf. ACM order no. 508930.

Chaudhuri:1996:DRC

[CR96] Shiva Chaudhuri and Jaikumar Radhakrishnan. Deterministic restrictions in circuit complexity. In ACM [ACM96], pages 30–36. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p30-chaudhuri/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p30-chaudhuri/p30-
chaudhuri.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Chari:1993:ROU

[CRS93] Suresh Chari, Pankaj Rohatgi, and Aravind Srinivasan. Randomness-optimal unique element isolation, with applications to perfect matching and related problems. In ACM [ACM93], pages 458–467. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Chari:1994:IAA

[CRS94] Suresh Chari, Pankaj Rohatgi, and Aravind Srinivasan. Improved algorithms via approximations of probability distributions (extended abstract). In ACM [ACM94], pages 584–592. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p584-chari/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p584-
                                        chari/p584-chari.pdf. ACM

order no. 508930.

Choy:1992:EFT

[CS92] Manhoi Choy and Ambuj K. Singh. Efficient fault tolerant algorithms for resource allocation in distributed systems. In ACM [ACM92], pages 593–602. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p593-choy/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p593-

choy/p593-choy.pdf. ACM or-

der no. 508920.

Cheriyan:1991:APV

[CT91] Joseph Cheriyan and Ramakrishna Thurimella. Algorithms for parallel k-vertex connectivity and sparse certificates. In ACM [ACM91], pages 391–

401. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p391-cheriyan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p391-cheriyan/p391-
cheriyan.pdf. ACM order no.

508910.

Cheriyan:1996:FAS

[CT96] Joseph Cheriyan and Ramakrishna Thurimella. Fast algorithms for k-shredders and knode connectivity augmentation (extended abstract). In ACM [ACM96], pages 37–

46. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p37-cheriyan/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p37-cheriyan/p37-cheriyan.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Chung:1994:NOA

[CY94] Fan R. K. Chung and S.-T.

Yau. A near optimal algorithm for edge separators (preliminary version). In ACM [ACM94], pages 1–8. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p1-chung/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/195058/p1-
chung/p1-chung.pdf. ACM or-

der no. 508930.

Chung:1997:EFS

[CY97] F. R. K. Chung and S.-T. Yau. Eigenvalues, flows and separators of graphs. In ACM [ACM97], page 749. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997. URL http://www.

acm.org/pubs/citations/proceedings/
stoc/258533/p749-chung/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p749-
chung/p749-chung.pdf. ACM

order no. 508970.

Dietzfelbinger:1990:HDD

[DadH90] M. Dietzfelbinger and F. Meyer auf der Heide. How to distribute a dictionary in a complete network. In ACM [ACM90], pages 117–127. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p117-dietzfelbinger/

. ACM order no. 508900.

Damaschke:1998:AVN

[Dam98] Peter Damaschke. Adaptive versus nonadaptive attributeefficient learning. In ACM [ACM98], pages 590–596. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998. URL http://www.

acm.org/pubs/citations/proceedings/
stoc/276698/p590-damaschke/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p590-damaschke/p590-
damaschke.pdf. ACM order

number 508980.

deBerg:1994:LRI

[dBDS94] Mark de Berg, Katrin Dobrindt, and Otfried Schwarzkopf. On lazy randomized incremental construction. In ACM [ACM94], pages 105–114. ISBN 0-89791-663-8. LCCN QA76 .A15 1994. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/195058/p105-
de_berg/p105-de_berg.pdf;
http://www.acm.org/pubs/
citations/proceedings/stoc/
195058/p105-de_berg/. ACM

order no. 508930.

deBerg:1993:PLP

[dBMS93] Mark de Berg, Jiří Matoušek, and Otfried Schwarzkopf. Piecewise linear paths among convex obstacles. In ACM [ACM93], pages 505–514. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

DeSantis:1994:HSF

[DDFY94] Alfredo De Santis, Yvo Desmedt, Yair Frankel, and Moti Yung. How to share a function securely. In ACM [ACM94], pages 522–533. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p522-de_santis/
;   http://www.acm.org/pubs/

articles/proceedings/stoc/
195058/p522-de_santis/p522-
de_santis.pdf. ACM order no.

508930.

Dolev:1991:NMC

[DDN91] Danny Dolev, Cynthia Dwork, and Moni Naor. Nonmalleable cryptography. In ACM [ACM91], pages 542–

552. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p542-dolev/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p542-
dolev/p542-dolev.pdf. ACM

order no. 508910.

Dinur:1999:PCN

[DFK^{+}99] Irit Dinur, Eldar Fischer, Guy Kindler, Ran Raz, and Shmuel Safra. PCP characterizations of NP: towards a polynomially-small errorprobability. In ACM [ACM99], pages 29–40. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p29-dinur/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p29-
dinur/p29-dinur.pdf. ACM

order number 508990.

Dey:1996:AMS

[DG96] Tamal K. Dey and Sumanta Guha. Algorithms for manifolds and simplicial complexes in Euclidean 3-space (preliminary version). In ACM [ACM96], pages 398– 407. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p398-dey/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/237814/p398-
dey/p398-dey.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Dooly:1998:TDA

[DGS98] Daniel R. Dooly, Sally A.

Goldman, and Stephen D. Scott. TCP dynamic acknowledgment delay (extended abstract): theory and practice. In ACM [ACM98], pages 389–398. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p389-dooly/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p389-
dooly/p389-dooly.pdf. ACM

order number 508980.

Dwork:1993:CSM

[DHW93] Cynthia Dwork, Maurice Herlihy, and Orli Waarts. Contention in shared memory algorithms. In ACM [ACM93], pages 174–183. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p174-dwork/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p174-
dwork/p174-dwork.pdf. ACM

order no. 508930.

Dimitriou:1996:TAL

[DI96] Tassos Dimitriou and Russell Impagliazzo. Towards an analysis of local optimization algorithms. In ACM [ACM96], pages 304– 313. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p304-dimitriou/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p304-dimitriou/p304-
                                        dimitriou.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

DiCrescenzo:1999:SPH

[DI99] Giovanni Di Crescenzo and Russell Impagliazzo. Securitypreserving hardness-amplification for any regular one-way function. In ACM [ACM99], pages 169–178. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p169-di_crescenzo/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p169-di_crescenzo/
                                        p169-di_crescenzo.pdf. ACM

order number 508990.

Dietzfelbinger:1997:LAP

[Die97] Martin Dietzfelbinger. The linear-array problem in communication complexity resolved. In ACM [ACM97], pages 373– 382. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p373-dietzfelbinger/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p373-dietzfelbinger/
p373-dietzfelbinger.pdf. ACM

order no. 508970.

DiCrescenzo:1998:NIN

[DIO98] Giovanni Di Crescenzo, Yuval Ishai, and Rafail Ostrovsky. Non-interactive and nonmalleable commitment. In ACM [ACM98], pages 141–

150. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p141-di_crescenzo/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p141-di_crescenzo/
p141-di_crescenzo.pdf. ACM

order number 508980.

Dahlhaus:1992:CMC

[DJP^{+}92] E. Dahlhaus, D. S. Johnson, C. H. Papadimitriou, P. D. Seymour, and M. Yannakakis. The complexity of multiway cuts (extended abstract). In ACM [ACM92], pages 241–251. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p241-dahlhaus/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p241-dahlhaus/p241-
dahlhaus.pdf. ACM order no.

508920.

Dodis:1999:DNB

[DK99] Yevgeniy Dodis and Sanjeev Khanna. Design networks with bounded pairwise distance. In ACM [ACM99], pages 750–759. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p750-dodis/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p750-
                                        dodis/p750-dodis.pdf. ACM

order number 508990.

Dolev:1995:BAR

[DKKP95] Shlomi Dolev, Evangelos Kranakis, Danny Krizanc, and David Peleg. Bubbles: adaptive routing scheme for highspeed dynamic networks. In ACM [ACM95], pages 528–

537. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p528-dolev/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p528-
                                        dolev/p528-dolev.pdf. ACM

order no. 508950.

Dwork:1996:DSS

[DLN96] Cynthia Dwork, Jeffrey Lotspiech, and Moni Naor. Digital signets: self-enforcing protection of digital information (preliminary version). In ACM [ACM96], pages 489–

498. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/237814/p489-dwork/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p489-
dwork/p489-dwork.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Deng:1991:IGR

[DM91] Xiaotie Deng and Sanjeev Mahajan. Infinite games: randomization, computability, and applications to online problems. In ACM [ACM91], pages 289–298. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p289-deng/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p289-
deng/p289-deng.pdf. ACM or-

der no. 508910.

Dinitz:1999:BCB

[DMR99] Yefim Dinitz, Shlomo Moran, and Sergio Rajsbaum. Bit complexity of breaking and achieving symmetry in chains and rings (extended abstract). In ACM [ACM99], pages 265–

274. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p265-dinitz/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p265-dinitz/p265-dinitz.

pdf. ACM order number 508990.

Dinitz:1995:LCM

[DN95] Yefim Dinitz and Zeev Nutov.

A 2-level cactus model for the system of minimum and minimum + 1 edge-cuts in a graph and its incremental maintenance. In ACM [ACM95], pages 509–518. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p509-dinitz/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p509-dinitz/p509-dinitz.

pdf. ACM order no. 508950.

Dwork:1998:CZK

[DNS98] Cynthia Dwork, Moni Naor, and Amit Sahai. Concurrent zero-knowledge. In ACM [ACM98], pages 409–

418. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p409-dwork/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p409-
dwork/p409-dwork.pdf. ACM

order number 508980.

Djidjev:1991:EAG

[DR91] Hristo Djidjev and John Reif. An efficient algorithm for the genus problem with explicit construction of forbidden subgraphs. In ACM [ACM91], pages 337–347. ISBN 0-89791-397-3. LCCN QA 76.6 A13

1991. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/103418/p337-djidjev/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p337-djidjev/p337-

djidjev.pdf. ACM order no. 508910.

Diaconis:1995:WDW

[DSC95] Persi Diaconis and Laurent Saloff-Coste. What do we know about the Metropolis algorithm? In ACM [ACM95], pages 112–129. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p112-diaconis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p112-diaconis/p112-
diaconis.pdf. ACM order no.

508950.

Dor:1992:GDN

[DT92] Dorit Dor and Michael Tarsi.

Graph decomposition is NPC

— a complete proof of Holyer’s conjecture. In ACM [ACM92], pages 252–263. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p252-dor/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p252-
dor/p252-dor.pdf. ACM order

no. 508920.

DiBattista:1993:APT

[DV93] Giuseppe Di Battista and Luca Vismara. Angles of planar triangular graphs. In ACM [ACM93], pages 431–437. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Dinitz:1994:CCV

[DV94] Yefim Dinitz and Alek Vainshtein. The connectivity carcass of a vertex subset in a graph and its incremental maintenance. In ACM [ACM94], pages 716–725. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p716-dinitz/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p716-dinitz/p716-dinitz.

pdf. ACM order no. 508930.

Dwork:1992:SEB

[DW92] Cynthia Dwork and Orli Waarts. Simple and efficient bounded concurrent timestamping or bounded concurrent timestamp systems are comprehensible! In ACM [ACM92], pages 655–666. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p655-dwork/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p655-
dwork/p655-dwork.pdf. ACM

order no. 508920.

Edmonds:1997:NCM

[ECBD97] Jeff Edmonds, Donald D. Chinn, Tim Brecht, and Xiaotie Deng. Non-clairvoyant multiprocessor scheduling of jobs with changing execution characteristics (extended abstract). In ACM [ACM97], pages 120–

129. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p120-edmonds/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p120-edmonds/p120-

edmonds.pdf. ACM order no. 508970.

Edmonds:1993:TST

[Edm93]  Jeff Edmonds.  Time-space

trade-offs for undirected stconnectivity on a JAG. In ACM [ACM93], pages 718–727. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Edmonds:1999:SD

[Edm99] Jeff Edmonds. Scheduling in the dark. In ACM [ACM99], pages 179–188. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p179-edmonds/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p179-edmonds/p179-

edmonds.pdf. ACM order number 508990.

Eppstein:1993:SBS

[EGIS93] David Eppstein, Zvi Galil, Giuseppe F. Italiano, and Thomas H. Spencer. Separator based sparsification for dynamic planar graph algorithms. In ACM [ACM93], pages 208– 217. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p208-eppstein/

;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p208-eppstein/p208-
eppstein.pdf. ACM order no.

508930.

Even:1992:AGI

[EGL^{+}92] Guy Even, Oded Goldreich, Michael Luby, Noam Nisan, and Boban Veličkovic. Approximations of general independent distributions. In ACM [ACM92], pages 10–16. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p10-even/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p10-
even/p10-even.pdf. ACM or-

der no. 508920.

Ergun:1998:SC

[EKK^{+}98] Funda Ergün, Sampath Kannan, S. Ravi Kumar, Ronitt Rubinfeld, and Mahesh Viswanathan. Spot-checkers. In ACM [ACM98], pages 259–268. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998. URL http://www.

acm.org/pubs/citations/proceedings/
stoc/276698/p259-ergun/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p259-
ergun/p259-ergun.pdf. ACM

order number 508980.

Ergun:1999:FAP

[EKR99] Funda Ergün, Ravi Kumar, and Ronitt Rubinfeld. Fast approximate PCPs. In ACM [ACM99], pages 41–50. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/

pubs/citations/proceedings/
stoc/301250/p41-ergun/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p41-
ergun/p41-ergun.pdf. ACM

order number 508990.

Edmonds:1995:NOT

[EP95] Jeff Edmonds and Chung-Keung Poon. A nearly optimal time-space lower bound for directed st-connectivity on the NNJAG model. In ACM [ACM95], pages 147–

156. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p147-edmonds/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p147-edmonds/p147-

edmonds.pdf. ACM order no. 508950.

Evans:1996:LBN

[EP96] William Evans and Nicholas Pippenger. Lower bounds for noisy Boolean decision trees. In ACM [ACM96], pages 620–

628. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p620-evans/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p620-
evans/p620-evans.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Eppstein:1995:GLB

[Epp95]  David Eppstein.  Geomet-

ric lower bounds for parametric matroid optimization. In ACM [ACM95], pages 662–

671. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p662-eppstein/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p662-eppstein/p662-
eppstein.pdf. ACM order no.

508950.

Even:1990:USY

[ER90] S. Even and S. Rajsbaum. The use of a synchronizer yields maximum computation rate in distributed networks. In ACM [ACM90], pages 95–

105. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p95-even/. ACM

order no. 508900.

Ergun:1995:TML

[Erg95] Funda Ergün. Testing multivariate linear functions: overcoming the generator bottleneck. In ACM [ACM95], pages 407–416. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p407-ergun/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p407-
ergun/p407-ergun.pdf. ACM

order no. 508950.

Feige:1995:RGP

[Fei95] Uriel Feige. Randomized graph products, chromatic numbers, and Lovasz j-function. In ACM [ACM95], pages 635–

640. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p635-feige/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p635-
feige/p635-feige.pdf. ACM

order no. 508950.

Feige:1996:TAS

[Fei96]  Uriel Feige.  A threshold

of ln n for approximating set cover (preliminary version). In ACM [ACM96], pages 314–

318. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p314-feige/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p314-
feige/p314-feige.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Feige:1998:ABV

[Fei98] Uriel Feige. Approximating the bandwidth via volume respecting embeddings (extended abstract). In ACM [ACM98], pages 90–99. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p90-feige/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p90-
feige/p90-feige.pdf. ACM

order number 508980.

Feige:1999:NPP

[Fei99] Uriel Feige. Nonmonotonic phenomena in packet routing. In ACM [ACM99], pages 583–591. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p583-feige/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p583-
                                        feige/p583-feige.pdf. ACM

order number 508990.

Ferragina:1995:FDD

[FG95] Paolo Ferragina and Roberto Grossi. A fully-dynamic data structure for external substring search. In ACM [ACM95], pages 693–702. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p693-ferragina/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p693-ferragina/p693-
                                        ferragina.pdf. ACM order no.

508950.

Frankel:1996:WBC

[FGY96] Yair Frankel, Peter Gemmell, and Moti Yung. Witness-based cryptographic program checking and robust function sharing. In ACM [ACM96], pages 499–508. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p499-frankel/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/

237814/p499-frankel/p499-

frankel.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Feder:1999:CGP

[FHKM99] Tomas Feder, Pavol Hell, Sulamita Klein, and Rajeev Motwani. Complexity of graph partition problems. In ACM [ACM99], pages 464–

472. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p464-feder/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p464-
feder/p464-feder.pdf. ACM

order number 508990.

Fill:1997:IAP

[Fil97] James Allen Fill. An interruptible algorithm for perfect sampling via Markov chains. In ACM [ACM97], pages 688–

695. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p688-fill/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p688-
fill/p688-fill.pdf. ACM or-

der no. 508970.

Feige:1994:TPP

[FK94] Uri Feige and Joe Kilian.

Two prover protocols: low error at affordable rates. In ACM [ACM94], pages 172–

183. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p172-feige/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p172-
                                        feige/p172-feige.pdf. ACM

order no. 508930.

Feige:1995:IRR

[FK95a] Uriel Feige and Joe Kilian. Impossibility results for recycling random bits in twoprover proof systems. In ACM [ACM95], pages 457–

468. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p457-feige/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p457-
                                        feige/p457-feige.pdf. ACM

order no. 508950.

Fiat:1995:RMP

[FK95b] Amos Fiat and Anna R. Karlin. Randomized and multipointer paging with locality of reference. In ACM [ACM95], pages 626–634. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p626-fiat/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p626-
                                        fiat/p626-fiat.pdf. ACM or-

der no. 508950.

Farach:1996:EAI

[FK96] Martin Farach and Sampath Kannan. Efficient algorithms for inverting evolution. In ACM [ACM96], pages 230– 236. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p230-farach/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p230-farach/p230-farach.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Feige:1997:MGS

[FK97] Uriel Feige and Joe Kilian. Making games short (extended abstract). In ACM [ACM97], pages 506–

516. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p506-feige/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p506-
feige/p506-feige.pdf. ACM

order no. 508970.

Fournier:1998:LBE

[FK98] Hervé Fournier and Pascal Koiran. Are lower bounds easier over the reals? In ACM [ACM98], pages 507–

513. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p507-fournier/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p507-fournier/p507-
fournier.pdf. ACM order

number 508980.

Feige:1994:MMS

[FKN94] Uri Feige, Joe Killian, and Moni Naor. A minimal model for secure computation (extended abstract). In ACM [ACM94], pages 554–563. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p554-feige/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p554-
                                        feige/p554-feige.pdf. ACM

order no. 508930.

Freund:1993:ELT

[FKR^{+}93] Yoav Freund, Michael Kearns, Dana Ron, Ronitt Rubinfeld, Robert E. Schapire, and Linda Sellie. Efficient learning of typical finite automata from random walks. In ACM [ACM93], pages 315–

324. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p315-freund/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p315-freund/p315-freund.

pdf. ACM order no. 508930.

Feldmann:1993:OOS

[FKST93] Anja Feldmann, Ming-Yang Kao, Jiří Sgall, and Shang-Hua Teng. Optimal online scheduling of parallel jobs with dependencies. In ACM [ACM93], pages 642–651. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Farach:1993:RMF

[FKW93] Martin Farach, Sampath Kannan, and Tandy Warnow. A robust model for finding optimal evolutionary trees. In ACM [ACM93], pages 137–

145. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p137-farach/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p137-farach/p137-farach.

pdf. ACM order no. 508930.

Feige:1992:TPO

[FL92a] Uriel Feige and László Lovász.

Two-prover one-round proof systems: their power and their problems (extended abstract). In ACM [ACM92], pages 733–744. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p733-feige/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p733-
feige/p733-feige.pdf. ACM

order no. 508920.

Feige:1992:HCP

[FL92b] Uriel Feige and Carsten Lund.

On the hardness of computing the permanent of random matrices (extended abstract). In ACM [ACM92], pages 643–654. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p643-feige/;

                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p643-
                                        feige/p643-feige.pdf. ACM

order no. 508920.

Fleischer:1993:DTO

[Fle93] Rudolf Fleischer. Decision trees: old and new results. In ACM [ACM93], pages 468–477. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Feder:1991:CPG

[FM91] Tomás Feder and Rajeev Motwani. Clique partitions, graph compression and speeding-up algorithms. In ACM [ACM91], pages 123–

133. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p123-feder/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p123-
                                        feder/p123-feder.pdf. ACM

order no. 508910.

Feder:1992:BM

[FM92] Tomás Feder and Milena Mihail. Balanced matroids. In ACM [ACM92], pages 26–38. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p26-feder/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p26-
                                        feder/p26-feder.pdf. ACM

order no. 508920.

Ferragina:1999:MMD

[FMdB99] Paolo Ferragina, S. Muthukrishnan, and Mark de Berg. Multimethod dispatching: a geometric approach with applications to string matching problems. In ACM [ACM99], pages 483–491. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p483-ferragina/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p483-ferragina/p483-
ferragina.pdf. ACM order

number 508990.

Fischer:1990:WP

[FMRT90] M. J. Fischer, S. Moran, R. Rudich, and G. Taubenfeld. The wakeup problem. In ACM [ACM90], pages 106–116. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p106-fischer/.

ACM order no. 508900.

Frankel:1998:RED

[FMY98] Yair Frankel, Philip D. MacKenzie, and Moti Yung. Robust efficient distributed RSA-key generation. In ACM [ACM98], pages 663–672. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p663-frankel/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p663-frankel/p663-

frankel.pdf. ACM order num-

ber 508980.

Fiat:1991:RTS

[FN91] Amos Fiat and Moni Naor. Rigorous time/space tradeoffs for inverting functions. In ACM [ACM91], pages 534–

541. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p534-fiat/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p534-
                                        fiat/p534-fiat.pdf. ACM or-

der no. 508910.

Feige:1990:CUI

[FPRU90] U. Feige, D. Peleg, P. Raghavan, and E. Upfal. Computing with unreliable information. In ACM [ACM90], pages 128–137. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p128-feige/. ACM

order no. 508900.

Frederickson:1990:ITB

[Fre90]  G. N. Frederickson.  The

information theory bound is tight for selection in a heap. In ACM [ACM90], pages 26–33. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p26-frederickson/

. ACM order no. 508900.

Fredman:1998:ITI

[Fre98a]  Michael L. Fredman.  In-

formation theoretic implications for pairing heaps. In ACM [ACM98], pages 319– 326. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p319-fredman/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p319-fredman/p319-

fredman.pdf. ACM order number 508980.

Freedman:1998:SGU

[Fre98b] Michael H. Freedman. Ksat on groups and undecidability. In ACM [ACM98], pages 572–576. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p572-freedman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p572-freedman/p572-
freedman.pdf. ACM order

number 508980.

Friedman:1996:CBN

[Fri96]  Joel Friedman.  Comput-

ing Betti numbers via combinatorial Laplacians. In ACM [ACM96], pages 386– 391. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p386-friedman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        237814/p386-friedman/p386-
                                        friedman.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Feige:1990:WIW

[FS90] U. Feige and A. Shamir. Witness indistinguishable and witness hiding protocols. In ACM [ACM90], pages 416– 426. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p416-feige/. ACM

order no. 508900.

Fortnow:1997:RPC

[FS97] Lance Fortnow and Michael Sipser. Retraction of probabilistic computation and linear time. In ACM [ACM97], page 750. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p750-fortnow/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p750-fortnow/p750-

fortnow.pdf. ACM order no. 508970.

Feige:1998:IBA

[FS98] Uriel Feige and Christian Scheideler. Improved bounds for acyclic job shop scheduling (extended abstract). In ACM [ACM98], pages 624– 633. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/276698/p624-feige/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p624-
feige/p624-feige.pdf. ACM

order number 508980.

Freund:1997:UCP

[FSSW97] Yoav Freund, Robert E. Schapire, Yoram Singer, and Manfred K. Warmuth. Using and combining predictors that specialize. In ACM [ACM97], pages 334–343. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p334-freund/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p334-freund/p334-freund.

pdf. ACM order no. 508970.

Farach:1995:SML

[FT95] Martin Farach and Mikkel Thorup. String matching in Lempel–Ziv compressed strings. In ACM [ACM95], pages 703–

712. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p703-farach/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p703-farach/p703-farach.

pdf. ACM order no. 508950.

Fu:1996:MCF

[Fu96]  Xudong Fu.  Modular col-

oring formulas are hard for cutting planes proofs. In ACM [ACM96], pages 595–

602. ISBN 0-89791-785-5.

LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p595-fu/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/237814/p595-
fu/p595-fu.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Feder:1993:MMS

[FV93] Tomás Feder and Moshe Y. Vardi. Monotone monadic SNP and constraint satisfaction. In ACM [ACM93], pages 612–622. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Fredman:1990:BTI

[FW90] M. L. Fredman and D. E. Willard. BLASTING through the information theoretic barrier with FUSION TREES. In ACM [ACM90], pages 1–7. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p1-fredman/. ACM

order no. 508900.

Felsner:1993:MCP

[FW93] Stefan Felsner and Lorenz Wernisch. Maximum k-chains in planar point sets: combinatorial structure and algorithms. In ACM [ACM93], pages 146– 153. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p146-felsner/;

 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p146-felsner/p146-

felsner.pdf. ACM order no. 508930.

Fortnow:1994:ODR

[FW94] Lance Fortnow and Duke Whang. Optimality and domination in repeated games with bounded players. In ACM [ACM94], pages 741– 749. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p741-fortnow/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p741-fortnow/p741-

fortnow.pdf. ACM order no. 508930.

Franklin:1992:CCS

[FY92] Matthew Franklin and Moti Yung. Communication complexity of secure computation (extended abstract). In ACM [ACM92], pages 699– 710. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p699-franklin/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p699-franklin/p699-
franklin.pdf. ACM order no.

508920.

Franklin:1995:SHP

[FY95] Matthew Franklin and Moti Yung. Secure hypergraphs: privacy from partial broadcast. In ACM [ACM95], pages 36–44. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p36-franklin/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p36-franklin/p36-franklin.

pdf. ACM order no. 508950.

Gabow:1991:MAF

[Gab91] Harold N. Gabow. A matroid approach to finding edge connectivity and packing arborescences. In ACM [ACM91], pages 112–122. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p112-gabow/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p112-
                                        gabow/p112-gabow.pdf. ACM

order no. 508910.

Gabow:1994:ESA

[Gab94] Harold N. Gabow. Efficient splitting off algorithms for graphs. In ACM [ACM94], pages 696–705. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p696-gabow/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p696-
                                        gabow/p696-gabow.pdf. ACM

order no. 508930.

Gil:1990:AKC

[GadHW90] J. Gil, F. Meyer auf der Heide, and A. Wigderson. Not all keys can be hashed in constant time. In ACM [ACM90], pages 244–253. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p244-gil/. ACM

order no. 508900.

Galil:1992:CTO

[Gal92] Zvi Galil. A constant-time optimal parallel string-matching algorithm. In ACM [ACM92], pages 69–76. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p69-galil/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p69-
galil/p69-galil.pdf. ACM

order no. 508920.

Gal:1998:CSP

[Gál98] Anna Gál. A characterization of span program size and improved lower bounds for monotone span programs. In ACM [ACM98], pages 429–

437. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p429-gal/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/276698/p429-
gal/p429-gal.pdf. ACM order

number 508980.

Gamarnik:1999:SAN

[Gam99]  David Gamarnik.  Stability

of adaptive and non-adaptive packet routing policies in adversarial queueing networks. In ACM [ACM99], pages 206–

214. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p206-gamarnik/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p206-gamarnik/p206-
gamarnik.pdf. ACM order

number 508990.

Goldreich:1998:LNA

[GG98] Oded Goldreich and Shafi Goldwasser. On the limits of non-approximability of lattice problems. In ACM [ACM98], pages 1–9. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p1-goldreich/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p1-goldreich/p1-goldreich.

pdf. ACM order number 508980.

Garg:1994:AAM

[GH94] Naveen Garg and Dorit S. Hochbaum. An O(log k) approximation algorithm for the k minimum spanning tree problem in the plane. In ACM [ACM94], pages 432–

438. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p432-garg/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p432-
garg/p432-garg.pdf. ACM or-

der no. 508930.

Goldmann:1995:MCC

[GH95] Mikael Goldmann and Johan Håstad. Monotone circuits for connectivity have depth (log n)^{2−o(1)} (extended abstract). In ACM [ACM95], pages 569–574. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p569-goldmann/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p569-goldmann/p569-
goldmann.pdf. ACM order no.

508950.

Grove:1992:ACP

[GHK92] Adam J. Grove, Joseph Y.

Halpern, and Daphne Koller. Asymptomatic conditional probabilities for first-order logic. In ACM [ACM92], pages 294–305. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p294-grove/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p294-
grove/p294-grove.pdf. ACM

order no. 508920.

Goel:1999:SDT

[GHPT99] Ashish Goel, Monika R. Henzinger, Serge Plotkin, and Eva Tardos. Scheduling data transfers in a network and the set scheduling problem. In ACM [ACM99], pages 189–

197. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/

                                        pubs/citations/proceedings/
                                        stoc/301250/p189-goel/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p189-
                                        goel/p189-goel.pdf. ACM or-

der number 508990.

Galil:1991:FDA

[GI91] Zvi Galil and Giuseppe F. Italiano. Fully dynamic algorithms for edge connectivity problems. In ACM [ACM91], pages 317–327. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p317-galil/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p317-
                                        galil/p317-galil.pdf. ACM

order no. 508910.

Gertner:1998:PDP

[GIKM98] Yael Gertner, Yuval Ishai, Eyal Kushilevitz, and Tal Malkin. Protecting data privacy in private information retrieval schemes. In ACM [ACM98], pages 151–

160. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p151-gertner/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p151-gertner/p151-

gertner.pdf. ACM order num-

ber 508980.

Galil:1992:FDP

[GIS92] Zvi Galil, Giuseppe F. Italiano, and Neil Sarnak. Fully dynamic planarity testing (extended abstract). In ACM [ACM92], pages 495–506. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p495-galil/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p495-
galil/p495-galil.pdf. ACM

order no. 508920.

Gore:1997:SWP

[GJ97] Vivek K. Gore and Mark R.

Jerrum. The Swendsen–Wang process does not always mix rapidly. In ACM [ACM97], pages 674–681. ISBN 0-89791-888-6. LCCN QA76.5 .A849

1997. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/258533/p674-gore/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p674-
gore/p674-gore.pdf. ACM or-

der no. 508970.

Goldmann:1993:STC

[GK93] Mikael Goldmann and Marek Karpinski. Simulating threshold circuits by majority circuits. In ACM [ACM93], pages 551–560. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Grigoriev:1997:RLB

[GK97] Dima Grigoriev and Marek Karpinski. Randomized Ω(n^{2}) lower bound for knapsack. In ACM [ACM97], pages 76–

85. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/

stoc/258533/p76-grigoriev/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p76-grigoriev/p76-
grigoriev.pdf. ACM order no.

508970.

Grigoriev:1998:ELB

[GK98] Dima Grigoriev and Marek Karpinski. An exponential lower bound for depth 3 arithmetic circuits. In ACM [ACM98], pages 577– 582. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p577-grigoriev/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p577-grigoriev/p577-
grigoriev.pdf. ACM order

number 508980.

Grigoriev:1996:LBR

[GKadHS96] Dima Grigoriev, Marek Karpinski, Friedhelm Meyer auf der Heide, and Roman Smolensky. A lower bound for randomized algebraic decision trees. In ACM [ACM96], pages 612– 619. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p612-grigoriev/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p612-grigoriev/p612-
grigoriev.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Guruswami:1999:NOH

[GKR^{+}99] Venkatesan Guruswami, Sanjeev Khanna, Rajmohan Rajaraman, Bruce Shepherd, and Mihalis Yannakakis. Nearoptimal hardness results and approximation algorithms for edge-disjoint paths and related problems. In ACM [ACM99], pages 19–28. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p19-guruswami/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p19-guruswami/p19-
guruswami.pdf. ACM order

number 508990.

Goddard:1990:ORA

[GKS90] W. Goddard, V. King, and L. Schulman. Optimal randomized algorithms for local sorting and set-maxima. In ACM [ACM90], pages 45–53. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p45-goddard/.

ACM order no. 508900.

Gabow:1999:UMM

[GKT99] Harold N. Gabow, Haim Kaplan, and Robert E. Tarjan. Unique maximum matching algorithms. In ACM [ACM99], pages 70–78. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p70-gabow/;

http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p70-
gabow/p70-gabow.pdf. ACM

order number 508990.

Grigoriev:1994:LBT

[GKV94] Dima Grigoriev, Marek Karpinski, and Nicolai Vorobjov. Lower bounds on testing membership to a polyhedron by algebraic decision trees. In ACM [ACM94], pages 635–

644. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p635-grigoriev/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p635-grigoriev/p635-
grigoriev.pdf. ACM order no.

508930.

Ghosh:1995:TAT

[GLM^{+}95] Bhaskar Ghosh, F. T. Leighton, Bruce M. Maggs, S. Muthukrishnan, C. Greg Plaxton, R. Rajaraman, Andréa W. Richa, Robert E. Tarjan, and David Zuckerman. Tight analyses of two local load balancing algorithms. In ACM [ACM95], pages 548–

558. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p548-ghosh/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p548-
ghosh/p548-ghosh.pdf. ACM

order no. 508950.

Gemmell:1991:STC

[GLR^{+}91] Peter Gemmell, Richard Lipton, Ronitt Rubinfeld, Madhu Sudan, and Avi Wigderson. Self-testing/correcting for polynomials and for approximate functions. In ACM [ACM91], pages 33–42. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p33-gemmell/.

ACM order no. 508910.

Garay:1993:FPB

[GM93] Juan A. Garay and Yoram Moses. Fully polynomial Byzantine agreement in t + 1 rounds. In ACM [ACM93], pages 31–41. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p31-garay/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p31-
garay/p31-garay.pdf. ACM

order no. 508930.

Gradel:1995:DCT

[GM95] Erich Grädel and Klaus Meer.

Descriptive complexity theory over the real numbers. In ACM [ACM95], pages 315–

324. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p315-gradel/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p315-gradel/p315-gradel.

pdf. ACM order no. 508950.

Guha:1999:ERP

[GMNS99] Sudipto Guha, Anna Moss, Joseph (Seffi) Naor, and Baruch Schieber. Efficient recovery from power outage (extended abstract). In ACM [ACM99], pages 574–

582. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p574-guha/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p574-
guha/p574-guha.pdf. ACM or-

der number 508990.

Goldberg:1993:PSP

[Gol93] Leslie Ann Goldberg. Polynomial space polynomial delay algorithms for listing families of graphs. In ACM [ACM93], pages 218–225. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p218-goldberg/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p218-goldberg/p218-
goldberg.pdf. ACM order no.

508930.

Goodrich:1992:PSP

[Goo92] Michael T. Goodrich. Planar separators and parallel polygon triangulation (preliminary version). In ACM [ACM92], pages 507–516. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p507-goodrich/

;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p507-goodrich/p507-
goodrich.pdf. ACM order no.

508920.

Goodrich:1996:CEP

[Goo96] Michael T. Goodrich. Communicationefficient parallel sorting (preliminary version). In ACM [ACM96], pages 247–256. ISBN 0-89791-785-5. LCCN QA 76.6

A13 1996. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/237814/p247-goodrich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p247-goodrich/p247-
goodrich.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Goldreich:1994:CCK

[GOP94] Oded Goldreich, Rafail Ostrovsky, and Erez Petrank. Computational complexity and knowledge complexity (extended abstract). In ACM [ACM94], pages 534–543. ISBN 0-89791-663-8. LCCN QA76 .A15 1994. URL http://www.

acm.org/pubs/citations/proceedings/
stoc/195058/p534-goldreich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p534-goldreich/p534-
goldreich.pdf. ACM order no.

508930.

Goldreich:1997:PTB

[GR97] Oded Goldreich and Dana Ron. Property testing in bounded degree graphs. In ACM [ACM97], pages 406– 415. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p406-goldreich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p406-goldreich/p406-
goldreich.pdf. ACM order no.

508970.

Goldreich:1998:SBT

[GR98] Oded Goldreich and Dana Ron. A sublinear bipartiteness tester for bounded degree graphs. In ACM [ACM98], pages 289– 298. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p289-goldreich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p289-goldreich/p289-
goldreich.pdf. ACM order

number 508980.

Gal:1999:TSA

[GR99] Anna Gál and Adi Rosén. A theorem on sensitivity and applications in private computation. In ACM [ACM99], pages 348–357. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p348-gal/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/301250/p348-
gal/p348-gal.pdf. ACM order

number 508990.

Grigoriev:1998:RCL

[Gri98]  D. Grigoriev.  Randomized

complexity lower bounds. In ACM [ACM98], pages 219–

223. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p219-grigoriev/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p219-grigoriev/p219-
grigoriev.pdf. ACM order

number 508980.

Grove:1991:HOS

[Gro91]  E. F. Grove.  The har-

monic online K-server algorithm is competitive. In ACM [ACM91], pages 260–

266. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p260-grove/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p260-
grove/p260-grove.pdf. ACM

order no. 508910.

Grolmusz:1994:WST

[Gro94] Vince Grolmusz. A weight-size trade-off for circuits with MOD m gates. In ACM [ACM94], pages 68–74. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p68-grolmusz/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p68-grolmusz/p68-grolmusz.

pdf. ACM order no. 508930.

Grover:1996:FQM

[Gro96]  Lov K. Grover.  A fast

quantum mechanical algorithm for database search. In ACM [ACM96], pages 212–

219. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p212-grover/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p212-grover/p212-grover.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Grover:1998:FFQ

[Gro98] Lov K. Grover. A framework for fast quantum mechanical algorithms. In ACM [ACM98], pages 53–62. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p53-grover/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p53-
                                        grover/p53-grover.pdf. ACM

order number 508980.

Goldreich:1999:CRE

[GRS99] Oded Goldreich, Dana Ron, and Madhu Sudan. Chinese remaindering with errors. In ACM [ACM99], pages 225–234. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p225-goldreich/
                                        ;   http://www.acm.org/pubs/

articles/proceedings/stoc/
301250/p225-goldreich/p225-
goldreich.pdf. ACM order

number 508990.

Goldreich:1998:HVS

[GSV98] Oded Goldreich, Amit Sahai, and Salil Vadhan. Honestverifier statistical zero-knowledge equals general statistical zeroknowledge. In ACM [ACM98], pages 399–408. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p399-goldreich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p399-goldreich/p399-
goldreich.pdf. ACM order

number 508980.

Goodrich:1991:DTD

[GT91] Michael T. Goodrich and Roberto Tamassia. Dynamic trees and dynamic point location. In ACM [ACM91], pages 523–533. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p523-goodrich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p523-goodrich/p523-
goodrich.pdf. ACM order no.

508910.

Gupta:1999:ETM

[Gup99] Anupam Gupta. Embedding tree metrics into low dimensional Euclidean spaces. In ACM [ACM99], pages 694–

700. ISBN 1-58113-067-8.

LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p694-gupta/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p694-
                                        gupta/p694-gupta.pdf. ACM

order number 508990.

Garg:1993:AMF

[GVY93] Naveen Garg, Vijay V. Vazirani, and Mihalis Yannakakis. Approximate max-flow min- (multi)cut theorems and their applications. In ACM [ACM93], pages 698–707. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Goemans:1994:AAM

[GW94a] Michel X. Goemans and David P. Williamson. .879-approximation algorithms for MAX CUT and MAX 2SAT. In ACM [ACM94], pages 422–431. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p422-goemans/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p422-goemans/p422-

goemans.pdf. ACM order no.

508930.

Goldreich:1994:TFF

[GW94b] Oded Goldreich and Avi Wigderson. Tiny families of functions with random properties (preliminary version): a quality-size trade-off for hashing. In ACM [ACM94], pages 574–584. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p574-goldreich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p574-goldreich/p574-
goldreich.pdf. ACM order no.

508930.

Galil:1995:SLV

[GY95] Zvi Galil and Xiangdong Yu. Short length versions of Menger’s theorem. In ACM [ACM95], pages 499–

508. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p499-galil/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p499-
galil/p499-galil.pdf. ACM

order no. 508950.

Hagerup:1991:CTP

[Hag91] Torben Hagerup. Constanttime parallel integer sorting. In ACM [ACM91], pages 299–

306. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p299-hagerup/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p299-hagerup/p299-

hagerup.pdf. ACM order no. 508910.

Hagerup:1994:OPS

[Hag94] Torben Hagerup. Optimal parallel string algorithms: sorting, merging and computing the minimum. In ACM [ACM94], pages 382–391. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p382-hagerup/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p382-hagerup/p382-

hagerup.pdf. ACM order no.

508930.

Harel:1991:HPI

[Har91]  David Harel.  Hamiltonian

paths in infinite graphs. In ACM [ACM91], pages 220–

229. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p220-harel/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p220-
                                        harel/p220-harel.pdf. ACM

order no. 508910.

Hariharan:1994:OPS

[Har94] Ramesh Hariharan. Optimal parallel suffix tree construction. In ACM [ACM94], pages 290–299. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p290-hariharan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p290-hariharan/p290-
                                        hariharan.pdf. ACM order no.

508930.

Haastad:1990:PRG

[Hås90] J. Håstad. Pseudo-random generators under uniform assumptions. In ACM [ACM90], pages 395–404. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p395-hastad/.

ACM order no. 508900.

Haastad:1996:TLC

[Hås96] Johan Håstad. Testing of the long code and hardness for clique. In ACM [ACM96], pages 11–19. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p11-hastad/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p11-
hastad/p11-hastad.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Haastad:1997:SOI

[Hås97] Johan Håstad. Some optimal inapproximability results. In ACM [ACM97], pages 1–

10. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p1-hastad/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p1-
hastad/p1-hastad.pdf. ACM

order no. 508970.

Harchol-Balter:1995:BDP

[HBW95] Mor Harchol-Balter and David Wolfe. Bounding delays in packet-routing networks. In ACM [ACM95], pages 248–

257. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p248-harchol-
                                        balter/;   http://www.acm.
                                        org/pubs/articles/proceedings/
                                        stoc/225058/p248-harchol-
                                        balter/p248-harchol-balter.

pdf. ACM order no. 508950.

Holm:1998:PLD

[HdLT98] Jacob Holm, Kristian de Lichtenberg, and Mikkel Thorup. Poly-logarithmic deterministic fully-dynamic algorithms for connectivity, minimum spanning tree, 2-edge, and biconnectivity. In ACM [ACM98], pages 79–89. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p79-holm/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p79-
                                        holm/p79-holm.pdf. ACM or-

der number 508980.

He:1997:FPG

[He97] Xin He. On floorplans of planar graphs. In ACM [ACM97], pages 426–435. ISBN 0-89791-888-6. LCCN QA76.5 .A849

                                        1997. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/258533/p426-he/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/258533/p426-
                                        he/p426-he.pdf. ACM order

no. 508970.

Hales:1999:QFS

[HH99] Lisa Hales and Sean Hallgren. Quantum Fourier sampling simplified. In ACM [ACM99], pages 330–338. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p330-hales/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p330-
hales/p330-hales.pdf. ACM

order number 508990.

Hart:1995:FPF

[HI95] William E. Hart and Sorin Istrail. Fast protein folding in the hydrophobic-hydrophilic model within three-eights of optimal. In ACM [ACM95], pages 157–168. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p157-hart/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p157-
hart/p157-hart.pdf. ACM or-

der no. 508950.

Harju:1990:DME

[HK90] T. Harju and J. Karhumäki.

Decidability of the multiplicity equivalence of multitape finite automata. In ACM [ACM90], pages 477–481. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p477-harju/. ACM

order no. 508900.

Henzinger:1995:RDG

[HK95] Monika Rauch Henzinger and Valerie King. Randomized dynamic graph algorithms with polylogarithmic time per operation. In ACM [ACM95], pages 519–527. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p519-henzinger/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p519-henzinger/p519-
henzinger.pdf. ACM order no.

508950.

Henzinger:1995:WDA

[HKPV95] Thomas A. Henzinger, Peter W. Kopke, Anuj Puri, and Pravin Varaiya. What’s decidable about hybrid automata? In ACM [ACM95], pages 373– 382. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p373-henzinger/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p373-henzinger/p373-
henzinger.pdf. ACM order no.

508950.

Hannenhalli:1995:TCT

[HP95] Sridhar Hannenhalli and Pavel Pevzner. Transforming cabbage into turnip: polynomial algorithm for sorting signed permutations by reversals. In ACM [ACM95], pages 178– 189. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p178-hannenhalli/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p178-hannenhalli/p178-

hannenhalli.pdf. ACM order

no. 508950.

Halldorsson:1994:GGA

[HR94] Magnús Halldórsson and Jaikumar Radhakrishnan. Greed is good: approximating independent sets in sparse and bounded-degree graphs. In ACM [ACM94], pages 439– 448. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p439-halldorsson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p439-halldorsson/p439-
halldorsson.pdf. ACM order

no. 508930.

Herlihy:1997:DDD

[HR97] Maurice Herlihy and Sergio Rajsbaum. The decidability of distributed decision tasks (extended abstract). In ACM [ACM97], pages 589– 598. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p589-herlihy/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p589-herlihy/p589-

herlihy.pdf. ACM order no. 508970.

Huang:1998:BBA

[HR98] Ming-Deh A. Huang and Ashwin J. Rao. A black box approach to the algebraic set decomposition problem. In ACM [ACM98], pages 497– 506. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p497-huang/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p497-
huang/p497-huang.pdf. ACM

order number 508980.

Hellerstein:1995:HMQ

[HRPW95] Lisa Hellerstein, Vijay Raghavan, Krishnan Pillaipakkamnatt, and Dawn Wilkins. How many queries are needed to learn? In ACM [ACM95], pages 190–199. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p190-hellerstein/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p190-hellerstein/p190-
hellerstein.pdf. ACM order

no. 508950.

Herlihy:1993:ACT

[HS93a] Maurice Herlihy and Nir Shavit. The asynchronous computability theorem for t-resilient tasks. In ACM [ACM93], pages 111–

120. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p111-herlihy/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p111-herlihy/p111-

herlihy.pdf. ACM order no. 508930.

Hershberger:1993:MSS

[HS93b] John Hershberger and Subhash Suri. Matrix searching with the shortest path metric. In ACM [ACM93], pages 485–494. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Herlihy:1994:SCC

[HS94] Maurice Herlihy and Nir Shavit. A simple constructive computability theorem for wait-free computation. In ACM [ACM94], pages 243– 252. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p243-herlihy/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p243-herlihy/p243-

herlihy.pdf. ACM order no. 508930.

Hromkovic:1996:NCL

[HS96] J. Hromkovic and G. Schnitger. Nondeterministic communication with a limited number of advice bits. In ACM [ACM96], pages 551– 560. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p551-hromkovic/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p551-hromkovic/p551-
hromkovic.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Huber:1998:ESA

[Hub98] Mark Huber. Exact sampling and approximate counting techniques. In ACM [ACM98], pages 31–40. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p31-huber/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p31-
huber/p31-huber.pdf. ACM

order number 508980.

Ishai:1999:IUB

[IK99] Yuval Ishai and Eyal Kushilevitz. Improved upper bounds on information-theoretic private information retrieval (extended abstract). In ACM [ACM99], pages 79–88. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p79-ishai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p79-
ishai/p79-ishai.pdf. ACM

order number 508990.

Indyk:1998:ANN

[IM98] Piotr Indyk and Rajeev Motwani. Approximate nearest neighbors: towards removing the curse of dimensionality. In ACM [ACM98], pages 604–613. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p604-indyk/;

http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p604-
indyk/p604-indyk.pdf. ACM

order number 508980.

Indyk:1997:LPH

[IMRV97] Piotr Indyk, Rajeev Motwani, Prabhakar Raghavan, and Santosh Vempala. Localitypreserving hashing in multidimensional spaces. In ACM [ACM97], pages 618–

625. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p618-indyk/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p618-
indyk/p618-indyk.pdf. ACM

order no. 508970.

Indyk:1999:ISF

[Ind99a]  Piotr Indyk.  Interpolation

of symmetric functions and a new type of combinatorial design. In ACM [ACM99], pages 736–740. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p736-indyk/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p736-
indyk/p736-indyk.pdf. ACM

order number 508990.

Indyk:1999:STA

[Ind99b] Piotr Indyk. Sublinear time algorithms for metric space problems. In ACM [ACM99], pages 428–434. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/

                                        pubs/citations/proceedings/
                                        stoc/301250/p428-indyk/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p428-
                                        indyk/p428-indyk.pdf. ACM

order number 508990.

Impagliazzo:1994:PNA

[INW94] Russell Impagliazzo, Noam Nisan, and Avi Wigderson. Pseudorandomness for network algorithms. In ACM [ACM94], pages 356–364. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p356-impagliazzo/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p356-impagliazzo/p356-
                                        impagliazzo.pdf. ACM order

no. 508930.

Impagliazzo:1993:SDT

[IPS93] Russell Impagliazzo, Ramamohan Paturi, and Michael E. Saks. Size-depth trade-offs for threshold circuits. In ACM [ACM93], pages 541–550. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Irani:1997:PRM

[Ira97] Sandy Irani. Page replacement with multi-size pages and applications to Web caching. In ACM [ACM97], pages 701–

710. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p701-irani/;
                                        http://www.acm.org/pubs/articles/

proceedings/stoc/258533/p701-
irani/p701-irani.pdf. ACM

order no. 508970.

Idury:1993:MMR

[IS93] Ramana M. Idury and Alejandro A. Schäffer. Multiple matching of rectangular patterns. In ACM [ACM93], pages 81–90. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p81-idury/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p81-
idury/p81-idury.pdf. ACM

order no. 508930.

Impagliazzo:1997:PBI

[IW97] Russell Impagliazzo and Avi Wigderson. P = BPP if E requires exponential circuits: derandomizing the XOR lemma. In ACM [ACM97], pages 220–

229. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p220-impagliazzo/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p220-impagliazzo/p220-
impagliazzo.pdf. ACM order

no. 508970.

Jiang:1993:OWH

[JL93] Tao Jiang and Ming Li. k oneway heads cannot do stringmatching. In ACM [ACM93], pages 62–70. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/167088/p62-jiang/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/167088/p62-
                                        jiang/p62-jiang.pdf. ACM

order no. 508930.

Jiang:1994:ASE

[JLW94] Tao Jiang, Eugene L. Lawler, and Lusheng Wang. Aligning sequences via an evolutionary tree: complexity and approximation. In ACM [ACM94], pages 760–769. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p760-jiang/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p760-
                                        jiang/p760-jiang.pdf. ACM

order no. 508930.

Jones:1993:CTF

[Jon93] Neil D. Jones. Constant time factors do matter. In ACM [ACM93], pages 602–611. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Jansen:1999:IAS

[JP99] Klaus Jansen and Lorant Porkolab. Improved approximation schemes for scheduling unrelated parallel machines. In ACM [ACM99], pages 408–

417. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p408-jansen/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p408-jansen/p408-jansen.

pdf. ACM order number 508990.

Jakoby:1994:CCW

[JRS94] Andreas Jakoby, Rüdiger Reischuk, and Christian Schindelhauer. Circuit complexity: from the worst case to the average case. In ACM [ACM94], pages 58–67. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p58-jakoby/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p58-
jakoby/p58-jakoby.pdf. ACM

order no. 508930.

Jacquet:1994:FEO

[JS94] Philippe Jacquet and Wojciech Szpankowski. A functional equation often arising in the analysis of algorithms (extended abstract). In ACM [ACM94], pages 780–789. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p780-jacquet/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p780-jacquet/p780-

jacquet.pdf. ACM order no. 508930.

Jansen:1999:MMJ

[JSOS99] Klaus Jansen, Roberto Solis-Oba, and Maxim Sviridenko. Makespan minimization in job shops: a polynomial time approximation scheme. In ACM [ACM99], pages 394–

399. ISBN 1-58113-067-8.

LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p394-jansen/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p394-jansen/p394-jansen.

pdf. ACM order number

508990.

Jiang:1994:THB

[JSV94] Tao Jiang, Joel I. Seiferas, and Paul M. B. Vitányi. Two heads are better than two tapes. In ACM [ACM94], pages 668–675. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p668-jiang/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p668-
                                        jiang/p668-jiang.pdf. ACM

order no. 508930.

Kahan:1991:MDM

[Kah91]  Simon Kahan.  A model

for data in motion. In ACM [ACM91], pages 265–

277. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p265-kahan/. ACM

order no. 508910.

Kaltofen:1991:ENI

[Kal91]  Erich Kaltofen.  Effective

Noether irreducibility forms and applications. In ACM [ACM91], pages 54–63. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/103418/p54-kaltofen/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p54-kaltofen/p54-kaltofen.

pdf. ACM order no. 508910.

Kalai:1992:SRS

[Kal92] Gil Kalai. A subexponential randomized simplex algorithm (extended abstract). In ACM [ACM92], pages 475–

482. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p475-kalai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p475-
kalai/p475-kalai.pdf. ACM

order no. 508920.

Kapoor:1999:ECG

[Kap99] Sanjiv Kapoor. Efficient computation of geodesic shortest paths. In ACM [ACM99], pages 770–779. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p770-kapoor/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p770-kapoor/p770-kapoor.

pdf. ACM order number 508990.

Karp:1991:PRR

[Kar91] Richard M. Karp. Probabilistic recurrence relations. In ACM [ACM91], pages 190–

197. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/103418/p190-karp/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p190-
                                        karp/p190-karp.pdf. ACM or-

der no. 508910.

Karp:1993:MGS

[Kar93]  Richard M. Karp.  Map-

ping the genome: some combinatorial problems arising in molecular biology. In ACM [ACM93], pages 278–

285. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p278-karp/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/167088/p278-
                                        karp/p278-karp.pdf. ACM or-

der no. 508930.

Karger:1994:RSC

[Kar94] David R. Karger. Random sampling in cut, flow, and network design problems. In ACM [ACM94], pages 648–

657. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p648-karger/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p648-karger/p648-karger.

pdf. ACM order no. 508930.

Karger:1995:RFP

[Kar95] David R. Karger. A randomized fully polynomial time approximation scheme for the all terminal network reliability problem. In ACM [ACM95], pages 11–17. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p11-karger/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p11-
karger/p11-karger.pdf. ACM

order no. 508950.

Karger:1996:MCN

[Kar96a] David R. Karger. Minimum cuts in near-linear time. In ACM [ACM96], pages 56–

63. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p56-karger/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p56-
karger/p56-karger.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Karloff:1996:HGG

[Kar96b] Howard Karloff. How good is the Goemans-Williamson MAX CUT algorithm? In ACM [ACM96], pages 427–

434. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p427-karloff/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p427-karloff/p427-

karloff.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Karger:1997:URS

[Kar97]  David R. Karger.  Us-

ing random sampling to find maximum flows in uncapacitated undirected graphs. In ACM [ACM97], pages 240–

249. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p240-karger/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p240-karger/p240-karger.

pdf. ACM order no. 508970.

Kosaraju:1995:LSA

[KD95] S. Rao Kosaraju and Arthur L. Delcher. Large-scale assembly of DNA strings and spaceefficient construction of suffix trees. In ACM [ACM95], pages 169–177. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p169-kosaraju/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p169-kosaraju/p169-
                                        kosaraju.pdf. ACM order no.

508950.

Kosaraju:1996:LSA

[KD96] S. Rao Kosaraju and Arthur L. Delcher. Large-scale assembly of DNA strings and spaceefficient construction of suffix trees. In ACM [ACM96], page 659. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/237814/p659-kosaraju/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p659-kosaraju/p659-
kosaraju.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Kearns:1993:ENT

[Kea93] Michael Kearns. Efficient noisetolerant learning from statistical queries. In ACM [ACM93], pages 392–401. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p392-kearns/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p392-kearns/p392-kearns.

pdf. ACM order no. 508930.

Kelsen:1992:PCC

[Kel92] Pierre Kelsen. On the parallel complexity of computing a maximal independent set in a hypergraph. In ACM [ACM92], pages 339–350. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p339-kelsen/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p339-kelsen/p339-kelsen.

pdf. ACM order no. 508920.

Kharitonov:1993:CHD

[Kha93] Michael Kharitonov. Cryptographic hardness of distributionspecific learning. In ACM [ACM93], pages 372–381. ISBN 0-89791-591-7. LCCN QA 76.6

                                        A13 1993. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/167088/p372-kharitonov/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p372-kharitonov/p372-
                                        kharitonov.pdf. ACM order

no. 508930.

Kilian:1991:GCT

[Kil91] Joe Kilian. A general completeness theorem for two party games. In ACM [ACM91], pages 553–560. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p553-kilian/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p553-kilian/p553-kilian.

pdf. ACM order no. 508910.

Kilian:1992:NEZ

[Kil92] Joe Kilian. A note on efficient zero-knowledge proofs and arguments (extended abstract). In ACM [ACM92], pages 723–732. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p723-kilian/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p723-kilian/p723-kilian.

pdf. ACM order no. 508920.

Kao:1990:TOT

[KK90] M.-Y. Kao and P. N. Klein. Towards overcoming the transitiveclosure bottleneck: efficient parallel algorithms for planar digraphs. In ACM [ACM90], pages 181–192. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p181-kao/. ACM

order no. 508900.

Kahn:1992:ES

[KK92] Jeff Kahn and Jeong Han Kim. Entropy and sorting. In ACM [ACM92], pages 178–187. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p178-kahn/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p178-
kahn/p178-kahn.pdf. ACM or-

der no. 508920.

Karger:1999:RAG

[KKS^{+}99] David R. Karger, Philip Klein, Cliff Stein, Mikkel Thorup, and Neal E. Young. Rounding algorithms for a geometric embedding of minimum multiway cut. In ACM [ACM99], pages 668–678. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p668-karger/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p668-karger/p668-karger.

pdf. ACM order number 508990.

Kantor:1990:CQG

[KL90] W. M. Kantor and E. M.

Luks. Computing in quotient groups. In ACM [ACM90], pages 524–534. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p524-kantor/.

ACM order no. 508900.

Klein:1996:EAA

[KL96] Philip Klein and Hsueh-I. Lu. Efficient approximation algorithms for semidefinite programs arising from MAX CUT and COLORING. In ACM [ACM96], pages 338–

347. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p338-klein/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p338-
klein/p338-klein.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Karger:1998:FMF

[KL98a] David R. Karger and Matthew S. Levine. Finding maximum flows in undirected graphs seems easier than bipartite matching. In ACM [ACM98], pages 69–78. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p69-karger/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p69-
karger/p69-karger.pdf. ACM

order number 508980.

Khanna:1998:BDP

[KL98b] Sanjeev Khanna and Vincenzo Liberatore. On broadcast disk paging. In ACM [ACM98], pages 634–643. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p634-khanna/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p634-khanna/p634-khanna.

pdf. ACM order number 508980.

Karp:1992:EPS

[KLadH92] Richard M. Karp, Michael Luby, and Friedhelm Meyer auf der Heide. Efficient PRAM simulation on a distributed memory machine. In ACM [ACM92], pages 318–

326. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p318-karp/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p318-
karp/p318-karp.pdf. ACM or-

der no. 508920.

Kleinberg:1997:TAN

[Kle97] Jon M. Kleinberg. Two algorithms for nearest-neighbor search in high dimensions. In ACM [ACM97], pages 599–

608. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p599-kleinberg/
;   http://www.acm.org/pubs/

                                        articles/proceedings/stoc/
                                        258533/p599-kleinberg/p599-
                                        kleinberg.pdf. ACM order no.

508970.

Kleinberg:1998:DAU

[Kle98] Jon M. Kleinberg. Decision algorithms for unsplittable flow and the half-disjoint paths problem. In ACM [ACM98], pages 530–539. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/276698/p530-kleinberg/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        276698/p530-kleinberg/p530-
                                        kleinberg.pdf. ACM order

number 508980.

Karger:1997:CHR

[KLL^{+}97] David Karger, Eric Lehman, Tom Leighton, Rina Panigrahy, Matthew Levine, and Daniel Lewin. Consistent hashing and random trees: distributed caching protocols for relieving hot spots on the World Wide Web. In ACM [ACM97], pages 654–663. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p654-karger/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p654-karger/p654-karger.

pdf. ACM order no. 508970.

Kahale:1995:LBS

[KLM^{+}95] Nabil Kahale, Tom Leighton, Yuan Ma, C. Greg Plaxton, Torsten Suel, and Endre Szemerédi. Lower bounds for sorting networks. In ACM [ACM95], pages 437–

446. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p437-kahale/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p437-kahale/p437-kahale.

pdf. ACM order no. 508950.

Kavraki:1995:RQP

[KLMR95] Lydia E. Kavraki, Jean-Claude Latombe, Rajeev Motwani, and Prabhakar Raghavan. Randomized query processing in robot path planning. In ACM [ACM95], pages 353–

362. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p353-kavraki/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p353-kavraki/p353-

kavraki.pdf. ACM order no. 508950.

Kushilevitz:1996:LAC

[KLO96] Eyal Kushilevitz, Nathan Linial, and Rafail Ostrovsky. The linear-array conjecture in communication complexity is false. In ACM [ACM96], pages 1–

10. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p1-kushilevitz/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        237814/p1-kushilevitz/p1-
                                        kushilevitz.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Kao:1997:GTC

[KLP^{+}97] Ming-Yang Kao, Tak-Wah Lam, Teresa M. Przytycka, Wing-Kin Sung, and Hing-Fung Ting. General techniques for comparing unrooted evolutionary trees. In ACM [ACM97], pages 54–65. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p54-kao/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/258533/p54-
                                        kao/p54-kao.pdf. ACM order

no. 508970.

Kushilevitz:1991:LDT

[KM91] Eyal Kushilevitz and Yishay Mansour. Learning decision trees using the Fourier spectrum. In ACM [ACM91], pages 455–464. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p455-kushilevitz/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p455-kushilevitz/p455-
                                        kushilevitz.pdf. ACM order

no. 508910.

Koller:1993:CSS

[KM93] Daphne Koller and Nimrod Megiddo. Constructing small sample spaces satisfying given constraints. In ACM [ACM93], pages 268–

277. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p268-koller/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p268-koller/p268-koller.

pdf. ACM order no. 508930.

Karger:1994:DTA

[KM94a] David R. Karger and Rajeev Motwani. Derandomization through approximation: an NC algorithm for minimum cuts. In ACM [ACM94], pages 497–506. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p497-karger/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p497-karger/p497-karger.

pdf. ACM order no. 508930.

Karloff:1994:CWI

[KM94b] Howard Karloff and Yishay Mansour. On construction of kwise independent random variables. In ACM [ACM94], pages 564–573. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p564-karloff/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p564-karloff/p564-

karloff.pdf. ACM order no. 508930.

Karpinski:1995:PBV

[KM95] Marek Karpinski and Angus Macintyre. Polynomial bounds for VC dimension of sigmoidal neural networks. In ACM [ACM95], pages 200–

208. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p200-karpinski/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p200-karpinski/p200-
karpinski.pdf. ACM order no.

508950.

Kearns:1996:BAT

[KM96a] Michael Kearns and Yishay Mansour. On the boosting ability of top-down decision tree learning algorithms. In ACM [ACM96], pages 459–

468. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p459-kearns/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p459-kearns/p459-kearns.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Khanna:1996:TSC

[KM96b] Sanjeev Khanna and Rajeev Motwani. Towards a syntactic characterization of PTAS. In ACM [ACM96], pages 329–

337. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p329-khanna/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p329-khanna/p329-khanna.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Kearns:1994:LDD

[KMR^{+}94] Michael Kearns, Yishay Mansour, Dana Ron, Ronitt Rubinfeld, Robert E. Schapire, and Linda Sellie. On the learnability of discrete distributions. In ACM [ACM94], pages 273–282. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p273-kearns/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p273-kearns/p273-kearns.

pdf. ACM order no. 508930.

Kushilevitz:1993:LBR

[KMRZ93] Eyal Kushilevitz, Yishay Mansour, Michael O. Rabin, and David Zuckerman. Lower bounds for randomized mutual exclusion. In ACM [ACM93], pages 154–163. ISBN 0-89791-591-7. LCCN QA 76.6 A13

1993. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/167088/p154-kushilevitz/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p154-kushilevitz/p154-

                                        kushilevitz.pdf. ACM order

no. 508930.

Kiwi:1999:ATR

[KMS99] Marcos Kiwi, Frédéric Magniez, and Miklos Santha. Approximate testing with relative error. In ACM [ACM99], pages 51–60. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p51-kiwi/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p51-
                                        kiwi/p51-kiwi.pdf. ACM or-

der number 508990.

Koller:1994:FAF

[KMvS94] Daphne Koller, Nimrod Megiddo, and Bernhard von Stengel. Fast algorithms for finding randomized strategies in game trees. In ACM [ACM94], pages 750–759. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p750-koller/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p750-koller/p750-koller.

pdf. ACM order no. 508930.

Kirkpatrick:1990:QST

[KMY90] D. G. Kirkpatrick, B. Mishra, and C. K. Yap. Quantitative Steinitz’s theorems with applications to multifingered grasping. In ACM [ACM90], pages 341–351. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/100216/p341-kirkpatrick/

. ACM order no. 508900.

Kremer:1995:ROR

[KNR95] Ilan Kremer, Noam Nisan, and Dana Ron. On randomized oneround communication complexity. In ACM [ACM95], pages 596–605. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p596-kremer/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p596-kremer/p596-kremer.

pdf. ACM order no. 508950.

Karlin:1994:FTB

[KNT94] Anna R. Karlin, Greg Nelson, and Hisao Tamaki. On the fault tolerance of the butterfly. In ACM [ACM94], pages 125–133. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p125-karlin/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p125-karlin/p125-karlin.

pdf. ACM order no. 508930.

Ko:1991:IES

[Ko91] Ker-I. Ko. Integral equations, systems of quadratic equations, and exponential time completeness. In ACM [ACM91], pages 10–20. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p10-ko/;   http:

//www.acm.org/pubs/articles/
proceedings/stoc/103418/p10-
ko/p10-ko.pdf. ACM order no.

508910.

Kushilevitz:1995:LSP

[KOR95] Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Log-space polynomial end-to-end communication. In ACM [ACM95], pages 559–568. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p559-kushilevitz/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p559-kushilevitz/p559-
kushilevitz.pdf. ACM order

no. 508950.

Kushilevitz:1996:CLS

[KOR96] Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Characterizing linear size circuits in terms of privacy. In ACM [ACM96], pages 541–

550. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p541-kushilevitz/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p541-kushilevitz/p541-
kushilevitz.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Kushilevitz:1998:ESA

[KOR98] Eyal Kushilevitz, Rafail Ostrovsky, and Yuval Rabani.

Efficient search for approximate nearest neighbor in high dimensional spaces. In ACM [ACM98], pages 614– 623. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p614-kushilevitz/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p614-kushilevitz/p614-
kushilevitz.pdf. ACM order

number 508980.

Kosaraju:1994:RTP

[Kos94] S. Rao Kosaraju. Real-time pattern matching and quasireal-time construction of suffix trees (preliminary version). In ACM [ACM94], pages 310–316. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p310-kosaraju/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p310-kosaraju/p310-
kosaraju.pdf. ACM order no.

508930.

Klugerman:1992:SDC

[KP92] Michael Klugerman and C. Greg Plaxton. Small-depth counting networks. In ACM [ACM92], pages 417–428. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p417-klugerman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p417-klugerman/p417-

                                        klugerman.pdf. ACM order no.

508920.

Kalyanasundaram:1994:FTS

[KP94a] Bala Kalyanasundaram and Kirk R. Pruhs. Fault-tolerant scheduling. In ACM [ACM94], pages 115–124. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p115-kalyanasundaram/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p115-kalyanasundaram/
                                        p115-kalyanasundaram.pdf.

ACM order no. 508930.

Koutsoupias:1994:SC

[KP94b] Elias Koutsoupias and Christos Papadimitriou. On the k-server conjecture. In ACM [ACM94], pages 507–511. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p507-koutsoupias/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p507-koutsoupias/p507-
                                        koutsoupias.pdf. ACM order

no. 508930.

Krause:1994:CPD

[KP94c] Matthias Krause and Pavel Pudlák. On the computational power of depth 2 circuits with threshold and modulo gates. In ACM [ACM94], pages 48–57. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p48-krause/;

http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p48-
krause/p48-krause.pdf. ACM

order no. 508930.

Karger:1995:AMC

[KP95] David Karger and Serge Plotkin.

Adding multiple cost constraints to combinatorial optimization problems, with applications to multicommodity flows. In ACM [ACM95], pages 18–25. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p18-karger/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p18-
karger/p18-karger.pdf. ACM

order no. 508950.

Klein:1993:EMN

[KPR93] Philip Klein, Serge A. Plotkin, and Satish Rao. Excluded minors, network decomposition, and multicommodity flow. In ACM [ACM93], pages 682–690. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Kleinberg:1998:SP

[KPR98] Jon Kleinberg, Christos Papadimitriou, and Prabhakar Raghavan. Segmentation problems. In ACM [ACM98], pages 473–482. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p473-kleinberg/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        276698/p473-kleinberg/p473-
                                        kleinberg.pdf. ACM order

number 508980.

Kedem:1992:EPT

[KPRR92] Z. M. Kedem, K. V. Palem, M. O. Rabin, and A. Raghunathan. Efficient program transformations for resilient parallel computation via randomization (preliminary version). In ACM [ACM92], pages 306–317. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p306-kedem/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p306-
                                        kedem/p306-kedem.pdf. ACM

order no. 508920.

Kedem:1991:CTD

[KPRS91] Z. M. Kedem, K. V. Palem, A. Raghunathan, and P. G. Spirakis. Combining tentative and definite executions for very fast dependable parallel computing. In ACM [ACM91], pages 381–390. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p381-kedem/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p381-
                                        kedem/p381-kedem.pdf. ACM

order no. 508910.

Kedem:1990:ERP

[KPS90] Z. M. Kedem, K. V. Palem, and P. G. Spirakis. Efficient robust parallel computations. In ACM [ACM90], pages 138–148. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p138-kedem/. ACM

order no. 508900.

Kilian:1997:PCP

[KPT97] Joe Kilian, Erez Petrank, and Gábor Tardos. Probabilistically checkable proofs with zero knowledge. In ACM [ACM97], pages 496–

505. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p496-kilian/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p496-kilian/p496-kilian.

pdf. ACM order no. 508970.

Khuller:1995:IAA

[KR95] Samir Khuller and Balaji Raghavachari. Improved approximation algorithms for uniform connectivity problems. In ACM [ACM95], pages 1–

10. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p1-khuller/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p1-
khuller/p1-khuller.pdf. ACM

order no. 508950.

Kumar:1999:CRP

[KR99] V. S. Anil Kumar and H. Ramesh.

Covering rectilinear polygons with axis-parallel rectangles.

In ACM [ACM99], pages 445–

454. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p445-anil_kumar/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p445-kumar/p445-kumar.

pdf. ACM order number 508990.

Karloff:1991:LBR

[KRR91] Howard Karloff, Yuval Rabani, and Yiftach Ravid. Lower bounds for randomized k-server and motionplanning algorithms. In ACM [ACM91], pages 279–

288. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p279-karloff/.

ACM order no. 508910.

Klein:1994:FSP

[KRRS94] Philip Klein, Satish Rao, Monika Rauch, and Sairam Subramanian. Faster shortestpath algorithms for planar graphs. In ACM [ACM94], pages 27–37. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p27-klein/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p27-
klein/p27-klein.pdf. ACM

order no. 508930.

Kenyon:1993:MLG

[KRS93] Claire Kenyon, Dana Randall, and Alistair Sinclair. Matchings in lattice graphs. In ACM [ACM93], pages 738–746. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Kleinberg:1997:ABB

[KRT97] Jon Kleinberg, Yuval Rabani, and Éva Tardos. Allocating bandwidth for bursty connections. In ACM [ACM97], pages 664–673. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p664-kleinberg/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p664-kleinberg/p664-
kleinberg.pdf. ACM order no.

508970.

Khuller:1994:LDS

[KRY94] Samir Khuller, Balaji Raghavachari, and Neal Young. Low degree spanning trees of small weight. In ACM [ACM94], pages 412–421. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p412-khuller/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p412-khuller/p412-

khuller.pdf. ACM order no. 508930.

Klein:1992:PRA

[KS92] Philip N. Klein and S. Sairam.

A parallel randomized approximation scheme for shortest paths. In ACM [ACM92], pages 750–758. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p750-klein/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p750-
klein/p750-klein.pdf. ACM

order no. 508920.

Karger:1993:AMC

[KS93] David R. Karger and Clifford Stein. An Õ(n^{2}) algorithm for minimum cuts. In ACM [ACM93], pages 757–765. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Kaltofen:1995:STF

[KS95] Erich Kaltofen and Victor Shoup. Subquadratic-time factoring of polynomials over finite fields. In ACM [ACM95], pages 398–406. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p398-kaltofen/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p398-kaltofen/p398-
kaltofen.pdf. ACM order no.

508950.

Kapur:1996:SCD

[KS96] Deepak Kapur and Tushar Saxena. Sparsity considerations in Dixon resultants. In ACM [ACM96], pages 184–

191. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/

stoc/237814/p184-kapur/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p184-
kapur/p184-kapur.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

King:1999:FDA

[KS99] Valerie King and Garry Sagert.

A fully dynamic algorithm for maintaining the transitive closure. In ACM [ACM99], pages 492–498. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p492-king/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p492-
king/p492-king.pdf. ACM or-

der number 508990.

Klein:1990:LRM

[KST90] P. Klein, C. Stein, and É. Tardos. Leighton-rao might be practical: faster approximation algorithms for concurrent flow with uniform capacities. In ACM [ACM90], pages 310–321. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p310-klein/. ACM

order no. 508900.

Khanna:1997:CCA

[KSW97] Sanjeev Khanna, Madhu Sudan, and David P. Williamson. A complete classification of the approximability of maximization problems derived from Boolean constraint satisfaction. In ACM [ACM97], pages 11–20. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p11-khanna/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p11-
khanna/p11-khanna.pdf. ACM

order no. 508970.

Klein:1994:RLT

[KT94] Philip N. Klein and Robert E. Tarjan. A randomized lineartime algorithm for finding minimum spanning trees. In ACM [ACM94], pages 9–15. ISBN 0-89791-663-8. LCCN QA76 .A15

1994. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/195058/p9-klein/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/195058/p9-
klein/p9-klein.pdf. ACM or-

der no. 508930.

Kaplan:1995:PLC

[KT95a] Haim Kaplan and Robert E. Tarjan. Persistent lists with catenation via recursive slowdown. In ACM [ACM95], pages 93–102. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p93-kaplan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p93-
kaplan/p93-kaplan.pdf. ACM

order no. 508950.

Kleinberg:1995:ADP

[KT95b] Jon Kleinberg and Éva Tardos. Approximations for the disjoint paths problem in high-diameter planar networks. In ACM [ACM95], pages 26–35. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p26-kleinberg/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p26-kleinberg/p26-
kleinberg.pdf. ACM order no.

508950.

Kaplan:1996:PFR

[KT96] Haim Kaplan and Robert E.

Tarjan. Purely functional representations of catenable sorted lists. In ACM [ACM96], pages 202–211. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p202-kaplan/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p202-kaplan/p202-kaplan.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Kfoury:1990:USU

[KTU90] A. J. Kfoury, J. Tiuryn, and P. Urzyczyn. The undecidability of the semi-unification problem. In ACM [ACM90], pages 468–476. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p468-kfoury/.

ACM order no. 508900.

Kellerer:1996:ANR

[KTW96] Hans Kellerer, Thomas Tautenhahn, and Gerhard J. Woeginger. Approximability and nonapproximability results for minimizing total flow time on a single machine. In ACM [ACM96], pages 418– 426. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p418-kellerer/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p418-kellerer/p418-
                                        kellerer.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Kurshan:1994:CV

[Kur94] R. P. Kurshan. The complexity of verification. In ACM [ACM94], pages 365– 371. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p365-kurshan/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p365-kurshan/p365-

kurshan.pdf. ACM order no. 508930.

Khuller:1992:BAG

[KV92] Samir Khuller and Uzi Vishkin. Biconnectivity approximations and graph carvings. In ACM [ACM92], pages 759– 770. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p759-khuller/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p759-khuller/p759-

khuller.pdf. ACM order no. 508920.

Kannan:1997:SLP

[KV97] Ravi Kannan and Santosh Vempala. Sampling lattice points. In ACM [ACM97], pages 696–

700. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p696-kannan/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p696-kannan/p696-kannan.

pdf. ACM order no. 508970.

Kupferman:1998:WAA

[KV98] Orna Kupferman and Moshe Y.

Vardi. Weak alternating automata and tree automata emptiness. In ACM [ACM98], pages 224–233. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p224-kupferman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p224-kupferman/p224-
kupferman.pdf. ACM order

number 508980.

Klivans:1999:GNS

[KvM99] Adam R. Klivans and Dieter van Melkebeek. Graph nonisomorphism has subexponential size proofs unless the polynomial-time hierarchy collapses. In ACM [ACM99], pages 659–667. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p659-klivans/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p659-klivans/p659-

klivans.pdf. ACM order number 508990.

Karp:1990:OAL

[KVV90] R. M. Karp, U. V. Vazirani, and V. V. Vazirani. An optimal algorithm for on-line bipartite matching. In ACM [ACM90], pages 352–358. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p352-karp/. ACM

order no. 508900.

Karchmer:1993:CND

[KW93] M. Karchmer and A. Wigderson. Characterizing nondeterministic circuit size. In ACM [ACM93], pages 532–540. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Kivinen:1995:AVE

[KW95] Jyrki Kivinen and Manfred K. Warmuth. Additive versus exponentiated gradient updates for linear prediction. In ACM [ACM95], pages 209– 218. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/

stoc/225058/p209-kivinen/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p209-kivinen/p209-

kivinen.pdf. ACM order no. 508950.

Khanna:1998:IDB

[KZ98] Sanjeev Khanna and Shiyu Zhou. On indexed data broadcast. In ACM [ACM98], pages 463–472. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p463-khanna/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p463-khanna/p463-khanna.

pdf. ACM order number 508980.

LaPoutre:1990:LBU

[La 90] J. A. La Poutré. Lower bounds for the union-find and the splitfind problem on pointer machines. In ACM [ACM90], pages 34–44. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p34-la_poutre/

. ACM order no. 508900.

LaPoutre:1994:AAI

[La 94] Johannes A. La Poutré. Alphaalgorithms for incremental planarity testing (preliminary version). In ACM [ACM94], pages 706–715. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p706-la_poutre/

;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p706-la_poutre/p706-
la_poutre.pdf. ACM order no.

508930.

Lakshman:1990:CCG

[Lak90] Y. N. Lakshman. On the complexity of computing a Gröbner basis for the radical of a zero dimensional ideal. In ACM [ACM90], pages 555–563. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p555-lakshman/

. ACM order no. 508900.

Leighton:1992:MMR

[Lei92] Tom Leighton. Methods for message routing in parallel machines. In ACM [ACM92], pages 77–96. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p77-leighton/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p77-leighton/p77-leighton.

pdf. ACM order no. 508920.

Lo:1997:AUS

[LH97] Wai-Kau Lo and Vassos Hadzilacos. All of us are smarter than any of us: wait-free hierarchies are not robust. In ACM [ACM97], pages 579–

588. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p579-lo/;   http:

//www.acm.org/pubs/articles/
proceedings/stoc/258533/p579-
lo/p579-lo.pdf. ACM order

no. 508970.

Lin:1992:FTP

[Lin92a]  Geng Lin.  Fault tolerant

planar communication networks. In ACM [ACM92], pages 133–139. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p133-lin/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p133-
lin/p133-lin.pdf. ACM order

no. 508920.

Lindell:1992:LAT

[Lin92b] Steven Lindell. A logspace algorithm for tree canonization (extended abstract). In ACM [ACM92], pages 400– 404. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p400-lindell/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p400-lindell/p400-

lindell.pdf. ACM order no. 508920.

Lovasz:1999:FMA

[LK99] László Lovász and Ravi Kannan. Faster mixing via average conductance. In ACM [ACM99], pages 282– 287. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/301250/p282-lovasz/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p282-lovasz/p282-lovasz.

pdf. ACM order number 508990.

Lenstra:1990:NFS

[LLMP90] A. K. Lenstra, H. W. Lenstra, M. S. Manasse, and J. M. Pollard. The number field sieve. In ACM [ACM90], pages 564–572. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p564-lenstra/.

ACM order no. 508900.

Linial:1993:ECS

[LLSZ93] Nati Linial, Michael Luby, Michael Saks, and David Zuckerman. Efficient construction of a small hitting set for combinatorial rectangles in high dimension. In ACM [ACM93], pages 258–267. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p258-linial/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p258-linial/p258-linial.

pdf. ACM order no. 508930.

Littlestone:1991:LLL

[LLW91] Nicholas Littlestone, Philip M. Long, and Manfred K. Warmuth. On-line learning of linear functions. In ACM [ACM91], pages 465–

475. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p465-littlestone/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p465-littlestone/p465-
littlestone.pdf. ACM order

no. 508910.

Luby:1997:PLR

[LMS^{+}97] Michael G. Luby, Michael Mitzenmacher, M. Amin Shokrollahi, Daniel A. Spielman, and Volker Stemann. Practical loss-resilient codes. In ACM [ACM97], pages 150–

159. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p150-luby/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p150-
luby/p150-luby.pdf. ACM or-

der no. 508970.

Linial:1998:TEM

[LMS98] Nathan Linial, Avner Magen, and Michael E. Saks. Trees and Euclidean metrics. In ACM [ACM98], pages 169–

175. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p169-linial/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p169-linial/p169-linial.

pdf. ACM order number 508980.

Luby:1998:ALD

[LMSS98] M. Luby, M. Mitzenmacher, A. Shokrollah, and D. Spielman. Analysis of low density codes and improved designs using irregular graphs. In ACM [ACM98], pages 249–

258. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p249-luby/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p249-
luby/p249-luby.pdf. ACM or-

der number 508980.

Li:1999:FSR

[LMW99] Ming Li, Bin Ma, and Lusheng Wang. Finding similar regions in many strings. In ACM [ACM99], pages 473–

482. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p473-li/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/301250/p473-
li/p473-li.pdf. ACM order

number 508990.

Linial:1990:AIE

[LN90] N. Linial and N. Nisan. Approximate inclusion-exclusion. In ACM [ACM90], pages 260–270. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p260-linial/.

ACM order no. 508900.

Luby:1993:PAA

[LN93] Michael Luby and Noam Nisan. A parallel approximation algorithm for positive linear programming. In ACM [ACM93], pages 448–457. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Levcopoulos:1998:EAC

[LNS98] Christos Levcopoulos, Giri Narasimhan, and Michiel Smid. Efficient algorithms for constructing fault-tolerant geometric spanners. In ACM [ACM98], pages 186–195. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p186-levcopoulos/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p186-levcopoulos/p186-
levcopoulos.pdf. ACM order

number 508980.

Lafferty:1997:STE

[LR97a] John D. Lafferty and Daniel N.

Rockmore. Spectral techniques for expander codes. In ACM [ACM97], pages 160–

167. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p160-lafferty/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p160-lafferty/p160-
lafferty.pdf. ACM order no.

508970.

Leonardi:1997:ATF

[LR97b] Stefano Leonardi and Danny Raz. Approximating total flow time on parallel machines. In ACM [ACM97], pages 110– 119. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p110-leonardi/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p110-leonardi/p110-
leonardi.pdf. ACM order no.

508970.

Linial:1996:NEH

[LS96] Nathan Linial and Ori Sasson. Non-expansive hashing. In ACM [ACM96], pages 509–518. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p509-linial/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p509-linial/p509-linial.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Leighton:1991:FAA

[LSM^{+}91] Tom Leighton, Clifford Stein, Fillia Makedon, Éva Tardos, Serge Plotkin, and Spyros Tragoudas. Fast approximation algorithms for multicommodity flow problems. In ACM [ACM91], pages 101–

111. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p101-leighton/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p101-leighton/p101-

leighton.pdf. ACM order no.

508910.

Linial:1998:DSP

[LSW98] Nathan Linial, Alex Samorodnitsky, and Avi Wigderson. A deterministic strongly polynomial algorithm for matrix scaling and approximate permanents. In ACM [ACM98], pages 644–652. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p644-linial/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p644-linial/p644-linial.

pdf. ACM order number 508980.

Luks:1999:HIS

[Luk99] Eugene M. Luks. Hypergraph isomorphism and structural equivalence of Boolean functions. In ACM [ACM99], pages 652–658. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p652-luks/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p652-
luks/p652-luks.pdf. ACM or-

der number 508990.

Luby:1991:DAD

[LV91] Michael Luby and Boban Veličkovic. On deterministic approximation of DNF. In ACM [ACM91], pages 430–

438. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/

pubs/citations/proceedings/
stoc/103418/p430-luby/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p430-
luby/p430-luby.pdf. ACM or-

der no. 508910.

Lin:1992:AMP

[LV92] Jyh-Han Lin and Jeffrey Scott Vitter. ϵ-approximations with minimum packing constraint violation (extended abstract). In ACM [ACM92], pages 771–782. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p771-lin/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p771-
lin/p771-lin.pdf. ACM order

no. 508920.

Luby:1997:ACF

[LV97] Michael Luby and Eric Vigoda. Approximately counting up to four (extended abstract). In ACM [ACM97], pages 682–

687. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p682-luby/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p682-
luby/p682-luby.pdf. ACM or-

der no. 508970.

Lewin:1998:CPI

[LV98] Daniel Lewin and Salil Vadhan. Checking polynomial identities over any field: towards a derandomization? In ACM [ACM98], pages 438– 447. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p438-lewin/;
http://www.acm.org/pubs/articles/
proceedings/stoc/276698/p438-
lewin/p438-lewin.pdf. ACM

order number 508980.

Lovasz:1995:ESR

[LW95] László Lovász and Peter Winkler. Efficient stopping rules for Markov chains. In ACM [ACM95], pages 76–

82. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p76-lovasz/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p76-
lovasz/p76-lovasz.pdf. ACM

order no. 508950.

Lee:1992:OMT

[LY92] David Lee and Mihalis Yannakakis. Online minimization of transition systems (extended abstract). In ACM [ACM92], pages 264–274. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p264-lee/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/129712/p264-
lee/p264-lee.pdf. ACM order

no. 508920.

Lund:1993:HAM

[LY93] Carsten Lund and Mihalis Yannakakis. On the hardness of approximating minimization problems. In ACM [ACM93], pages 286–293. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p286-lund/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p286-
lund/p286-lund.pdf. ACM or-

der no. 508930.

Lipton:1994:SSL

[LY94] Richard J. Lipton and Neal E. Young. Simple strategies for large zero-sum games with applications to complexity theory. In ACM [ACM94], pages 734–740. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p734-lipton/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p734-lipton/p734-lipton.

pdf. ACM order no. 508930.

Ma:1996:SFT

[Ma96]  Yuan Ma.  An O(n log n)-

size fault-tolerant sorting network (extended abstract). In ACM [ACM96], pages 266–

275. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p266-ma/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/237814/p266-
ma/p266-ma.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Maass:1993:BCP

[Maa93]  Wolfgang Maass.  Bounds

for the computational power and learning complexity of analog neural nets. In ACM [ACM93], pages 335–

344. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p335-maass/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p335-
maass/p335-maass.pdf. ACM

order no. 508930.

Malitz:1992:ARP

[Mal92]  Seth Malitz.  On the an-

gular resolution of planar graphs. In ACM [ACM92], pages 527–538. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p527-malitz/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p527-malitz/p527-malitz.

pdf. ACM order no. 508920.

Matousek:1991:AOG

[Mat91] Jiří Matoušek. Approximations and optimal geometric divide-and-conquer. In ACM [ACM91], pages 505–

511. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p505-matousek/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p505-matousek/p505-

                                        matousek.pdf. ACM order no.

508910.

Maurer:1991:PCS

[Mau91] Ueli M. Maurer. Perfect cryptographic security from partially independent channels. In ACM [ACM91], pages 561–

571. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p561-maurer/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p561-maurer/p561-maurer.

pdf. ACM order no. 508910.

McCormick:1996:FAP

[McC96] S. Thomas McCormick. Fast algorithms for parametric scheduling come from extensions to parametric maximum flow. In ACM [ACM96], pages 319–

328. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p319-mccormick/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p319-mccormick/p319-
                                        mccormick.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Marathe:1994:ASP

[MHSR94] M. V. Marathe, H. B. Hunt, R. E. Stearns, and V. Radhakrishnan. Approximation schemes for PSPACE-complete problems for succinct specifications (preliminary version). In ACM [ACM94], pages 468–477. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p468-marathe/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p468-marathe/p468-

marathe.pdf. ACM order no. 508930.

Micciancio:1997:ODS

[Mic97]  Daniele Micciancio.  Obliv-

ious data structures: applications to cryptography. In ACM [ACM97], pages 456– 464. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p456-micciancio/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p456-micciancio/p456-
micciancio.pdf. ACM order

no. 508970.

Miltersen:1994:LBU

[Mil94] Peter Bro Miltersen. Lower bounds for union-split-find related problems on random access machines. In ACM [ACM94], pages 625– 634. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p625-miltersen/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p625-bro_miltersen/

                                        p625-bro_miltersen.pdf. ACM

order no. 508930.

Milenkovic:1996:TPC

[Mil96] Victor J. Milenkovic. Translational polygon containment and minimal enclosure using linear programming based restriction. In ACM [ACM96], pages 109–118. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p109-milenkovic/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p109-milenkovic/p109-
                                        milenkovic.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Matousek:1992:HSC

[MLS92] Jiří Matoušek, Chi-Yuan Lo, and William Steiger. Hamsandwich cuts in R^{d}. In ACM [ACM92], pages 539– 545. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p539-matousek/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p539-lo/p539-lo.pdf.

ACM order no. 508920.

Miltersen:1995:DSA

[MNSW95] Peter Bro Miltersen, Noam Nisan, Shmuel Safra, and Avi Wigderson. On data structures and asymmetric communication complexity. In ACM [ACM95], pages 103– 111. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p103-miltersen/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p103-miltersen/p103-
miltersen.pdf. ACM order no.

508950.

Mansour:1990:CCU

[MNT90] Y. Mansour, N. Nisan, and P. Tiwari. The computational complexity of universal hashing. In ACM [ACM90], pages 235–243. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p235-mansour/.

ACM order no. 508900.

Mansour:1994:TOB

[MNV94] Yishay Mansour, Noam Nisan, and Uzi Vishkin. Trade-offs between communication throughput and parallel time. In ACM [ACM94], pages 372– 381. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p372-mansour/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p372-mansour/p372-

mansour.pdf. ACM order no. 508930.

Mohar:1996:EGA

[Moh96]  Bojan Mohar.  Embedding

graphs in an arbitrary surface in linear time. In ACM [ACM96], pages 392–

397. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p392-mohar/;
http://www.acm.org/pubs/articles/
proceedings/stoc/237814/p392-
mohar/p392-mohar.pdf. ACM

order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Mayer:1992:SSS

[MOOY92] Alain Mayer, Yoram Ofek, Rafail Ostrovsky, and Moti Yung. Self-stabilizing symmetry breaking in constantspace (extended abstract). In ACM [ACM92], pages 667–

678. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p667-mayer/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p667-
mayer/p667-mayer.pdf. ACM

order no. 508920. Muthukrishnan:1994:NSS

[MP94] S. Muthukrishnan and Krishna Palem. Non-standard stringology: algorithms and complexity. In ACM [ACM94], pages 770–779. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p770-muthukrishnan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p770-muthukrishnan/

p770-muthukrishnan.pdf. ACM

order no. 508930.

Maciel:1997:AFP

[MP97] Alexis Maciel and Toniann Pitassi. On ACC^{0}[p^{k}] Frege proofs. In ACM [ACM97], pages 720–729. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p720-maciel/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p720-maciel/p720-maciel.

pdf. ACM order no. 508970.

Mourrain:1998:AAS

[MP98] Bernard Mourrain and Victor Y. Pan. Asymptotic acceleration of solving multivariate polynomial systems of equations. In ACM [ACM98], pages 488–496. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p488-mourrain/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p488-mourrain/p488-
mourrain.pdf. ACM order

number 508980.

MacKenzie:1994:CRP

[MPR94] P. D. MacKenzie, C. G. Plaxton, and R. Rajaraman. On contention resolution protocols and associated probabilistic phenomena. In ACM [ACM94], pages 153–162. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/195058/p153-mackenzie/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p153-mackenzie/p153-
                                        mackenzie.pdf. ACM order no.

508930.

Mansour:1995:MOP

[MPS95] Yishay Mansour and Boaz Patt-Shamir. Many-to-one packet routing on grids. In ACM [ACM95], pages 258–

267. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p258-mansour/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p258-mansour/p258-

mansour.pdf. ACM order no.

508950.

Muthukrishnan:1999:CGL

[MPS.S99] S. Muthukrishnan, Mike Paterson, Süleyman Cenk S.ahinalp, and Torsten Suel. Compact grid layouts of multi-level networks. In ACM [ACM99], pages 455–463. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p455-muthukrishnan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p455-muthukrishnan/
                                        p455-muthukrishnan.pdf. ACM

order number 508990.

Malkhi:1997:BQS

[MR97] Dahlia Malkhi and Michael Reiter. Byzantine quorum systems. In ACM [ACM97], pages 569–578. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p569-malkhi/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p569-malkhi/p569-malkhi.

pdf. ACM order no. 508970.

Molloy:1998:FAA

[MR98] Michael Molloy and Bruce Reed. Further algorithmic aspects of the local lemma. In ACM [ACM98], pages 524–

529. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p524-molloy/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p524-molloy/p524-molloy.

pdf. ACM order number 508980.

McCuaig:1997:PPO

[MRST97] William McCuaig, Neil Robertson, P. D. Seymour, and Robin Thomas. Permanents, Pfaffian orientations, and even directed circuits (extended abstract). In ACM [ACM97], pages 402–

405. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p402-mccuaig/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p402-mccuaig/p402-

mccuaig.pdf. ACM order no. 508970.

Maggs:1992:SAR

[MS92] Bruce M. Maggs and Ramesh K. Sitaraman. Simple algorithms for routing on butterfly networks with bounded queues. In ACM [ACM92], pages 150–161. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p150-maggs/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p150-
                                        maggs/p150-maggs.pdf. ACM

order no. 508920.

Macintyre:1993:FRS

[MS93a] Angus Macintyre and Eduardo D. Sontag. Finiteness results for sigmoidal “neural” networks. In ACM [ACM93], pages 325–334. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p325-macintyre/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p325-macintyre/p325-
                                        macintyre.pdf. ACM order no.

508930.

Matousek:1993:DAT

[MS93b] Jiří Matoušek and Otfried Schwarzkopf. A deterministic algorithm for the threedimensional diameter problem. In ACM [ACM93], pages 478–

484. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Miller:1990:STT

[MT90] G. L. Miller and W. Thurston.

Separators in two and three dimensions. In ACM [ACM90], pages 300–309. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p300-miller/.

ACM order no. 508900.

Miller:1995:DBN

[MTTW95] Gary L. Miller, Dafna Talmor, Shang-Hua Teng, and Noel Walkington. A Delaunay based numerical method for three dimensions: generation, formulation, and partition. In ACM [ACM95], pages 683–692. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p683-miller/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p683-miller/p683-miller.

pdf. ACM order no. 508950.

Mulmuley:1990:OSC

[Mul90] K. Mulmuley. Output sensitive construction of levels and Voronoi diagrams in R^{d} of order 1 to k. In ACM [ACM90], pages 322–330. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p322-mulmuley/

. ACM order no. 508900.

Mulmuley:1991:HSR

[Mul91] Ketan Mulmuley. Hidden surface removal with respect to a moving view point. In ACM [ACM91], pages 512– 522. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p512-mulmuley/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p512-mulmuley/p512-
mulmuley.pdf. ACM order no.

508910.

Mulmuley:1994:LBP

[Mul94]  Ketan Mulmuley.  Lower

bounds for parallel linear programming and other problems. In ACM [ACM94], pages 603–614. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p603-mulmuley/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p603-mulmuley/p603-
mulmuley.pdf. ACM order no.

508930.

Mulmuley:1997:TAP

[Mul97] Ketan Mulmuley. Is there an algebraic proof for P ≠ NC (extended abstract)? In ACM [ACM97], pages 210– 219. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p210-mulmuley/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p210-mulmuley/p210-

mulmuley.pdf. ACM order no.

508970.

Matias:1991:CHP

[MV91] Yossi Matias and Uzi Vishkin.

Converting high probability into nearly-constant time — with applications to parallel hashing. In ACM [ACM91], pages 307–316. ISBN 0-89791-397-3. LCCN QA 76.6 A13

1991. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/103418/p307-matias/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
103418/p307-matias/p307-matias.

pdf. ACM order no. 508910.

Maggs:1997:IRS

[MV97] Bruce M. Maggs and Berthold Vöcking. Improved routing and sorting on multibutterflies. In ACM [ACM97], pages 517–530. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p517-maggs/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p517-
maggs/p517-maggs.pdf. ACM

order no. 508970.

Menezes:1991:REC

[MVO91] Alfred Menezes, Scott Vanstone, and Tatsuaki Okamoto. Reducing elliptic curve logarithms to logarithms in a finite field. In ACM [ACM91], pages 80–89. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/103418/p80-menezes/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p80-menezes/p80-menezes.

pdf. ACM order no. 508910.

Ma:1994:CCR

[MvzG94] Keju Ma and Joachim von zur Gathen. The computational complexity of recognizing permutation functions. In ACM [ACM94], pages 392–

401. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p392-ma/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p392-
                                        ma/p392-ma.pdf. ACM order

no. 508930.

Naor:1996:EME

[Nao96]  Moni Naor.  Evaluation

may be easier than generation (extended abstract). In ACM [ACM96], pages 74–

83. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p74-naor/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/237814/p74-
                                        naor/p74-naor.pdf. ACM or-

der no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Nisan:1997:PJR

[NBY97] Noam Nisan and Ziv Bar-Yossef. Pointer jumping requires concurrent read. In ACM [ACM97], pages 549–

558. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p549-nisan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p549-
nisan/p549-nisan.pdf. ACM

order no. 508970.

Nagamochi:1996:DTE

[NI96] Hiroshi Nagamochi and Toshihide Ibaraki. Deterministic Õ(nm) time edge-splitting in undirected graphs. In ACM [ACM96], pages 64–

73. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p64-nagamochi/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p64-nagamochi/p64-
nagamochi.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Nisan:1990:PGS

[Nis90] N. Nisan. Pseudorandom generators for space-bounded computations. In ACM [ACM90], pages 204–212. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p204-nisan/. ACM

order no. 508900.

Nisan:1991:LBN

[Nis91] Noam Nisan. Lower bounds for non-commutative computation. In ACM [ACM91], pages 410–418. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p410-nisan/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/103418/p410-
                                        nisan/p410-nisan.pdf. ACM

order no. 508910.

Nisan:1992:RS

[Nis92] Noam Nisan. RL ⊆ SC. In ACM [ACM92], pages 619–

623. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p619-nisan/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p619-
                                        nisan/p619-nisan.pdf. ACM

order no. 508920.

Naor:1990:SBP

[NN90] J. Naor and M. Naor. Smallbias probability spaces: efficient constructions and applications. In ACM [ACM90], pages 213–223. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p213-naor/. ACM

order no. 508900.

Naor:1999:OTP

[NP99] Moni Naor and Benny Pinkas. Oblivious transfer and polynomial evaluation. In ACM [ACM99], pages 245–254. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. URL http://www.

                                        acm.org/pubs/citations/proceedings/

stoc/301250/p245-naor/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p245-
naor/p245-naor.pdf. ACM or-

der number 508990.

Naor:1997:CPR

[NR97] Moni Naor and Omer Reingold. On the construction of pseudo-random permutations: Luby-Rackoff revisited (extended abstract). In ACM [ACM97], pages 189–

199. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p189-naor/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p189-
naor/p189-naor.pdf. ACM or-

der no. 508970.

Nisan:1999:AMD

[NR99] Noam Nisan and Amir Ronen. Algorithmic mechanism design (extended abstract). In ACM [ACM99], pages 129–

140. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p129-nisan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p129-
nisan/p129-nisan.pdf. ACM

order number 508990.

Nisan:1992:DBF

[NS92] Noam Nisan and Mario Szegedy.

On the degree of Boolean functions as real polynomials. In ACM [ACM92], pages 462–467. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p462-nisan/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p462-
                                        nisan/p462-nisan.pdf. ACM

order no. 508920.

Naor:1993:WCC

[NS93] Moni Naor and Larry Stockmeyer. What can be computed locally? In ACM [ACM93], pages 184–193. ISBN 0-89791-591-7. LCCN QA 76.6 A13

                                        1993. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/167088/p184-naor/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/167088/p184-
                                        naor/p184-naor.pdf. ACM or-

der no. 508930.

Newman:1996:PVP

[NS96] Ilan Newman and Mario Szegedy. Public vs. private coin flips in one round communication games (extended abstract). In ACM [ACM96], pages 561–

570. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p561-newman/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p561-newman/p561-newman.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Natanzon:1998:PAA

[NSS98] Assaf Natanzon, Ron Shamir, and Roded Sharan. A polynomial approximation algorithm for the minimum fill-in problem. In ACM [ACM98], pages 41–47. ISBN 0-89791-962-

9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p41-natanzon/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p41-natanzon/p41-natanzon.

pdf. ACM order number 508980.

Nisan:1995:SLC

[NTS95] Noam Nisan and Amnon Ta-Shma. Symmetric logspace is closed under complement. In ACM [ACM95], pages 140–

146. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p140-nisan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p140-
nisan/p140-nisan.pdf. ACM

order no. 508950.

Nisan:1991:RCC

[NW91] Noam Nisan and Avi Widgerson. Rounds in communication complexity revisited. In ACM [ACM91], pages 419–

429. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/103418/p419-nisan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/103418/p419-
nisan/p419-nisan.pdf. ACM

order no. 508910.

Nisan:1995:CBF

[NW95] Noam Nisan and Avi Wigderson. On the complexity of bilinear forms: dedicated to the memory of Jacques Morgenstern. In ACM [ACM95], pages 723–732. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p723-nisan/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p723-
                                        nisan/p723-nisan.pdf. ACM

order no. 508950.

Nayak:1999:QQC

[NW99] Ashwin Nayak and Felix Wu. The quantum query complexity of approximating the median and related statistics. In ACM [ACM99], pages 384–

393. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p384-nayak/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p384-
                                        nayak/p384-nayak.pdf. ACM

order number 508990.

Naor:1990:PKC

[NY90] M. Naor and M. Yung. Publickey cryptosystems provably secure against chosen ciphertext attacks. In ACM [ACM90], pages 427–437. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p427-naor/. ACM

order no. 508900.

Nisan:1993:MDS

[NZ93] Noam Nisan and David Zuckerman. More deterministic simulation in logspace. In ACM [ACM93], pages 235–

244. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p235-nisan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/167088/p235-
nisan/p235-nisan.pdf. ACM

order no. 508930.

Odlyzko:1994:SMR

[Odl94] Andrew M. Odlyzko. Search for the maximum of a random walk. In ACM [ACM94], pages 336–345. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p336-odlyzko/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p336-odlyzko/p336-

odlyzko.pdf. ACM order no. 508930.

Ogihara:1996:PHC

[Ogi96]  Mitsunori Ogihara.  The

PL hierarchy collapses. In ACM [ACM96], pages 84–

88. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p84-ogihara/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p84-ogihara/p84-ogihara.

pdf. ACM order no. 508960.

Also known as Federated Computing Research Conference (FCRS ’96).

Okamoto:1996:RBS

[Oka96] Tatsuaki Okamoto. On relationships between statistical zero-knowledge proofs. In ACM [ACM96], pages 649– 658. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p649-okamoto/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p649-okamoto/p649-

okamoto.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Ostrovsky:1997:UCD

[OR97] Rafail Ostrovsky and Yuval Rabani. Universal O(congestion + dilation+log^{1+ϵ} N) local control packet switching algorithms. In ACM [ACM97], pages 644– 653. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p644-ostrovsky/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p644-ostrovsky/p644-
                                        ostrovsky.pdf. ACM order no.

508970.

Ostrovsky:1994:SEL

[ORV94] Rafail Ostrovsky, Sridhar Rajagopalan, and Umesh Vazirani. Simple and efficient leader election in the full information model. In ACM [ACM94], pages 234–242. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p234-ostrovsky/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p234-ostrovsky/p234-
ostrovsky.pdf. ACM order no.

508930.

Ostrovsky:1997:PIS

[OS97] Rafail Ostrovsky and Victor Shoup. Private information storage (extended abstract). In ACM [ACM97], pages 294– 303. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p294-ostrovsky/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p294-ostrovsky/p294-
ostrovsky.pdf. ACM order no.

508970.

Olshevsky:1999:DAE

[OS99] Vadim Olshevsky and M. Amin Shokrollahi. A displacement approach to efficient decoding of algebraic-geometric codes. In ACM [ACM99], pages 235– 244. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p235-olshevsky/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p235-olshevsky/p235-

                                        olshevsky.pdf. ACM order

number 508990.

Ostrovsky:1990:ECO

[Ost90] R. Ostrovsky. Efficient computation on oblivious RAMs. In ACM [ACM90], pages 514–523. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p514-ostrovsky/

. ACM order no. 508900.

Ogiwara:1990:PTB

[OW90] M. Ogiwara and O. Watanabe. On polynomial time bounded truth-table reducibility of NP sets to sparse sets. In ACM [ACM90], pages 457–467. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p457-ogiwara/.

ACM order no. 508900.

Pan:1995:OPF

[Pan95] Victor Y. Pan. Optimal (up to polylog factors) sequential and parallel algorithms for approximating complex polynomial zeros. In ACM [ACM95], pages 741–750. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p741-pan/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p741-
                                        pan/p741-pan.pdf. ACM order

no. 508950.

Pan:1997:FSK

[Pan97] Victor Y. Pan. Faster solution of the key equation for decoding BCH error-correcting codes. In ACM [ACM97], pages 168–175. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p168-pan/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/258533/p168-
pan/p168-pan.pdf. ACM order

no. 508970.

Paturi:1992:DPA

[Pat92] Ramamohan Paturi. On the degree of polynomials that approximate symmetric Boolean functions (preliminary version). In ACM [ACM92], pages 468–474. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p468-paturi/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p468-paturi/p468-paturi.

pdf. ACM order no. 508920.

Pan:1999:CME

[PC99] Victor Y. Pan and Zhao Q.

Chen. The complexity of the matrix eigenproblem. In ACM [ACM99], pages 507–

516. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p507-pan/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/301250/p507-

                                        pan/p507-pan.pdf. ACM order

number 508990.

Pellegrini:1994:PLM

[Pel94] Marco Pellegrini. On point location and motion planning among simplices. In ACM [ACM94], pages 95– 104. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p95-pellegrini/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p95-pellegrini/p95-
                                        pellegrini.pdf. ACM order

no. 508930.

Phillips:1993:NIP

[Phi93] Cynthia A. Phillips. The network inhibition problem. In ACM [ACM93], pages 776–785. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Pippenger:1993:SRS

[Pip93]  Nicholas Pippenger.  Self-

routing superconcentrators. In ACM [ACM93], pages 355– 361. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p355-pippenger/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        167088/p355-pippenger/p355-
                                        pippenger.pdf. ACM order no.

508930.

Peinado:1998:RGE

[PL98] Marcus Peinado and Thomas Lengauer. Random generation of embedded graphs and an extension to Dobrushin uniqueness (extended abstract). In ACM [ACM98], pages 176– 185. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p176-peinado/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p176-peinado/p176-

peinado.pdf. ACM order number 508980.

Plaxton:1992:HSN

[Pla92]  C. Greg Plaxton.  A hy-

percubic sorting network with nearly logarithmic depth. In ACM [ACM92], pages 405– 416. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p405-plaxton/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p405-plaxton/p405-

plaxton.pdf. ACM order no. 508920.

Plandowski:1999:SWE

[Pla99]  Wojciech Plandowski.  Sat-

isfiability of word equations with constants is in NEXP-TIME. In ACM [ACM99], pages 721–725. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/301250/p721-plandowski/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p721-plandowski/p721-
                                        plandowski.pdf. ACM order

number 508990.

Ponzio:1995:LBI

[Pon95]  Stephen Ponzio.  A lower

bound for integer multiplication with read-once branching programs. In ACM [ACM95], pages 130–139. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p130-ponzio/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p130-ponzio/p130-ponzio.

pdf. ACM order no. 508950.

Park:1993:FMQ

[PP93] James K. Park and Cynthia A. Phillips. Finding minimumquotient cuts in planar graphs. In ACM [ACM93], pages 766–

775. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Panconesi:1990:QA

[PR90] A. Panconesi and D. Ranjan. Quantifiers and approximation. In ACM [ACM90], pages 446–456. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p446-panconesi/

. ACM order no. 508900.

Pudlak:1993:MRT

[PR93] P. Pudlák and V. Rödl. Modified ranks of tensors and the size of circuits. In ACM [ACM93], pages 523–531. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Ponzio:1999:CCP

[PRV99] Stephen J. Ponzio, Jaikumar Radhakrishnan, and S. Venkatesh. The communication complexity of pointer chasing: applications of entropy and sampling. In ACM [ACM99], pages 602–611. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p602-ponzio/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p602-ponzio/p602-ponzio.

pdf. ACM order number 508990.

Parnafes:1997:DPR

[PRW97] Itzhak Parnafes, Ran Raz, and Avi Wigderson. Direct product results and the GCD problem, in old and new communication models. In ACM [ACM97], pages 363–

372. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p363-parnafes/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p363-parnafes/p363-
parnafes.pdf. ACM order no.

508970.

Panconesi:1992:IDA

[PS92] Alessandro Panconesi and Aravind Srinivasan. Improved distributed algorithms for coloring and network decomposition problems. In ACM [ACM92], pages 581–592. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p581-panconesi/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p581-panconesi/p581-
panconesi.pdf. ACM order no.

508920.

Polishchuk:1994:NLS

[PS94] Alexander Polishchuk and Daniel A. Spielman. Nearlylinear size holographic proofs. In ACM [ACM94], pages 194–203. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p194-polishchuk/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p194-polishchuk/p194-
polishchuk.pdf. ACM order

no. 508930.

Patt-Shamir:1994:TCS

[PSR94] Boaz Patt-Shamir and Sergio Rajsbaum. A theory of clock synchronization (extended abstract). In ACM [ACM94], pages 810–819. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p810-patt-shamir/

;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p810-patt-shamir/p810-
patt-shamir.pdf. ACM order

no. 508930.

Phillips:1997:OTC

[PSTW97] Cynthia A. Phillips, Cliff Stein, Eric Torng, and Joel Wein. Optimal time-critical scheduling via resource augmentation (extended abstract). In ACM [ACM97], pages 140–

149. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p140-phillips/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p140-phillips/p140-
phillips.pdf. ACM order no.

508970.

Papadimitriou:1990:CLS

[PSY90] C. H. Papadimitriou, A. A.

Schäffer, and M. Yannakakis. On the complexity of local search. In ACM [ACM90], pages 438–445. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p438-papadimitriou/

. ACM order no. 508900.

Paturi:1997:ELB

[PSZ97] Ramamohan Paturi, Michael E.

Saks, and Francis Zane. Exponential lower bounds for depth 3 Boolean circuits. In ACM [ACM97], pages 86–

91. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p86-paturi/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/258533/p86-
                                        paturi/p86-paturi.pdf. ACM

order no. 508970.

Plotkin:1993:IBM

[PT93] Serge A. Plotkin and Éva Tardos. Improved bounds on the max-flow min-cut ratio for multicommodity flows. In ACM [ACM93], pages 691–697. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Pandurangan:1999:SDE

[PU99] Gopal Pandurangan and Eli Upfal. Static and dynamic evaluation of QoS properties. In ACM [ACM99], pages 566–573. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p566-pandurangan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p566-pandurangan/p566-
                                        pandurangan.pdf. ACM order

number 508990.

Pulleyblank:1995:TST

[Pul95] William R. Pulleyblank. Two Steiner tree packing problems. In ACM [ACM95], pages 383–

387. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p383-pulleyblank/
                                        ;   http://www.acm.org/pubs/

articles/proceedings/stoc/
225058/p383-pulleyblank/p383-
pulleyblank.pdf. ACM order

no. 508950.

Phillips:1993:OLB

[PW93] Steven Phillips and Jeffery Westbrook. Online load balancing and network flow. In ACM [ACM93], pages 402–

411. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p402-phillips/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p402-phillips/p402-
phillips.pdf. ACM order no.

508930.

Papadimitriou:1993:LPM

[PY93] Christos H. Papadimitriou and Mihalis Yannakakis. Linear programming without the matrix. In ACM [ACM93], pages 121–129. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p121-papadimitriou/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p121-papadimitriou/
p121-papadimitriou.pdf. ACM

order no. 508930.

Papadimitriou:1994:CBR

[PY94] Christos H. Papadimitriou and Mihalis Yannakakis. On complexity as bounded rationality (extended abstract). In ACM [ACM94], pages 726–

733. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p726-papadimitriou/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p726-papadimitriou/
                                        p726-papadimitriou.pdf. ACM

order no. 508930.

Paterson:1992:SMC

[PZ92] Michael S. Paterson and Uri Zwick. Shallow multiplication circuits and wise financial investments. In ACM [ACM92], pages 429–438. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p429-paterson/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p429-paterson/p429-
                                        paterson.pdf. ACM order no.

508920.

Rao:1992:FAF

[Rao92] Satish B. Rao. Faster algorithms for finding small edge cuts in planar graphs. In ACM [ACM92], pages 229–

240. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p229-rao/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p229-
                                        rao/p229-rao.pdf. ACM order

no. 508920.

Rauch:1994:IDS

[Rau94] Monika Rauch. Improved data structures for fully dynamic biconnectivity. In ACM [ACM94], pages 686–695. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p686-rauch/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p686-
rauch/p686-rauch.pdf. ACM

order no. 508930.

Raz:1995:PRT

[Raz95] Ran Raz. A parallel repetition theorem. In ACM [ACM95], pages 447–456. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p447-raz/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/225058/p447-
raz/p447-raz.pdf. ACM order

no. 508950.

Raz:1999:ESQ

[Raz99] Ran Raz. Exponential separation of quantum and classical communication complexity. In ACM [ACM99], pages 358–367. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p358-raz/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/301250/p358-
raz/p358-raz.pdf. ACM order

number 508990.

Reed:1992:FAS

[Ree92] Bruce A. Reed. Finding approximate separators and computing tree width quickly. In ACM [ACM92], pages 221– 228. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p221-reed/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p221-
                                        reed/p221-reed.pdf. ACM or-

der no. 508920.

Reif:1995:WEP

[Rei95] John H. Reif. Work efficient parallel solution of Toeplitz systems and polynomial GCD. In ACM [ACM95], pages 751–

761. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p751-reif/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p751-
                                        reif/p751-reif.pdf. ACM or-

der no. 508950.

Reif:1997:ACP

[Rei97] John H. Reif. Approximate complex polynomial evaluation in near constant work per point. In ACM [ACM97], pages 30–39. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p30-reif/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/258533/p30-
                                        reif/p30-reif.pdf. ACM or-

der no. 508970.

Ravi:1993:MBO

[RMR^{+}93] R. Ravi, M. V. Marathe, S. S. Ravi, D. J. Rosenkrantz, and H. B. Hunt. Many birds with one stone: multi-objective approximation algorithms. In ACM [ACM93], pages 438–447. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Rojas:1999:CDG

[Roj99]  J. Maurice Rojas.  On

the complexity of Diophantine geometry in low dimensions (extended abstract). In ACM [ACM99], pages 527–

536. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p527-rojas/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p527-
rojas/p527-rojas.pdf. ACM

order number 508990.

Rompel:1990:OWF

[Rom90] J. Rompel. One-way functions are necessary and sufficient for secure signatures. In ACM [ACM90], pages 387–394. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p387-rompel/.

ACM order no. 508900.

Razborov:1994:NP

[RR94] Alexander A. Razborov and Steven Rudich. Natural proofs. In ACM [ACM94], pages 204–213. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p204-razborov/

                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p204-razborov/p204-
                                        razborov.pdf. ACM order no.

508930.

Raz:1999:RRS

[RR99] Ran Raz and Omer Reingold. On recycling the randomness of states in space bounded computation. In ACM [ACM99], pages 159–168. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p159-raz/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p159-
                                        raz/p159-raz.pdf. ACM order

number 508990.

Rabani:1995:CVP

[RRS95] Yuval Rabani, Yuri Rabinovich, and Alistair Sinclair. A computational view of population genetics. In ACM [ACM95], pages 83–92. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p83-rabani/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p83-
                                        rabani/p83-rabani.pdf. ACM

order no. 508950.

Raz:1999:EAR

[RRV99] Ran Raz, Omer Reingold, and Salil Vadhan. Extracting all the randomness and reducing the error in Trevisan’s extractors. In ACM [ACM99], pages 149–158. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p149-raz/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/301250/p149-
raz/p149-raz.pdf. ACM order

number 508990.

Rackoff:1993:CDA

[RS93] Charles Rackoff and Daniel R.

Simon. Cryptographic defense against traffic analysis. In ACM [ACM93], pages 672–681. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Rajagopalan:1994:CTD

[RS94] Sridhar Rajagopalan and Leonard Schulman. A coding theorem for distributed computation. In ACM [ACM94], pages 790–799. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p790-rajagopalan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p790-rajagopalan/p790-
rajagopalan.pdf. ACM order

no. 508930.

Raschle:1995:RGT

[RS95] Thomas Raschle and Klaus Simon. Recognition of graphs with threshold dimension two. In ACM [ACM95], pages 650– 661. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p650-raschle/;
http://www.acm.org/pubs/

articles/proceedings/stoc/
225058/p650-raschle/p650-

raschle.pdf. ACM order no. 508950.

Raz:1997:SCE

[RS97] Ran Raz and Shmuel Safra. A sub-constant error-probability low-degree test, and a subconstant error-probability PCP characterization of NP. In ACM [ACM97], pages 475– 484. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p475-raz/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/258533/p475-
raz/p475-raz.pdf. ACM order

no. 508970.

Rao:1998:AGG

[RS98] Satish B. Rao and Warren D. Smith. Approximating geometrical graphs via “spanners” and “banyans”. In ACM [ACM98], pages 540– 550. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p540-rao/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/276698/p540-
rao/p540-rao.pdf. ACM order

number 508980.

Robertson:1996:EFC

[RSST96] Neil Robertson, Daniel P. Sanders, Paul Seymour, and Robin Thomas. Efficiently fourcoloring planar graphs. In ACM [ACM96], pages 571– 575. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p571-robertson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p571-robertson/p571-
robertson.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Russell:1999:LBL

[RSZ99] Alexander Russell, Michael Saks, and David Zuckerman. Lower bounds for leader election and collective coin-flipping in the perfect information model. In ACM [ACM99], pages 339–347. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p339-russell/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p339-russell/p339-

russell.pdf. ACM order number 508990.

Rabani:1996:DPS

[RT96] Yuval Rabani and Éva Tardos. Distributed packet switching in arbitrary networks. In ACM [ACM96], pages 366– 375. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p366-rabani/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        237814/p366-rabani/p366-rabani.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Raghavan:1994:ERA

[RU94] Prabhakar Raghavan and Eli Upfal. Efficient routing in all-optical networks. In ACM [ACM94], pages 134–

143. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p134-raghavan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p134-raghavan/p134-
                                        raghavan.pdf. ACM order no.

508930.

Raghavan:1995:SCR

[RU95] Prabhakar Raghavan and Eli Upfal. Stochastic contention resolution with short delays. In ACM [ACM95], pages 229–

237. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p229-raghavan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        225058/p229-raghavan/p229-
                                        raghavan.pdf. ACM order no.

508950.

Raz:1990:MCM

[RW90] R. Raz and A. Wigderson. Monotone circuits for matching require linear depth. In ACM [ACM90], pages 287–

292. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p287-raz/. ACM

order no. 508900.

Razborov:1997:ROB

[RWY97] Alexander Razborov, Avi Wigderson, and Andrew Yao. Readonce branching programs, rectangular proofs of the pigeonhole principle and the transversal calculus. In ACM [ACM97], pages 739–

748. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p739-razborov/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p739-razborov/p739-
razborov.pdf. ACM order no.

508970.

Safra:1992:EDA

[Saf92] Shmuel Safra. Exponential determinization for ω-automata with strong-fairness acceptance condition (extended abstract). In ACM [ACM92], pages 275–282. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p275-safra/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p275-
safra/p275-safra.pdf. ACM

order no. 508920.

Schulman:1992:SSU

[Sch92] Leonard J. Schulman. Sample spaces uniform on neighborhoods. In ACM [ACM92], pages 17–25. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p17-schulman/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p17-schulman/p17-schulman.

pdf. ACM order no. 508920.

Schulman:1993:DCI

[Sch93] Leonard J. Schulman. Deterministic coding for interactive communication. In ACM [ACM93], pages 747–756. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Schaefer:1999:GRT

[Sch99a] Marcus Schaefer. Graph Ramsey theory and the polynomial hierarchy. In ACM [ACM99], pages 592–601. ISBN 1-58113-067-8. LCCN QA75.5 .A14

                                        1999. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/301250/p592-schaefer/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p592-schaefer/p592-
                                        schaefer.pdf. ACM order

number 508990.

Schaeffer:1999:RSL

[Sch99b] Gilles Schaeffer. Random sampling of large planar maps and convex polyhedra. In ACM [ACM99], pages 760–

769. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/301250/p760-schaeffer/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p760-schaeffer/p760-
schaeffer.pdf. ACM order

number 508990.

Seidel:1992:APS

[Sei92]  Raimund Seidel.  On the

all-pairs-shortest-path problem. In ACM [ACM92], pages 745–749. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p745-seidel/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p745-seidel/p745-seidel.

pdf. ACM order no. 508920.

Servedio:1999:CSC

[Ser99]  Rocco A. Servedio.  Com-

putational sample complexity and attribute-efficient learning. In ACM [ACM99], pages 701–710. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p701-servedio/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p701-servedio/p701-
servedio.pdf. ACM order

number 508990.

Shapiro:1991:SCL

[Sha91] Ehud Shapiro. Separating concurrent languages with categories of language embeddings. In ACM [ACM91], pages 198–208. ISBN 0-89791-397-

3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p198-shapiro/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        103418/p198-shapiro/p198-

shapiro.pdf. ACM order no.

508910.

Shamir:1993:GMP

[Sha93] Adi Shamir. On the generation of multivariate polynomials which are hard to factor. In ACM [ACM93], pages 796–804. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Shoup:1990:SPR

[Sho90] V. Shoup. Searching for primitive roots in finite fields. In ACM [ACM90], pages 546–

554. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p546-shoup/. ACM

order no. 508900.

Sipser:1992:HSP

[Sip92] Michael Sipser. The history and status of the P versus NP question. In ACM [ACM92], pages 603–618. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p603-sipser/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        129712/p603-sipser/p603-sipser.

pdf. ACM order no. 508920.

Sitharam:1994:PGL

[Sit94]  Meera Sitharam.  Pseudo-

random generators and learning algorithms for AC. In ACM [ACM94], pages 478–

486. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p478-sitharam/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p478-sitharam/p478-
sitharam.pdf. ACM order no.

508930.

Slavik:1996:TAG

[Sla96] Petr Slavík. A tight analysis of the greedy algorithm for set cover. In ACM [ACM96], pages 435–441. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p435-slavik/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p435-slavik/p435-slavik.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Spielman:1995:LTE

[Spi95] Daniel A. Spielman. Lineartime encodable and decodable error-correcting codes. In ACM [ACM95], pages 388–

397. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p388-spielman/

;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p388-spielman/p388-
spielman.pdf. ACM order no.

508950.

Spielman:1996:FIT

[Spi96] Daniel A. Spielman. Faster isomorphism testing of strongly regular graphs. In ACM [ACM96], pages 576–584. ISBN 0-89791-785-5. LCCN QA 76.6

A13 1996. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/237814/p576-spielman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p576-spielman/p576-
spielman.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Srinivasan:1995:IAP

[Sri95]  Aravind Srinivasan.  Im-

proved approximations of packing and covering problems. In ACM [ACM95], pages 268–

276. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p268-srinivasan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
225058/p268-srinivasan/p268-
srinivasan.pdf. ACM order

no. 508950.

Schmidt:1990:ACH

[SS90a] J. P. Schmidt and A. Siegel. The analysis of closed hashing under limited randomness. In ACM [ACM90], pages 224–234. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p224-schmidt/.

ACM order no. 508900.

Schrift:1990:DLV

[SS90b] A. W. Schrift and A. Shamir.

The discrete log is very discreet. In ACM [ACM90], pages 405–415. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p405-schrift/.

ACM order no. 508900.

Simon:1992:CRV

[SS92] Janos Simon and Mario Szegedy.

On the complexity of RAM with various operation sets. In ACM [ACM92], pages 624–631. ISBN 0-89791-511-

9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p624-simon/;
http://www.acm.org/pubs/articles/
proceedings/stoc/129712/p624-
simon/p624-simon.pdf. ACM

order no. 508920.

Saks:1995:EDP

[SSZ95] Michael Saks, Aravind Srinivasan, and Shiyu Zhou. Explicit dispersers with polylog degree. In ACM [ACM95], pages 479–488. ISBN 0-89791-718-

9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/225058/p479-saks/;
http://www.acm.org/pubs/articles/

                                        proceedings/stoc/225058/p479-
                                        saks/p479-saks.pdf. ACM or-

der no. 508950.

Sundar:1990:UBS

[ST90] R. Sundar and R. E. Tarjan. Unique binary search tree representations and equalitytesting of sets and sequences. In ACM [ACM90], pages 18–25. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p18-sundar/. ACM

order no. 508900.

Srinivasan:1997:CFA

[ST97] Aravind Srinivasan and Chung-Piaw Teo. A constant-factor approximation algorithm for packet routing, and balancing local vs. global criteria. In ACM [ACM97], pages 636– 643. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/258533/p636-srinivasan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        258533/p636-srinivasan/p636-
                                        srinivasan.pdf. ACM order

no. 508970.

Shmoys:1997:AAF

[STA97] David B. Shmoys, Éva Tardos, and Karen Aardal. Approximation algorithms for facility location problems (extended abstract). In ACM [ACM97], pages 265–274. ISBN 0-89791-888-6. LCCN QA76.5 .A849

1997. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/258533/p265-shmoys/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p265-shmoys/p265-shmoys.

pdf. ACM order no. 508970.

Sudan:1999:PGX

[STV99] Madhu Sudan, Luca Trevisan, and Salil Vadhan. Pseudorandom generators without the XOR Lemma (extended abstract). In ACM [ACM99], pages 537–546. ISBN 1-58113-067-8. LCCN QA75.5 .A14

1999. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/301250/p537-sudan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p537-
sudan/p537-sudan.pdf. ACM

order number 508990.

Szegedy:1993:LBG

[SV93] Márió Szegedy and Sundar Vishwanathan. Locality based graph coloring. In ACM [ACM93], pages 201–207. ISBN 0-89791-591-7. LCCN QA 76.6

A13 1993. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/167088/p201-szegedy/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p201-szegedy/p201-

szegedy.pdf. ACM order no. 508930.

Sahinalp:1994:SBS

[SV94] Sühleyman Cenk Sahinalp and Uzi Vishkin. Symmetry breaking for suffix tree construction. In ACM [ACM94], pages 300–309. ISBN 0-89791-663-

8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p300-sahinalp/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
195058/p300-sahinalp/p300-
sahinalp.pdf. ACM order no.

508930.

Scheideler:1999:SDR

[SV99a] Christian Scheideler and Berthold Vöcking. From static to dynamic routing: efficient transformations of store-and-forward protocols. In ACM [ACM99], pages 215–224. ISBN 1-58113-067-8. LCCN QA75.5 .A14

1999. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/301250/p215-scheideler/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p215-scheideler/p215-
scheideler.pdf. ACM order

number 508990.

Schulman:1999:MSH

[SV99b] Leonard J. Schulman and Umesh V. Vazirani. Molecular scale heat engines and scalable quantum computation. In ACM [ACM99], pages 322–329. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p322-schulman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p322-schulman/p322-
schulman.pdf. ACM order

number 508990.

Schulman:1999:MEA

[SV99c] Leonard J. Schulman and Vijay V. Vazirani. Majorizing estimators and the approximation of #p-complete problems. In ACM [ACM99], pages 288–294. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p288-schulman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p288-schulman/p288-
schulman.pdf. ACM order

number 508990.

Shokrollahi:1998:DAG

[SW98] M. Amin Shokrollahi and Hal Wasserman. Decoding algebraic-geometric codes beyond the error-correction bound. In ACM [ACM98], pages 241–248. ISBN 0-89791-962-9. LCCN QA75.5 .A14

1998. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/276698/p241-shokrollahi/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p241-shokrollahi/p241-
shokrollahi.pdf. ACM order

number 508980.

Skutella:1999:PMW

[SW99] Martin Skutella and Gerhard J. Woeginger. A PTAS for minimizing the weighted sum of job completion times on parallel machines. In ACM [ACM99], pages 400–

407. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p400-skutella/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        301250/p400-skutella/p400-
                                        skutella.pdf. ACM order

number 508990.

Saks:1993:WFS

[SZ93] Michael Saks and Fotios Zaharoglou. Wait-free k-set agreement is impossible: the topology of public knowledge. In ACM [ACM93], pages 101–

110. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/167088/p101-saks/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/167088/p101-
                                        saks/p101-saks.pdf. ACM or-

der no. 508930.

Szegedy:1990:FBS

[Sze90] M. Szegedy. Functions with bounded symmetric communication complexity and circuits with mod m gates. In ACM [ACM90], pages 278–

286. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/100216/p278-szegedy/.

ACM order no. 508900.

Thathachar:1998:SRK

[Tha98] Jayram S. Thathachar. On separating the read-k-times branching program hierarchy. In ACM [ACM98], pages 653–

662. ISBN 0-89791-962-9.

LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p653-thathachar/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p653-thathachar/p653-
thathachar.pdf. ACM order

number 508980.

Tanaka:1994:CNL

[TN94] Keisuke Tanaka and Tetsuro Nishino. On the complexity of negation-limited Boolean networks. In ACM [ACM94], pages 38–47. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p38-tanaka/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p38-
tanaka/p38-tanaka.pdf. ACM

order no. 508930.

Trevisan:1997:WHM

[Tre97] Luca Trevisan. When Hamming meets Euclid: the approximability of geometric TSP and MST (extended abstract). In ACM [ACM97], pages 21–

29. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p21-trevisan/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
258533/p21-trevisan/p21-trevisan.

pdf. ACM order no. 508970.

Trevisan:1998:RQP

[Tre98]  Luca Trevisan.  Recycling

queries in PCPs and in linearity tests (extended abstract). In ACM [ACM98], pages 299–

308. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p299-trevisan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p299-trevisan/p299-
trevisan.pdf. ACM order

number 508980.

Trevisan:1999:CEU

[Tre99] Luca Trevisan. Construction of extractors using pseudo-random generators (extended abstract). In ACM [ACM99], pages 141–

148. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p141-trevisan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
301250/p141-trevisan/p141-
trevisan.pdf. ACM order

number 508990.

Ta-Shma:1996:ERW

[TS96]  Amnon Ta-Shma.  On ex-

tracting randomness from weak random sources (extended abstract). In ACM [ACM96], pages 276–285. ISBN 0-89791-785-5. LCCN QA 76.6 A13

1996. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/237814/p276-ta-shma/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p276-ta-shma/p276-

ta-shma.pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Ta-Shma:1998:AOD

[TS98] Amnon Ta-Shma. Almost optimal dispersers. In ACM [ACM98], pages 196–202. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998. URL http://www.

acm.org/pubs/citations/proceedings/
stoc/276698/p196-ta-shma/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p196-ta-shma/p196-

ta-shma.pdf. ACM order number 508980.

Therien:1998:WTV

[TW98] Denis Thérien and Thomas Wilke. Over words, two variables are as powerful as one quantifier alternation. In ACM [ACM98], pages 234–

240. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p234-therien/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p234-therien/p234-

therien.pdf. ACM order number 508980.

Valiant:1999:RL

[Val99] Leslie G. Valiant. Robust logics. In ACM [ACM99], pages 642–651. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p642-valiant/;
http://www.acm.org/pubs/
articles/proceedings/stoc/

301250/p642-valiant/p642-

valiant.pdf. ACM order number 508990.

Vardy:1997:ACC

[Var97] Alexander Vardy. Algorithmic complexity in coding theory and the minimum distance problem. In ACM [ACM97], pages 92–109. ISBN 0-89791-888-

6. LCCN QA76.5 .A849 1997.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/258533/p92-vardy/;
http://www.acm.org/pubs/articles/
proceedings/stoc/258533/p92-
vardy/p92-vardy.pdf. ACM

order no. 508970.

Vishkin:1990:DSN

[Vis90]  U. Vishkin.  Determinis-

tic sampling — a new technique for fast pattern matching. In ACM [ACM90], pages 170–180. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p170-vishkin/.

ACM order no. 508900.

Vorobjov:1998:CLD

[Vor98]  Nicolai Vorobjov.  Com-

puting local dimension of a semialgebraic set. In ACM [ACM98], pages 483–

487. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p483-vorobjov/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
276698/p483-vorobjov/p483-

vorobjov.pdf. ACM order

number 508980.

Venkatesan:1992:ACI

[VR92] Ramarathnam Venkatesan and Sivaramakrishnan Rajagopalan. Average case intractability of matrix and Diophantine problems (extended abstract). In ACM [ACM92], pages 632– 642. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/129712/p632-venkatesan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
129712/p632-venkatesan/p632-
venkatesan.pdf. ACM order

no. 508920.

Vitter:1990:ODP

[VS90] J. S. Vitter and E. A. M.

Shriver. Optimal disk I/O with parallel block transfer. In ACM [ACM90], pages 159–169. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p159-vitter/.

ACM order no. 508900.

Vengroff:1996:EDR

[VV96] Darren Erik Vengroff and Jeffrey Scott Vitter. Efficient 3-D range searching in external memory. In ACM [ACM96], pages 192– 201. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p192-vengroff/

                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p192-vengroff/p192-
                                        vengroff.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Vavasis:1994:AIP

[VY94] Stephen A. Vavasis and Yinyu Ye. An accelerated interior point method whose running time depends only on A (extended abstract). In ACM [ACM94], pages 512– 521. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p512-vavasis/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p512-vavasis/p512-

vavasis.pdf. ACM order no. 508930.

vonzurGathen:1993:CCT

[vzGKS93] Joachim von zur Gathen, Marek Karpinski, and Igor Shparlinski. Counting curves and their projections. In ACM [ACM93], pages 805–812. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

vonzurGathen:1992:CFM

[vzGS92] Joachim von zur Gathen and Victor Shoup. Computing Frobenius maps and factoring polynomials. In ACM [ACM92], pages 97–105. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/

stoc/129712/p97-von_zur_gathen/
p97-von_zur_gathen.pdf;
http://www.acm.org/pubs/citations/
proceedings/stoc/129712/p97-
von_zur_gathen/. ACM order

no. 508920.

Wang:1995:ACC

[Wan95] Jie Wang. Average-case completeness of a word problem for groups. In ACM [ACM95], pages 325–334. ISBN 0-89791-718-9. LCCN QA 76.6 A13

1995. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/225058/p325-wang/;
http://www.acm.org/pubs/articles/
proceedings/stoc/225058/p325-
wang/p325-wang.pdf. ACM or-

der no. 508950.

Wayne:1999:PCA

[Way99] Kevin D. Wayne. A polynomial combinatorial algorithm for generalized minimum cost flow. In ACM [ACM99], pages 11–18. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/301250/p11-wayne/;
http://www.acm.org/pubs/articles/
proceedings/stoc/301250/p11-
wayne/p11-wayne.pdf. ACM

order number 508990.

Williamson:1993:PDA

[WGMV93] David P. Williamson, Michel X.

Goemans, Milena Mihail, and Vijay V. Vazirani. A primaldual approximation algorithm for generalized Steiner network problems. In ACM [ACM93], pages 708–717. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

Wigderson:1994:APP

[Wig94] Avi Wigderson. The amazing power of pairwise independence (abstract). In ACM [ACM94], pages 645–647. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/195058/p645-wigderson/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        195058/p645-wigderson/p645-
                                        wigderson.pdf. ACM order no.

508930.

Wilson:1996:GRS

[Wil96] David Bruce Wilson. Generating random spanning trees more quickly than the cover time. In ACM [ACM96], pages 296–303. ISBN 0-89791-785-

5. LCCN QA 76.6 A13 1996.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/237814/p296-wilson/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        237814/p296-wilson/p296-wilson.

pdf. ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

Wigderson:1993:EBE

[WZ93] Avi Wigderson and David Zuckerman. Expanders that beat the eigenvalue bound: explicit construction and applications. In ACM [ACM93], pages 245–251. ISBN 0-89791-591-

7. LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/167088/p245-wigderson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
167088/p245-wigderson/p245-
wigderson.pdf. ACM order no.

508930.

Yang:1994:TBM

[YA94] Jae-Heon Yang and James H.

Anderson. Time bounds for mutual exclusion and related problems. In ACM [ACM94], pages 224–233. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p224-yang/;
http://www.acm.org/pubs/articles/
proceedings/stoc/195058/p224-
yang/p224-yang.pdf. ACM or-

der no. 508930.

Yao:1990:CFP

[Yao90]  A. C.-C. Yao.  Coherent

functions and program checkers. In ACM [ACM90], pages 84–94. ISBN 0-89791-361-

2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p84-yao/. ACM

order no. 508900.

Yao:1994:DTC

[Yao94] Andrew Chi-Chih Yao. Decision tree complexity and Betti numbers. In ACM [ACM94], pages 615–624. ISBN 0-89791-663-8. LCCN QA76 .A15 1994.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/195058/p615-yao/;   http:

                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/195058/p615-
                                        yao/p615-yao.pdf. ACM order

no. 508930.

Yao:1995:SQP

[Yao95] Andrew Chi-Chih Yao. Security of quantum protocols against coherent measurements. In ACM [ACM95], pages 67– 75. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/225058/p67-yao/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/225058/p67-
                                        yao/p67-yao.pdf. ACM order

no. 508950.

Cai:1992:PCH

[yC92] Jin yi Cai. Parallel computation over hyperbolic groups. In ACM [ACM92], pages 106–115. ISBN 0-89791-511-9. LCCN QA76.A15 1992.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/129712/p106-cai/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/129712/p106-
                                        cai/p106-cai.pdf. ACM order

no. 508920.

Yannakakis:1991:TFS

[YL91] Mihalis Yannakakis and David Lee. Testing finite state machines. In ACM [ACM91], pages 476–485. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/103418/p476-yannakakis/
                                        ;   http://www.acm.org/pubs/

articles/proceedings/stoc/
103418/p476-yannakakis/p476-
yannakakis.pdf. ACM order

no. 508910.

Zuckerman:1990:TLB

[Zuc90] D. Zuckerman. A technique for lower bounding the cover time. In ACM [ACM90], pages 254–259. ISBN 0-89791-361-2. LCCN QA76.A15 1990.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/100216/p254-zuckerman/

. ACM order no. 508900.

Zuckerman:1996:ROS

[Zuc96] David Zuckerman. Randomnessoptimal sampling, extractors, and constructive leader election. In ACM [ACM96], pages 286–295. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/237814/p286-zuckerman/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
237814/p286-zuckerman/p286-
zuckerman.pdf. ACM order no.

508960. Also known as Federated Computing Research Conference (FCRS ’96).

Zwick:1998:FAS

[Zwi98] Uri Zwick. Finding almostsatisfying assignments. In ACM [ACM98], pages 551– 560. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/276698/p551-zwick/;

                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/276698/p551-
                                        zwick/p551-zwick.pdf. ACM

order number 508980.

Zwick:1999:APL

[Zwi99a] Uri Zwick. All pairs lightest shortest paths. In ACM [ACM99], pages 61–69. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. URL http://www.

                                        acm.org/pubs/citations/proceedings/
                                        stoc/301250/p61-zwick/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p61-
                                        zwick/p61-zwick.pdf. ACM

order number 508990.

Zwick:1999:ORT

[Zwi99b] Uri Zwick. Outward rotations: a tool for rounding solutions of semidefinite programming relaxations, with applications to MAX CUT and other problems. In ACM [ACM99], pages 679–687. ISBN 1-58113-067-

8. LCCN QA75.5 .A14 1999.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/301250/p679-zwick/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/301250/p679-
                                        zwick/p679-zwick.pdf. ACM

order number 508990.